### v1

Effect of State and Local Sexual Orientation Anti-Discrimination Laws on Labor Market Differentialsthanks: We would like to thank Sandra Black, Rich Murphy, Steve Trejo, Gerald Oettinger, Dan Hamermesh, Kitt Carpenter, Dario Sansone, Chris Jepsen, Luis Faundez, and all the seminar participants of the University of Texas at Austin. No financial support was given to support this research. Correspondence: d.vamossy@pitt.edu.
Scott Delhommer University of Texas at Austin; sdelhommer@utexas.edu    Domonkos F. Vamossy University of Pittsburgh, d.vamossy@pitt.edu
(April 4, 2024)
This paper presents the first quasi-experimental research examining the effect of both local and state anti-discrimination laws on sexual orientation on the labor supply and wages of lesbian, gay, and bisexual (LGB) workers. To do so, we use the American Community Survey data on household composition to infer sexual orientation and combine this with a unique panel dataset on local anti-discrimination laws. Using variation in law implementation across localities over time and between same-sex and different-sex couples, we find that anti-discrimination laws significantly reduce gaps in labor force participation rate, employment, and the wage gap for gay men relative to straight men. These laws also significantly reduce the labor force participation rate, employment, and wage premium for lesbian women relative to straight women. One explanation for the reduced labor supply and wage premium is that lesbian couples begin to have more children in response to the laws. Finally, we present evidence that state anti-discrimination laws significantly and persistently increased support for same-sex marriage. This research shows that anti-discrimination laws can be an effective policy tool for reducing labor market inequalities across sexual orientation and improving sentiment toward LGB Americans.
JEL Codes: J31, J71, J78, K31
1 Introduction
On April 15, 2020, the Supreme Court of the United States ruled that federal employment discrimination law applied to lesbian, gay, bisexual, and transgender (LGBT) individuals. Before this ruling, the legislation in 25 states made it legal for private businesses to fire an employee for being LGBT. Unlike race, sex, age, and religion, federal law did not include sexual orientation as a protected class for federal anti-discrimination laws until 2020. As a result, laws extending protected class status based on sexual orientation varied between states and localities. The first state to pass an anti-discrimination law based on sexual orientation was Wisconsin in 1982, and the most recent state to pass such a law was Pennsylvania and Michigan in 2023.111Since August 2018, Pennsylvania has interpreted discrimination based on sexual orientation and gender identity as being covered under the “sex” category of the Pennsylvania Human Relations Act. This allowed LGBT individuals to file complaints for discrimination in employment, housing, education, and public accommodations. Following Michigan’s lead in May 2018, Pennsylvania became the second state to implement statewide LGBT protections in this manner. There has been a renewed push to enact sexual orientation protections at the legislative level as Democrats in the House of Representatives introduced “The Equality Act” in early 2019, which would add sexual orientation and gender identity as federally protected characteristics. The focus of this paper is the effect of local and state sexual orientation anti-discrimination legislation on labor market differentials for lesbian, gay, and bisexual (LGB) workers.
We exploit the differential roll-out of state and local laws from 2005-2019 in a modern triple difference-in-differences framework to analyze how these anti-discrimination laws impact the wages and labor supply of LGB workers. The economics literature on LGB workers consistently finds that gay/bisexual men have a pay and labor supply penalty and lesbian/bisexual women have a pay and labor supply premium over their heterosexual counterparts222(Badgett (1995); Klawitter and Flatt (1998); Allegretto and Arthur (2001); Black et al. (2003); Carpenter (2005); Black et al. (2007); Antecol et al. (2008); Klawitter (2015); Jepsen and Jepsen (2017); Carpenter and Eppink (2017)). The differences in pay are attributed to a host of factors including discrimination and intra-household labor allocation and specialization. We test the role of discrimination by examining how the passage of anti-discrimination laws affects the labor supply and wage differences between LGB and heterosexual workers.
There have been previous studies examining how these state and local laws affect the pay gap, but these studies have been limited in their ability to identify causal effects (Klawitter and Flatt (1998); Gates (2009)). Klawitter and Flatt (1998) compared people in same-sex partnerships to those in different-sex partnerships in the 1990 Census and found that same-sex couples have no significant difference in wages in places with anti-discrimination laws. Gates (2009) uses the 2000 Census and conducts a similar analysis. Gates (2009) finds that gay/bisexual men in places with anti-discrimination laws have a 3% wage premium over gay/bisexual men in places without these laws, and lesbian/bisexual women have a 2% wage premium over lesbian/bisexual women in places without these laws [^3] Tilcsik (2011) found that resumes signaling LGB status received significantly fewer callbacks in localities without sexual orientation anti-discrimination laws. These past studies on anti-discrimination laws, while informative, fail to fully account for changes over time.
Previous research has leveraged state-level sexual orientation anti-discrimination laws within a difference-in-differences framework. [^4] The Williams Institute found that sexual orientation discrimination occurs at a similar rate to sex-based discrimination using state-level complaint data following state anti-discrimination laws (Ramos et al., 2008). Martell (2013) study demonstrated that such laws have significantly reduced wage disparities for gay men by 20%.555 Since Martell (2013) and Klawitter and Flatt (1998) there has been a booming literature on difference-in-differences. Accordingly, unlike Martell (2013) and Klawitter and Flatt (1998), we employ modern difference-in-difference estimators to alleviate biases stemming from staggered policy design. Our research extends beyond the scope of previous studies by incorporating an examination of local laws, thereby offering a more comprehensive understanding of the legislative landscape. Furthermore, this study also estimates the effects of such laws on women, a demographic previously under-explored in this context. Our study expands the scope of Klawitter and Flatt (1998) by leveraging a larger dataset, utilizing American Community Survey (ACS) data spanning from 2004 to 2019, in contrast to Klawitter and Flatt (1998) which relied on the 1990 U.S. census data encompassing approximately 13,000 married couples, 14,500 different-sex unmarried couples, and 6,800 same-sex couples. The ACS data allows for a robust analysis across a broader temporal spectrum, capturing a wide array of demographic and economic changes over 15 years. This extensive dataset, combined with the application of modern difference-in-difference estimators, enables a more detailed and dynamic examination of the effects of anti-discrimination policies on the earnings and household incomes of same-sex couples. Additionally, Burn (2018)’s more recent examination in a triple difference-in-differences framework confirms these findings but also underscores a gap in the literature: the lack of consideration for local anti-discrimination ordinances. Previous studies have not fully addressed potential biases arising from variations in the timing of these laws’ enactment. To bridge this gap, our research employs the methodology proposed by Borusyak et al. (2021), comparing it with the conventional triple difference-in-differences estimator, finding both approaches yield similar results.
This paper is the first to examine both local and state-level sexual orientation anti-discrimination laws in a quasi-experimental design. This paper is also the first to analyze the effects of any local anti-discrimination laws on sex, race, or sexual orientation in a modern difference-in-difference framework [^6] Using state variation in anti-discrimination laws has also been analyzed to understand racial- and sex-based discrimination (Neumark and Stock (2006); (Donohue III and Heckman, 1991);((Margo, 1995); Goldin and Margo (1992)). However, these studies on sex and race anti-discrimination laws also omit local laws, potentially skewing their estimates..
We use the 2005-2019 American Community Survey (ACS) and household composition to infer sexual orientation and create a unique and novel panel dataset on the passage of local and state anti-discrimination laws. We collected information on local laws from a host of sources including media reports, FOIA requests, and an advocacy group. We find a significant reduction in differences between LGB workers and heterosexual workers across labor supply and wage measures, due to anti-discrimination laws. Anti-discrimination laws significantly reduce the gap in labor force participation and employment of gay men by 2.1 p.p. and 1.5 p.p., respectively. The laws also significantly increase hourly wages for gay men by 6.2%. The results differ for women, with the laws significantly reducing their labor force participation and employment premium over straight women by 1.7 p.p., and 2.3 p.p., respectively. Using an event study plot, we show that the outcome trends are parallel before the implementation of the anti-discrimination laws suggesting the workers in treatment and control regions are reasonable comparisons. [^7] In Appendix A, we replicate the findings of past research that finds gay/bisexual men have lower labor force participation and employment rates and make 8-11% less than their employed straight counterparts using hourly wages and annual earnings. We also replicate findings that lesbian/bisexual women have a higher labor force participation and employment rate and earn 5-15% more than their employed straight counterparts.
We explore theories for the differing effects of anti-discrimination laws on gay men and lesbian women in the Discussion section, using the Becker (1981) model of household specialization. In the traditional Becker (1981) model of household specialization, men typically specialize in market production, and women typically specialize in household production, in part due to differences in biology where women birth and care for children, resulting in a one-earner household. These differences in household specialization are less pronounced in same-sex partnerships, but they may become more similar for women in same-sex partnerships following the passage of an anti-discrimination law if it gives greater protection to the higher wage earner. We show empirically that the difference in hours worked between partners within a lesbian household goes up relative to gay households following anti-discrimination laws, suggesting that lesbian households could become more specialized following anti-discrimination laws with one woman working more hours and the other woman working fewer to focus on household production. We also show that lesbian households have significantly more children than gay households after the passage of an anti-discrimination law. More children could induce lesbian households to further specialize in the intra-household division of labor, adopting a more traditional household model to help care for additional children.
Finally, our last contribution is the use of polling data on the support for same-sex marriage to examine the relationship between anti-discrimination laws and state sentiment toward LGB workers. One may expect that sentiment toward LGB workers would increase right before the passage of the laws, creating a selection issue. Alternatively, the anti-discrimination laws may normalize being a sexual minority, and improve sentiments toward LGB workers. We collected every poll on same-sex marriage for each state and year from the Pew Research Center. We show that the passage of state-wide sexual orientation anti-discrimination laws persistently increases favorability toward LGB people through increased support for same-sex marriage. The states that passed anti-discrimination laws had parallel pre-trends in support of same-sex marriage before the passage of the laws. Afterward, those states had a significant and persistent increase in their support for same-sex marriage. The observed increase in support following the enactment of the law, rather than before, implies that policy changes can indeed influence public opinion. This notion aligns with the findings of Sansone (2019), suggesting that legislative actions can lead to shifts in societal attitudes. This is in contrast to the thermostatic model proposed by Wlezien (1995), which posits that public sentiment drives policy changes rather than the other way around. Increased favorability toward LGB people may act in conjunction with greater job security to improve labor market outcomes for LGB Americans.
2 Data
A common issue in analyzing the pay gap/premium between homosexual and heterosexual workers is a lack of high-quality data about sexual orientation, wages, and employment. We follow the literature (Klawitter and Flatt (1998); Gates (2009); Jepsen and Jepsen (2017)) in inferring sexual orientation by looking at household composition. Specifically, we infer a person’s sexual orientation to be homosexual/bisexual if they have an “unmarried partner” or a spouse that is the same sex as themselves. [^8] We exclude individuals whose sex or relationship status was allocated following (Gates and Steinberger, 2009).. The ACS references a close personal relationship for unmarried partners as well as giving an option for “partner/roommate”, which reduces the likelihood that straight roommates would misclassify as unmarried partners.
The comparisons that we make are between people in same-sex relationships, defined as being in an unmarried partnership or married with someone of the same sex, and those in different-sex relationships. This comparison based on household composition, though standard in the literature, is not equivalent to comparing LGB and heterosexual workers. It compares those who are in a same-sex relationship to those who are in a different-sex relationship. Notably, this comparison excludes all workers who are not in a cohabitating relationship and classifies bisexual people in different-sex relationships as members of the control group. For example, if a person is misclassified as straight instead of bisexual, and bisexual workers have lower pay than their straight counterparts, then the misclassification would bias our pay gap/premium results toward zero.
We use the 2005-2019 yearly ACS from (Ruggles et al., 2019). We use wages, defined as salaried wages from an employer. The ACS reports wage earnings in annual terms. We use the annual figures and convert them to hourly wage figures using variables on average weekly hours worked and weeks worked999“Weeks worked” in the ACS is a categorical variable giving a range of weeks worked. We take the median value given in the range to compute weeks worked.. Finally, we limit our sample to prime-age working adults and only examine those who are ages 25-65. We use 25 as the lower age cutoff to allow workers to complete college and have more fully developed human capital, which is typically thought of as an important factor for the differences in wages between LGB and straight workers (Black et al., 2007).
We obtained data on the passage of state anti-discrimination laws from LGBTMap.org, an LGBT advocacy group. The website gives information on which states passed anti-discrimination laws and when. We focus solely on sexual orientation anti-discrimination laws that give protection in employment. LGBTMap.org also provided incomplete data detailing the passage of local city and county laws with many cities missing years for the start of their anti-discrimination laws. We completed the dataset for the relevant years using old media reports, correspondence with local officials, and FOIA requests. This dataset is the first complete city-level dataset on sexual orientation anti-discrimination laws in the U.S. We focus on cities reported in the ACS and match those city laws with their corresponding counties to merge with the ACS. The ACS only reports county of residence for those in metro areas, so any rural counties are lumped together. We make the assumption in our analysis that the impact of these laws is the same across states and localities, estimating an average effect.
We report the state, city, and county laws and the year they were enacted in Table 1. In Figure 1, we show which counties had sexual orientation anti-discrimination laws in 2005 and 2019, respectively. In our analysis, laws are coded based on their initial application to private employment and housing, focusing on broad protections rather than sector-specific ones. For instance, while Delaware extended protections to state employees in 1999, our coding reflects 2009, the year broader protections were enacted. This approach could overlook the incremental impact of earlier, narrower protections on societal attitudes. The same principle applies to local laws, where initial protections for public employees in some cities might precede more comprehensive legislation, potentially influencing early shifts in public sentiment towards sexual minorities. The enforcement mechanisms of anti-discrimination laws, such as fines, mandatory training, and public rebukes, serve as significant deterrents, encouraging the adoption of more inclusive employment policies. The rise in successful legal challenges against companies concerning discrimination highlights the real legal and financial consequences they encounter, reinforcing the laws’ effectiveness in promoting workplace equality (for Human Resource Management (2023)).
The enactment of anti-discrimination legislation is not random. The states that have these protections are generally considered friendlier to LGB workers than those without these laws and have a larger share of same-sex partnerships than those states without these laws. Also, many counties with anti-discrimination laws have large cities with a larger concentration of LGB workers than rural counties. However, this is not universal. For instance, Utah extended protection to LGB workers despite being a relatively conservative state, and certain liberal cities like Houston [^10] Houston passed a sexual orientation and gender identity anti-discrimination ordinance, but it was only in effect for 3 months before being challenged. The law was put up to a public vote and lost, repealing the law. are noticeably absent from extending protection to LGB workers. One potential concern in this analysis is the endogenous adoption of laws. The areas that are friendly and less hostile to LGB workers may be the areas that are most likely to adopt sexual orientation anti-discrimination laws. The main threat to identification is that the timing of anti-discrimination laws is correlated with another factor causing an omitted variable bias. It could be that positive sentiment toward LGB workers causes both the law change and any change in labor market outcomes. We mitigate this by adopting a saturated triple-difference model, including all two-way fixed effects (county-year, year-SSP, and county-SSP), as proposed by (Burn, 2018). The inclusion of county-year fixed effects allows us not to include county controls hence eliminating the need to control for unavailable local levels of support for same-sex marriage. We refine the approach of (Burn, 2018) by employing the methodology of (Borusyak et al., 2021), particularly suited for analyzing staggered policy adoptions like ours. This method accurately estimates the effects of anti-discrimination laws, considering their varied implementation times. For additional validation, we include results from a simpler difference-in-differences estimator in the appendix, providing robustness to our findings.
In Figure 2, we present the distribution of same-sex partnerships by state from 2005 to 2019, and in Table 2, we present the 10 counties with the largest share of same-sex partnerships. Unsurprisingly, the distribution of same-sex partnerships by state and by county is skewed toward more progressive states and counties with large cities in them that are known for having a large LGB population like San Francisco, the District of Columbia, New York City, and Boston. The LGB population varies from the straight population in many dimensions including geography.
We present descriptive statistics showing the differences in the LGB population compared to the straight population broken down by education level and sex in Table 3. Descriptively, there is a clear difference in labor market outcomes and characteristics between men and women in same-sex partnerships and those in different-sex partnerships when controlling for education.
3 Identification Strategy & Estimation
The differential roll-out of anti-discrimination laws by state and locality over time and between same-sex and different-sex couples lends itself to a saturated triple difference-in-differences framework with the main outcomes of interest as the change in labor supply and pay. This strategy uses the variation presented in Figure 1, which shows how the laws changed over time by state and county. Formally, we estimate this equation on labor supply:
$\begin{split}Y_{ict}=\alpha+\beta_{0}(SSP_{ict}\times ENDA_{ct}\times I_{\textrm{Sex=Female})}+\beta_{1}(SSP_{ict}\times ENDA_{ct}\times I_{\textrm{Sex=Male})}\\ +(I_{c}\times I_{t}\times I_{\textrm{Sex=Male}}))\gamma_{sct}+(SSP_{ict}\times I_{c}\times I_{\textrm{Sex=Male}}))\theta_{S,s,c}\\ +(SSP_{ict}\times I_{t}I_{\textrm{Sex=Male}}))\theta_{S,s,t}+X^{\prime}_{ict}\delta+\epsilon_{ict}\end{split}$ (1)
In our regression, equation (1) defines the labor market outcomes $Y_{ict}$ for individual $i$, in county $c$, at time $t$ with sex $s$ and same-sex partnership status $S$. These include the labor force participation, employment status, or the real value of wages, in 1999 dollars. The binary variable $SSP_{ict}$ indicates if an individual is in a same-sex partnership. The model explores the interaction between being in a same-sex partnership and the presence of county-level Employment Non-Discrimination Acts $(SSP_{ict}\times ENDA_{ct})$. [^11] Treatment status is computed using the earliest year of ENDAs at the county or state level, with individuals remaining classified as “treated” thereafter throughout the dataset. Given the ACS’s 23-month reference period for earnings, as discussed by Hansen et al. (2020), there’s a potential for bias in treatment assignment. This extended timeframe might not precisely align with the exact implementation dates of ENDAs, potentially biasing the immediate effects of these policies towards zero. The fixed effects for county-year-sex interactions $(\gamma_{ct})$, control for unobserved heterogeneity. Interactions between the same-sex partnership status, sex and county $(SSP_{ict}\times I_{c}\times I_{\textrm{Sex=Male}})$ and year fixed effects $(SSP_{ict}\times I_{t}\times I_{\textrm{Sex=Male}})$ capture varying impacts over time and across different regions by sex. $X^{\prime}_{ict}$ is a vector of person-specific covariates including race, children, age, and education level - each represented up to a quadratic term to capture potential nonlinear effects. The error term $\epsilon_{ict}$ is clustered at the county level to account for within-group correlation. [^12] Since the ACS only identifies metro counties, the county fixed effect is close to a pseudo-county fixed effect. The rural counties are all compared together with a true county fixed effect for metro area counties. This specification aims to isolate the effects of anti-discrimination laws on the labor market outcomes for individuals in same-sex partnerships by sex.
Our sample is limited to people aged 25 to 65, and we estimate this model jointly by sex using the ACS person weights. When examining wages, the $\ln(Y_{ict}+1)$ specification is adopted to include both employed and unemployed individuals in the analysis, recognizing employment as endogenous to anti-discrimination laws. This approach ensures a comprehensive evaluation of the law’s impact on the entire income spectrum.
In equation (1), the differential effect of being a cohabiting same-sex couple by sex in the United States over time is captured by $\theta_{S,s,t}$. The county fixed effects for cohabiting same-sex couples by sex $\theta_{S,s,c}$ capture how the wage differential varies across counties by sex. The parameters of interest is $\beta_{0}$ and $\beta_{1}$, which capture how the differential between homosexuals and heterosexuals in counties with ENDAs changed after the law’s passage by sex. Hence, $\beta_{1}$, will give the effect of ENDAs on the labor supply and pay gap/premium for males, and $\beta_{0}$ for females. [^13] We recognize the heterogeneous impact of anti-discrimination laws across different regions, yet our analysis aims to estimate an average effect for those directly subject to the laws.
One concern is that the distribution of LGB workers is not random, and states/counties that pass anti-discrimination laws have a higher concentration of LGB workers. However, using county fixed effects will make comparisons within a given county, so San Francisco will be compared to San Francisco. A potential concern would be if there is sorting of LGB workers to areas with anti-discrimination laws protecting them, changing the composition of the area. We show that there is little evidence to support this in the robustness check section.
To interpret $\beta_{0}$ and $\beta_{1}$ as being the causal effect of enacting employment protections for sexual orientation, it must be the case that there are no other factors related to changes in the gay/lesbian-heterosexual wage differential in states that do and do not have LGBT employment protections. To accurately estimate $\beta_{0}$ and $\beta_{1}$ while considering the staggered implementation of ENDAs, we employ the methodology from (Borusyak et al., 2021). This approach uses observations only for units and periods that are not yet treated to infer the never-treated potential outcome for each treated unit.
Our identifying assumption in this framework is that, in the absence of anti-discrimination laws, the trends in labor supply and wage disparities in counties with such laws would have mirrored those without them. This difference-in-differences framework hinges on parallel trends across all groups and periods, assuming no prior anticipation of the laws. While directly testing this parallel trend assumption is challenging, examining pre-treatment trends can provide some indication of its validity. We examine the outcomes relative to the timing of the laws by estimating an event study model, where we estimate coefficients for each period relative to the treatment time. This approach allows us to assess the impact of the laws over time and ensure a thorough understanding of their effects before and after implementation.
We plot the effect of the law on the labor supply and pay gap/premium in each year relative to the start date. The effect of the law in the year before the law is enacted is normalized to zero. We report the coefficients estimated via Borusyak et al. (2021) in the main text, and standard TWFE in the Appendix.
This paper’s main contribution is applying a staggered policy design approach, providing an understanding of the average effects of anti-discrimination laws. Our approach highlights the importance of considering local contexts when assessing the efficacy of legislative measures.
4 Primary Findings
4.1 Extensive & Intensive Margins, Wages
In Table 4, we present the regression on the extensive & intensive margins of labor supply with $\beta_{1}$ giving the effect of anti-discrimination laws on gay men’s labor supply and $\beta_{0}$ giving the effect for lesbian women. Columns (1-2) show that anti-discrimination laws have a significant effect in reducing the labor force participation and employment gap for gay men. The labor force participation rate gap is reduced by 2.1 p.p., and the employment gap is reduced by 1.5 p.p. Conversely, the labor force participation and employment premium that lesbian women have over their straight counterparts is reduced by 1.7 p.p. and 2.3 p.p., respectively. Columns (3-6) present the regression results on the intensive margin of labor supply as measured by weekly hours worked and weeks worked. We find contrasting effects of anti-discrimination laws on male and female outcomes. We find that the intensive labor supply gap for gay men as measured by weekly hours worked and weeks worked is insignificantly reduced, and the intensive labor supply premium for lesbian women is significantly reduced when measured in weeks worked. We later show that the positive impact for lesbian women on weekly hours and the negative effect on weeks worked for gay men are gone once controlled for occupational fixed effects.
In Table 5, we present the regression results on the wage gap as measured by hourly wages and annual wage earnings. These variables received a $\ln(Y_{ict}+1)$ transformation, to handle zero values. [^14] We see small differences in effect sizes in a log-only specification on the employed-only population. We run the regressions jointly by sex and present the regressions showing how the sexual orientation anti-discrimination laws impact both employed and unemployed people in columns (1) and (2) and limit the regressions to only those who are employed in columns (3) and (4). Our preferred specification is to look at both employed and unemployed people in the sample as employment is endogenous. Conditioning on employed workers would create an endogeneity problem since labor force participation and employment are significantly impacted by anti-discrimination laws. Given Table 4, the sample of employed workers could be changing as a result of the law changes, which would lead to a change in the sample composition. We present the regressions limited to employed workers to replicate past work on the wage gap/premium where it is common to only examine employed workers.
Gay men experience a notable increase in hourly wages of 6.2%, which is statistically significant at the 95% confidence level. Additionally, there is a substantial increase of 22.2% concerning their annual earnings, indicating significant progress towards income equality. On the other hand, lesbian women see an insignificant change in their annual earnings premium and an insignificant reduction in their hourly wage premium. The pay gap/premium is reduced quite significantly once the sample is limited to employed workers in columns (3) and (4), which is unsurprising. The effects also turn insignificant, which shows that they come from increased labor force participation and employment rates.
4.2 Occupation Fixed Effects
LGB workers tend to sort to different occupations and industries based on sex and sexual orientation (Black et al., 2007). They show that gay men tend to sort to more female-dominated fields and lesbian women sort to more male-dominated fields relative to their heterosexual counterparts. It is typical to not consider those factors because discrimination affects occupation and industry choice, which would make occupation and industry choice endogenous.
While controlling for occupation will bias the wage gap, it can still be instructive for seeing the effect of anti-discrimination laws on the wage and labor supply gap. We run these regressions with occupation fixed effects using four-digit occupation codes in the ACS to see how the pay gap/premium changes after anti-discrimination laws are passed. Using occupation-fixed effects will examine the effect of anti-discrimination laws within an occupation. One potential mechanism for anti-discrimination laws is that gay workers sort into more in-demand occupations, lowering their relative labor supply and wage gap, while lesbian workers do the opposite. However, we show that even when controlling for occupation there is a significant change in relative labor supply for men and women in same-sex partnerships and a significant change in relative pay for women in same-sex partnerships, indicating there is a significant effect of these laws even within an occupation. We present the extensive labor supply and pay regressions with occupation fixed effects in Table 6 and 7.
In these tables, there is a smaller pay gap for gay men and a smaller pay premium for lesbian women when controlling for occupation or industry. However, qualitatively the results are similar to the main results in Table 4 and Table 5. Gay men see a significant reduction in the labor supply gap, and lesbian women see a significant reduction in their labor supply and annual wage premium. These findings suggest that anti-discrimination laws have a significant effect even within occupations. The effect does not come from workers shifting into occupations with less discrimination, but even when conditional on occupation choice, there is a significant reduction in differences in labor supply across sexual orientations for men and women.
4.3 Impacts Over Time
We use Equation (1) with the inclusion of occupation-fixed effects and conduct event study regressions via Borusyak et al. (2021). This approach allows us to analyze how the impact estimates evolve. First, we present the event study regressions for male and female extensive labor supply outcomes in Figure 3. Panel (a) illustrates the effects on labor force participation rates, whereas Panel (b) shows the effects on employment levels. The plots reveal an initial jump in both employment and labor force participation rates among gay men, a trend that appears to wane over subsequent years. This pattern hints at an early beneficial impact that seems to moderate with time. Conversely, the trajectory for women follows a different pattern, indicating initial adverse effects on employment and labor force participation rates. The pre-trends look fairly parallel, which, while not directly confirming the parallel trends assumption for the post-period, lends it greater plausibility. These visualizations indicate a temporary increase in labor force participation and employment rates for gay men following the enactment of ENDAs, contrasting with a negative impact on the same metrics for lesbian women.
We present the event study regressions for male and female pay penalty/premium outcomes in Figure 4. The figures suggest a delayed response in wage parity gains for gay employees compared to their heterosexual counterparts, with a notable but transient increase in relative gay wages observed a year post-ENDA implementation. This delayed effect may come from implementation lags and the labor market outcomes captured in the ACS, which reflect earnings over a 23-month reference period. Conversely, there is a marked decline in relative wages for lesbian women following the enactment of the law, becoming significant one year after its introduction.
5 Robustness Checks
5.1 Endogenous Adoption of Anti-Discrimination Laws
One concern in this analysis is the endogenous adoption of anti-discrimination laws. Anti-discrimination laws are not randomly distributed. Locally, anti-discrimination laws are concentrated in larger cities, and state laws are to be concentrated in more liberal states that presumably are more accepting of and more favorable to LGB workers. The main concern is that there is an unobservable factor like general sentiment toward LGB workers that affects both the passage of laws as well as the labor market outcomes for LGB workers. In our main regressions, we alleviate this concern by using county-year-sex fixed effects.
Nonetheless, we investigate the endogenous adoption of anti-discrimination laws at the state level. Specifically, we use state-level polling information on support for same-sex marriage as a proxy for general sentiment toward LGB workers. It is possible that controlling for state-level polling is not the best way to capture sentiment toward LGB workers since it is possible to discriminate against people based on sexual orientation and still support their right to marry. However, it seems plausible that the changes in state-level support for same-sex marriage are highly correlated with changes in sentiment toward LGB workers such that it will suffice for a suitable proxy. To better get at the question of endogenous adoption of laws, we create an event-study plot showing how state laws change support for same-sex marriages. Specifically, we estimate this equation:
$\text{Support}_{st}=\sum_{j=-5,j\neq-1}^{j=5}\psi_{j}\mathbbm{1}(\text{Laws}_{st}=j)+\phi_{s}+\delta_{t}+\epsilon_{ts}$ (2)
[…]
The event study supports the idea that anti-discrimination laws at the state level significantly increase the support for same-sex marriage at the state level. The trends before the legislation exhibit similar trends, with a noticeable jump in the fraction of individuals favoring same-sex marriage. This pattern suggests that states, irrespective of their legal stance on anti-discrimination, shared similar trajectories in support for same-sex marriage before the enactment of such laws, and only began to diverge following the implementation of the legislation. It might be argued that the passage of these laws was a consequence of shifting attitudes towards same-sex marriage at the time of enactment. Still, if the adoption of these laws were endogenous, one would expect to see a steady climb in support for same-sex marriage leading up to the law’s passage, with the legislation itself having minimal impact on this trend. However, the distinct and significant jump in support concurrent with the introduction of anti-discrimination laws lends more weight to the argument that it is the legislative changes that are driving the increase in support for same-sex marriage and, likely, increasing the overall sentiment towards LGB individuals, rather than the other way around.
The jump seen in state-level sentiment toward LGB workers following state anti-discrimination law passages could suggest that there is a corresponding jump in sentiment at the local level that is unobservable and potentially acting as a mechanism, affecting the labor supply and pay for LGB workers.
5.2 Sorting and Increased Reporting
One finding from Klawitter and Flatt (1998) suggests that LGB workers sort to areas with anti-discrimination laws. If LGB workers were sorting to areas that recently passed anti-discrimination laws, then it could violate the identifying assumption. It would be concerning if high-wage LGB workers migrated from unprotected areas to areas that recently passed anti-discrimination laws since it would lower the average wage in untreated counties and increase the average wage in treated counties.
Another potential concern would be increased reporting for same-sex partnerships. Same-sex partnerships, while less stigmatized in 2005-2019 than in previous years, were still heavily stigmatized. Anti-discrimination laws could signal to LGB people that their community is more accepting, and people may be more likely to declare that they are in a same-sex partnership. This change in reporting could affect the composition of our sample, which in turn could violate the identifying assumption.
To address this concern, we run the regression below to see if there is an increase in same-sex partnerships in a given county after an anti-discrimination law is passed:
$\text{SSP}_{ict}=\beta_{0}\text{Laws}_{ct}+\beta_{1}X_{i}+\beta_{2}\gamma_{st}+\mu_{c}+\delta_{t}+\epsilon_{itcs}$ (3)
In this regression, $\beta_{0}$ is the coefficient of interest, and it will give the effect of anti-discrimination laws passing on the number of same-sex partnerships in a given county whether that is from in-migration or increased reporting of same-sex partnerships. We present these results in Table 8.
We find no significant effect of anti-discrimination laws on the number of same-sex partnerships at the county level, suggesting there is minimal sorting or change in reporting following the passage of anti-discrimination laws.
6 Discussion
6.1 Differences in Response by Sex
The results paint an interesting picture of the effect of anti-discrimination laws on the workforce, in that they have differing effects on lesbian and gay workers. Lesbian workers see a reduction in the labor supply and pay premium while gay workers see a reduction in their labor supply and pay gap. The anti-discrimination laws appear to push the labor market outcome to a greater level of equality across sexual orientation, where lesbian women move toward the labor market outcomes of straight women and gay men move toward the labor market outcomes of straight men.
One potential explanation for this convergence of labor market outcomes is that employers stop differentiating between workers based on sexual orientation after anti-discrimination laws are passed and that results in greater equality. However, much of the labor market differences arise from differences in human capital accumulation and decisions that come before many of the anti-discrimination laws such as occupation/industry choice and desired number of children (Black et al., 2007).
Another explanation for different responses to anti-discrimination laws for same-sex couples lies in the canonical Becker (1981) theory of the family and division of labor, which is typically used to explain differences in LGB labor market outcomes from their straight counterparts (Black et al., 2007). In a male-female household with a “traditional” division of labor, the man specializes in market production, selling his labor for money to buy market goods, and the woman specializes in household production, using her labor to rear children, cook, and take care of the home. However, LGB people are less likely to end up in a traditional Beckerian household. This expectation of future household composition will induce gay men to be less specialized in market production and lesbian women to be more specialized in market production than their straight counterparts. This theory of household specialization is typically used to explain why gay men make less than straight men and lesbian women make more than straight women (Black et al., 2007).
This household model can also be used to think about responses to anti-discrimination laws. One potential explanation for lesbian women becoming more likely to leave the workforce and earn less is that a lesbian household may be more likely to shift to a more traditional Beckerian household where one parent specializes in household production and the other shifts to a more market orientation. If anti-discrimination laws decreased the risk of the primary earner getting fired for their sexual orientation, it could induce the secondary earner to switch to more part-time work or exit the workforce to specialize in household production or focus on child-rearing. Alternatively, both women in a female same-sex partnership could switch to reducing hours to spend more time in household production as a result of increased job security.
Households consisting of lesbian women typically have more children than households consisting of gay men. Given lesbian women have more children than gay men, it would be unsurprising to see lesbian households switch more to a specialization of their household labor than gay men would, which would help explain why anti-discrimination laws have different effects on lesbian women and gay men.
To test the different responses by men and women in same-sex partnerships to anti-discrimination laws and the implications of Becker’s household model, we implement an alternative difference-in-difference model. Instead of comparing those in same-sex partnerships to those in different-sex partnerships, we compare men and women in same-sex partnerships to examine their differential responses at the household level. Do women in same-sex partnerships change their household division of labor relative to men in same-sex partnerships following anti-discrimination laws? To answer this question, we collapse the data to the household level and examine households instead of individuals.
Specifically, we estimate this model:
$y_{jct}=\alpha_{0}Law_{ct}+\alpha_{1}FemSSP_{j}+\alpha_{2}Law_{ct}*FemSSP_{j}+\mu_{c}+\delta_{t}+\epsilon_{jct}$
The notation is the same as before with two main differences. Our interaction term uses an indicator for if household $j$ is a female same-sex partnership, denoted by $FemSSP_{j}$. $y_{jct}$ is a variable for outcomes at the household level. $\alpha_{2}$ is our coefficient of interest and will show how lesbian households differ in their response to anti-discrimination laws compared to gay households. The first outcome examined is an indicator for if the household is a one-earner family. We also examine the difference in absolute value between the two partners in terms of weekly hours worked. These labor supply measures inform how intra-household labor is divided between partners along the extensive and intensive margin of labor supply. Finally, we examine how the anti-discrimination laws may affect child rearing by examining if the households have any children and the number of children they have.
We present the results in Table 9. Anti-discrimination laws do not affect the likelihood of same-sex partnerships becoming a one-earner household, but there is an effect on the intensive margin. There is a significant effect of anti-discrimination laws in the intra-household difference in hours worked for gay and lesbian households. An increase in the difference of hours worked within the household could suggest a greater specialization in market production for one person and a greater specialization in household production and working fewer hours for the other person in the partnership. Lesbian households do see a significant increase in the likelihood of having any child and the number of children they have relative to gay households.
This evidence tied with Becker’s theoretical work suggests that men and women in same-sex partnerships are responding differently to these anti-discrimination laws. Lesbian households begin to have more children than gay households. It may appear in the main results that anti-discrimination laws hurt lesbian women because their labor supply and pay premium over straight women shrinks. However, their response in the labor market could be driven by a greater desire for children and subsequent changes in household labor allocation due to an increase in job security. The corresponding changes in the labor market could represent an increase in welfare for lesbian households if they value their new children more than their lost hours worked and subsequently lost wages. Consistent with the theory and empirical findings, lesbian households could become more secure with one woman being the primary earner following anti-discrimination laws.
6.2 Mechanisms
The results of this research indicate that sexual orientation anti-discrimination laws could significantly impact the labor market gaps for gay/bisexual men as well as the labor market premiums for lesbian/bisexual women. It is possible that these laws are binding and eliminate significant discrimination in the labor market. However, it is relatively easy for an employer to create a reason to fire an employee rather than firing someone specifically for being gay. It is also difficult to accurately determine someone’s sexual orientation by looking at them, unlike race and sex. Although it may be easier to determine someone’s sexual orientation conditional on them being in a same-sex relationship. Sexual orientation anti-discrimination laws may not be binding in a meaningful way that eliminates discrimination.
One mechanism that the laws could be affecting labor market outcomes is through increased positive sentiment toward LGB workers. We show in Figure 5 that following state anti-discrimination laws there is a significant and persistent increase in the percentage of people in that state that support same-sex marriage. Our regression results already condition on the change in the county-level support for same-sex marriage via fixed effects. In Appendix A, we show that local laws matter significantly, so it seems plausible that local law changes could significantly change local sentiment and impact the labor supply and pay of LGB workers. Changing public sentiment in conjunction with greater protection for LGB workers seems more plausible as the mechanism than simply greater protection in the workplace.
6.3 Threats to External Validity
One limitation of this research is that we are unable to view all LGB workers. The analysis is specifically conducted on individuals in same-sex partnerships and different-sex partnerships. The data does not allow for the identification of single LGB workers, whose outcomes may be significantly different than LGB workers in partnerships. Another pitfall to using partnerships to infer sexual orientation is that one could incorrectly infer someone’s sexual orientation from a partnership. Bisexual people are a larger group than lesbian or gay people, and this specification could erroneously assign bisexual people to be “straight” through being in a different-sex partnership (Gates, 2011). If bisexual workers appear more similar to their lesbian and gay counterparts then the results of quantifying the labor supply and pay gap/premium would likely be biased toward zero and the change in those labor market gaps/premiums would also be biased toward zero.
Another potential concern in using partnerships to infer sexual orientation is that people in a same-sex partnership may be the group that is most at risk of discrimination. It seems likely that single LGB workers can more plausibly stay in the closet with their co-workers compared to their counterparts in a same-sex partnership. This analysis may be capturing the effect of anti-discrimination laws on the group that is most likely to be affected. These results may have a larger effect size than what one would find in examining the generalized LGB population.
These biases are for the broader LGB population and external validity. Notably, these potential biases do not affect internal validity. Translating the effect of anti-discrimination laws for people in same-sex partnerships to all other LGB workers is not immediately obvious. It is dependent on several factors that are unknowable in this analysis such as the wages and labor supply for single LGB workers and bisexual workers in different-sex partnerships and how discrimination affects LGB people for those in same-sex partnerships.
7 Conclusion
This analysis is the first quasi-experimental research examining how both local and state anti-discrimination laws on sexual orientation affect the labor supply and pay gap/premium between LGB and straight workers. We construct a novel panel dataset on local anti-discrimination laws to capture the granular nature of anti-discrimination laws. We find that anti-discrimination laws appear to decrease sexual orientation inequality in the labor market. Specifically, we find that anti-discrimination laws significantly reduce the gay labor force participation gap by 2.1 p.p., the employment gap by 1.5 p.p., and increase hourly wages by 6.2%. Anti-discrimination laws also reduce the lesbian labor force participation premium by 1.7 p.p., and the employment premium by 2.3 p.p. We also show that one potential mechanism that anti-discrimination laws work through is increasing positive sentiments toward LGB Americans as measured by the support of same-sex marriage. Finally, we explain the differential response to these laws between gay and lesbian households through Becker’s household specialization model and support the explanation with empirical evidence.
Before 2020, it was legal in some states to fire someone solely based on their sexual orientation, and states have been less inclined to pass anti-discrimination laws recently. Arkansas and Tennessee even passed anti-anti-discrimination laws preventing cities from enacting anti-discrimination laws. From 2009 to 2019, more states have barred cities from protecting their residents than the number of states extending protection. Our research gives a comprehensive look at a policy that Congress is considering in “The Equality Act” to extend federal protection to sexual orientation and gender identity. Our research gives the most comprehensive look at policy and can meaningfully inform the policy discussion around anti-discrimination laws. Overall, this research broadly informs the policy decision of extending employment protections to the LGB community by the Supreme Court of the United States.
Appendix A TWFE Estimates
In the original version of our paper, we employed a traditional difference-in-differences (DiD) setup, incorporating a range of controls to assess the impact of anti-discrimination legislation on labor market outcomes for LGB workers. This conventional approach served as the foundational analysis for our initial findings. To ensure transparency and thoroughness in our research, we are including these initial estimates in this section. Furthermore, in response to recent advancements in DiD methodologies, we use Borusyak et al. [2021] to estimate the impact of anti-discrimination laws. Our analysis reveals that the application of these modern DiD techniques does not substantially alter the point estimates we initially obtained. This consistency across methodological frameworks underscores the robustness of our findings, reinforcing the reliability of our conclusions regarding the effects of anti-discrimination laws on LGB labor market outcomes.
In the TWFE setup, we control for sentiment toward LGB workers in a given state at a given year by including polling information on support for same-sex marriages. Unfortunately, we cannot observe polling at a local level, as such the polling measure is an incomplete proxy for sentiment toward LGB workers. We collected every poll that Pew Research Center published from 2005-2016 to pull each poll regarding support for same-sex marriage. The 28 polls asking about same-sex marriage were aggregated to get the percentage of people that supported same-sex marriage by state by year. The polling information contains data on every state for every year, except for Alaska and Hawaii, which are missing polling information for 2005-2008.
A.1 Identification Strategy & Estimation
The differential roll-out of anti-discrimination laws by state and locality over time lends itself to a difference-in-differences framework with the main outcomes of interest as the change in labor supply and pay between two groups, LGB workers and heterosexual workers. This strategy uses the variation presented in Figure 1, which shows how the laws changed over time by state and county.
Formally, we estimate this equation on labor supply:
$LS_{ict}=\alpha_{0}Law_{ct}+\alpha_{1}SSP_{i}+\alpha_{2}Law_{ct}\cdot SSP_{i}+\alpha_{3}X_{i}+\alpha_{4}\gamma_{st}+\mu_{c}+\delta_{t}+\epsilon_{itcs}$ (4)
Where $LS_{ict}$ is the labor force participation or employment status for person $i$ in county $c$ in year $t$. $Law_{ct}$ is an indicator for if county $c$ has an employment anti-discrimination law on sexual orientation or is covered by a state-wide anti-discrimination law in year $t$, and $SSP_{it}$ is an indicator for if person $i$ is in a same-sex partnership. $X_{i}$ is a vector of person-specific covariates including race, children, along with age and education level - each represented up to a quadratic term to capture potential nonlinear effects. $\gamma_{st}$ is a vector for state $s$ covariates on the legality of same-sex marriage in year $t$ as well as polling for the support of same-sex marriage. It includes the percentage of people that “strongly support,” “support,” “oppose,” and “strongly oppose” same-sex marriage in a given state in a given year. Finally, $\mu_{c}$ gives county fixed effects, $\delta_{t}$ gives year fixed effects, and $\epsilon_{itcs}$ is the error term [^15] Since the ACS only identifies metro counties, the county fixed effect is close to a pseudo-county fixed effect. The rural counties are all compared together with a true county fixed effect for metro area counties.. Our sample is limited to people aged 25 to 65, and we estimate this model separately by sex using the ACS person weights. We cluster the standard errors at the county level, as that is the level of the treatment in this specification. [^16] Treatment status is computed using the earliest year of ENDAs at the county or state level, with individuals remaining classified as “treated” thereafter throughout the dataset. Given the ACS’s 23-month reference period for earnings, as discussed by Hansen et al. [2020], there’s a potential for bias in treatment assignment. This extended timeframe might not precisely align with the exact implementation dates of ENDAs, potentially biasing the immediate effects of these policies towards zero. The coefficients of interest are $\alpha_{2}$, which will give the effect of the law on the labor supply gap/premium, and $\alpha_{1}$, which gives the labor supply gap/premium in the absence of anti-discrimination law. Our specification assumes that the effect of the laws is the same across states and localities, and we will be estimating an average effect. We estimate a similar equation for wages:
$\ln(Y_{ict}+1)=\beta_{0}\text{Law}_{ct}+\beta_{1}\text{SSP}_{i}+\beta_{2}\text{Law}_{ct}\cdot\text{SSP}_{i}+\beta_{3}X_{i}+\beta_{4}\gamma_{st}+\mu_{c}+\delta_{t}+\epsilon_{itcs}$ (5)
Where $Y_{ict}$ is the real value, in 1999 dollars, of person $i$’s wages in county $c$ year $t$. The rest of the notation is the same as above. The coefficients of interest are $\beta_{2}$, which will give the effect of the law on the pay gap/premium, and $\beta_{1}$, which gives the pay gap/premium. [^17] The $\ln(Y_{ict}+1)$ specification is adopted to include both employed and unemployed individuals in the analysis, recognizing employment as endogenous to anti-discrimination laws. This approach ensures a comprehensive evaluation of the law’s impact on the entire income spectrum, capturing nuances missed by focusing solely on employed workers. We report the coefficients estimated via Borusyak et al. [2021] in the main text, and standard TWFE in the Appendix.
A.2 TWFE Results
In Table A1, we present the regression on the extensive margin of labor supply with Panel A giving the effect of anti-discrimination laws on gay men’s labor supply and Panel B giving the effect for lesbian women. We present the full specification in columns (1) and (2) and include a specification without accounting for local anti-discrimination laws in columns (3) and (4) to show how failing to account for local laws would lead to incorrect inference.
Consistent with the previous literature, we find that lesbian women have a labor supply premium and gay men have a labor supply penalty on the extensive margin. Lesbian women are 7.3% more likely to be employed and gay men are 7.9% less likely to be employed than their straight peers. Anti-discrimination laws have a significant effect in reducing the labor force participation and employment gap for gay men. The labor force participation rate gap is reduced by 1.3 p.p. (18%), and the employment gap is reduced by 1.4 p.p. (18%). Conversely, the labor force participation and employment premium that lesbian women have over their straight counterparts is reduced by 1.3 p.p. (16%) and 1 p.p. (14%), respectively. As seen in columns (3) and (4) of Panel A, estimating the effect of anti-discrimination laws only using state laws will lead one to erroneously conclude that these laws have no significant effect on the gay labor force and employment gap. However, the effect on the extensive margin of labor supply for lesbian women looks similar when estimating the effect of just state anti-discrimination laws.
In Table A2, we present the regression results on the intensive margin of labor supply as measured by weekly hours worked and weeks worked. Similar to Table 4, we include the results when only estimating the effects of the state laws in columns (3) and (4) to show the contribution of including local laws in estimating the effect of anti-discrimination laws.
We replicate past findings where gay men supply significantly less labor and lesbian women supply significantly more labor on the intensive margin. As with Table A1, we find contrasting effects of anti-discrimination laws on male and female outcomes. We find that the intensive labor supply gap for gay men as measured by weekly hours worked and weeks worked is significantly reduced, and the intensive labor supply premium for lesbian women is significantly reduced as well. The effect of anti-discrimination laws on weekly hours worked becomes insignificant for gay men and lesbian women when only estimating off of state-wide anti-discrimination laws.
In Table A3, we present the regression results on the wage gap as measured by hourly wages and annual wage earnings. These variables received a log-transformation, to handle zero values so the changes can be interpreted in percentage terms. We run the regressions separately by sex and present the regressions showing how the sexual orientation anti-discrimination laws impact both employed and unemployed people in columns (1) and (2) and limit the regressions to only those who are employed in columns (3) and (4). Our preferred specification is to look at both employed and unemployed people in the sample as employment is endogenous. Conditioning on employed workers would create an endogeneity problem since labor force participation and employment are significantly impacted by anti-discrimination laws. Given Table A1, the sample of employed workers could be changing as a result of the law changes, which would lead to a change in the sample composition. We present the regressions limited to employed workers to replicate past work on the wage gap/premium where it is common to only examine employed workers. We include results by only using state anti-discrimination laws, excluding local laws. We present those results in columns (5) - (8).
Gay men see a significant decline in the hourly wage gap by 2.8 p.p. (11%) at the 90% significance level and see an insignificant decline in the annual earnings gap by 5.8 p.p. On the other hand, lesbian women see a significant change in their annual earnings premium by 13 p.p. (16%) and an insignificant reduction in their hourly wage premium. Columns (3) and (4) replicate the past literature where employed gay men make about 8% to 11% less than their straight counterparts and employed lesbian women make about 5% to 15% more than their straight counterparts.
The pay gap/premium is reduced quite significantly once the sample is limited to employed workers, which is unsurprising. As shown in Table A1 and Table A2, there are significant differences in labor supply for LGB and straight workers. Gay men and straight women are far more likely to be unemployed, out of the labor force, and work fewer hours and weeks, so including their zero wages in the regression will increase the gap/premium.
There is a significant decrease in the hourly wage gap now and an insignificant decrease in the annual wage gap. There is also a significant decrease in the annual earnings premium for lesbian women but no change in their hourly wage premium. This result is consistent with the labor supply results in Table A1 and A2 where their intensive and extensive margin of labor supply premium are declining, so the annual earnings decline, but hourly wage does not.
A.3 Occupation Fixed Effects
We show that even when controlling for occupation there is a significant change in relative labor supply for men and women in same-sex partnerships and a significant change in relative pay for women in same-sex partnerships, indicating there is a significant effect of these laws even within occupation. We present the extensive labor supply and pay regressions with occupation fixed effects in Table A4 and A5.
In these tables, there is a smaller pay gap for gay men and a smaller pay premium for lesbian women when controlling for occupation or industry. However, qualitatively the results seem to be similar to the main results in Table A1 and Table A3 . Gay men see a significant reduction in the labor supply gap, and lesbian women see a significant reduction in their labor supply and annual wage premium. There is no longer a significant effect on the wage gap for gay men. These findings suggest that anti-discrimination laws have a significant effect even within occupations. The effect does not come from workers shifting into occupations with less discrimination, but even when conditional on occupation choice, there is a significant reduction in differences in labor supply across sexual orientations for men and women.
Appendix Tables
Table A1: Effect of Anti-Discrimination Laws: Extensive Margin of Labor Supply
Panel A: Men
(1) (2) (3) (4)
Labor Force Employed Labor Force Employed
Laws*SSP 0.013*** 0.014** 0.007 0.005
(0.005) (0.005) (0.005) (0.006)
SSP -0.074*** -0.079*** -0.068*** -0.073***
(0.003) (0.004) (0.003) (0.004)
State Laws Only X X
Observations 6,287,441 6,287,441 6,287,441 6,287,441
R-squared 0.135 0.114 0.135 0.114
Panel B: Women
(1) (2) (3) (4)
Labor Force Employed Labor Force Employed
Laws*SSP -0.013** -$0$016*** -0.013**
(0.006) (0.006) (0.006) (0.006)
SSP 0.081*** 0.073*** 0.080*** 0.072***
(0.004) (0.005) (0.004) (0.004)
State Laws Only X X
Observations 6,569,373 6,569,373 6,569,373 6,569,373
R-squared 0.091 0.088 0.091 0.088
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex. All of the outcome are binary taking a value of 0 or 1. The first row of coefficients shows the effect of anti-discrimination on the labor supply gap or premium, and the second row of coefficients gives the labor supply gap or premium. Columns (3) and (4) present results when estimating only using state-wide anti-discrimination laws for comparison. Standard errors are clustered at the county level. $*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table A2: Effect of Anti-Discrimination Laws: Intensive Margin of Labor Supply
Panel A: Men
(1) (2) (3) (4)
Weekly Hours Weeks Worked Weekly Hours Weeks Worked
Laws*SSP 0.273* 0.089 0.180 -0.0744
(0.160) (0.080) (0.164) (0.0820)
SSP -2.196*** -0.575*** -2.106*** -0.477***
(0.114) (0.056) (0.0923) (0.0480)
State Laws Only X X
Observations 5,244,258 5,244,258 5,244,258 5,244,258
R-squared 0.034 0.011 0.034 0.011
Panel B: Women
(1) (2) (3) (4)
Weekly Hours Weeks Worked Weekly Hours Weeks Worked
Laws*SSP -0.305** -0.184** -0.213 -0.171**
(0.136) (0.084) (0.134) (0.0790)
SSP 2.792*** 0.604*** 2.706*** 0.574***
(0.105) (0.066) (0.0924) (0.0577)
State Laws Only X X
Observations 4,366,603 4,366,603 4,366,603 4,366,603
R-squared 0.038 0.015 0.038 0.015
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex. All of the outcome are binary taking a value of 0 or 1. The first row of coefficients shows the effect of anti-discrimination on the labor supply gap or premium, and the second row of coefficients gives the labor supply gap or premium. Columns (3) and (4) present results when estimating only using state-wide anti-discrimination laws for comparison. Standard errors are clustered at the county level. $*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table A3: Effect of Anti-Discrimination Laws: Wages and Earnings
Panel A: Men
(1) (2) (3) (4) (5) (6) (7) (8)
Hourly Wage Annual Earnings Hourly Wage Annual Earnings Hourly Wage Annual Earnings Hourly Wage Annual Earnings
Laws*SSP 0.028* 0.059 0.011 -0.041 0.0139 -0.00404 0.00639 -0.0581
(0.017) (0.050) (0.015) (0.035) (0.0184) (0.0525) (0.0168) (0.0373)
SSP -0.244*** -0.757*** -0.082*** -0.114*** -0.233*** -0.716*** -$0**$112***
(0.012) (0.038) (0.009) (0.023) (0.0108) (0.0354) (0.00745) (0.0184)
Employed Only X X X X
State Laws Only X X X X
Observations 6,287,441 6,287,441 5,244,258 5,244,258 5,244,258 6,287,441 5,244,258 6,287,441
R-squared 0.157 0.121 0.127 0.041 0.127 0.157 0.041 0.121
Panel B: Women
(1) (2) (3) (4) (5) (6) (7) (8)
Hourly Wage Annual Earnings Hourly Wage Annual Earnings Hourly Wage Annual Earnings Hourly Wage Annual Earnings
Laws*SSP -0.019 -0.129** -0.024** -0.062** -0.0203 -0.153** -0.0144 -0.0453
(0.020) (0.066) (0.011) (0.029) (0.0202) (0.0668) (0.0112) (0.0297)
SSP 0.220*** 0.815*** 0.053*** 0.149*** 0.218*** 0.809*** 0.0448*** 0.132***
(0.014) (0.049) (0.007) (0.021) (0.0116) (0.0409) (0.00635) (0.0180)
Employed Only X X X X
State Laws Only X X X X
Observations 6,569,373 6,569,373 4,366,603 4,366,603 6,569,373 6,569,373 4,366,603 4,366,603
R-squared 0.136 0.106 0.123 0.036 0.136 0.106 0.123 0.036
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex. Columns (3) and (4) are limited to those who are employed. All of the are in terms of 1999 dollars and received the log(x+1) transformation to have the interpretation of the coefficient be in terms of percent. The first row of coefficients show the effect of anti-discrimination laws on the pay gap or premium, and the second row of coefficients give the pay gap or premium. Columns (5) - (8) present results when estimating only using state-wide anti-discrimination laws for comparison. Standard errors are clustered at the county level. $*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table A4: Anti-Discrimination Laws on Labor Supply with Occupation FEs
Men Women
(1) (2) (1) (2)
Labor Force Employed Labor Force Employed
Laws*SSP 0.009*** 0.009** -0.009*** -0.009***
(0.003) (0.004) (0.003) (0.004)
SSP -0.037*** -0.045*** 0.039*** 0.036***
(0.002) (0.003) (0.002) (0.003)
Observations 6,287,441 6,287,441 6,569,373 6,569,373
R-squared 0.372 0.461 0.517 0.461
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex and include four-digit occupation fixed effects. All of the outcome are binary taking a value of 0 or 1. The first row of coefficients show the effect of anti-discrimination on the labor supply gap or premium, and the second row of coefficients give the labor supply gap or premium. Standard errors are clustered at the county level. $*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table A5: Anti-Discrimination Laws on Pay with Occupation FEs
Panel A: Men
(1) (2) (3) (4)
Hourly Wage Annual Earnings Hourly Wage Annual Earnings
Laws*SSP 0.013 0.018 0.011 -0.028
(0.013) (0.035) (0.013) (0.033)
SSP -0.113*** -0.351*** -0.043*** -0.051**
(0.008) (0.025) (0.008) (0.022)
Employed only X X
Observations 6,287,441 6,287,441 5,244,258 5,244,258
R-squared 0.372 0.372 0.248 0.169
Panel B: Women
(1) (2) (3) (4)
Hourly Wage Annual Earnings Hourly Wage Annual Earnings
Laws*SSP -0.015 -0.104*** -0.024*** -0.063**
(0.011) (0.034) (0.009) (0.025)
SSP 0.117*** 0.437*** 0.056*** 0.172***
(0.008) (0.028) (0.007) (0.018)
Employed only X X
Observations 6,569,373 6,569,373 4,366,603 4,366,603
R-squared 0.513 0.527 0.303 0.233
  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex and include four-digit occupation fixed effects. All of the are in terms of 1999 dollars and received the log(x+1) transformation to have the interpretation of the coefficient be in terms of percent. The first row of coefficients show the effect of anti-discrimination on the pay gap or premium, and the second row of coefficients give the pay gap or premium. Standard errors are clustered at the county level. $*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table A6: Sorting and Reporting
(1) (2) (3)
All SSP Male SSP Female SSP
Laws 0.000170 0.000633 -0.000304
(0.000489) (0.000572) (0.000573)
Observations 12,872,572 6,295,028 6,577,544
R-squared 0.068 0.072 0.065
  • Notes: Data comes from the 2005-2016 yearly ACS seeing how the number of same-sex partnerships change in a county after the passage of an anti-discrimination law. The first column looks at both men and women with the next two columns separating the sexes. Standard errors are clustered at the county level. $*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
Table A7: Differential Responses between Male and Female Same-Sex Partnerships
(1) (2) (3) (4)
One Earner HH Diff in Hours Worked Any Children Number of Children
Laws*FemSSP -0.00308 0.947*** 0.0344*** 0.0544***
(0.00868) (0.364) (0.00846) (0.0195)
FemSSP 0.00445 -0.729** 0.123*** 0.193***
(0.00717) (0.317) (0.00743) (0.0176)
Observations 73,181 73,181 73,181 73,181
R-squared 0.046 0.043 0.122 0.111
  • Notes: Data comes from the 2005-2016 yearly ACS comparing women in same-sex partnerships to men in same-sex partnerships. The first column examines if the household has only one-earner. The second column tests the difference in absolute terms of hours worked between the two partners. The third and fourth columns examines how households differ with having children. The first row of coefficients show the differential effect of anti-discrimination between lesbian and gay households. Standard errors are clustered at the county level. $*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.

### v2

@@
 Scott Delhommer University of Texas at Austin; sdelhommer@utexas.edu    Domonkos F. Vamossy University of Pittsburgh, d.vamossy@pitt.edu
-(April 4, 2024)
+(April 1, 2025)
 This paper presents the first quasi-experimental research examining the effect of both local and state anti-discrimination laws on sexual orientation on the labor supply and wages of lesbian, gay, and bisexual (LGB) workers.
-To do so, we use the American Community Survey data on household composition to infer sexual orientation and combine this with a unique panel dataset on local anti-discrimination laws.
-Using variation in law implementation across localities over time and between same-sex and different-sex couples, we find that anti-discrimination laws significantly reduce gaps in labor force participation rate, employment, and the wage gap for gay men relative to straight men.
-These laws also significantly reduce the labor force participation rate, employment, and wage premium for lesbian women relative to straight women.
-One explanation for the reduced labor supply and wage premium is that lesbian couples begin to have more children in response to the laws.
+To do so, we use the American Community Survey data on household composition to infer sexual orientation and combine this with a unique panel dataset on state and local anti-discrimination laws.
+Leveraging variation in law implementation across localities over time and between same-sex and different-sex couples, we find that anti-discrimination laws not only significantly narrow gaps in labor force participation and employment for male same-sex couples relative to men in different-sex couples, but also boost their percentile rank in the wage distribution.
+Our analysis reveals mostly null effects for female same-sex couples; however, in metropolitan areas these laws significantly reduce their employment compared to women in different-sex couples.
+One explanation for the reduced labor supply is that female same-sex couples begin to have more children in response to the laws.
 Finally, we present evidence that state anti-discrimination laws significantly and persistently increased support for same-sex marriage.
@@
 The focus of this paper is the effect of local and state sexual orientation anti-discrimination legislation on labor market differentials for lesbian, gay, and bisexual (LGB) workers.
+[^2] For a review of the impacts of legal access to same-sex marriage, see Badgett et al.
+(2024).
+For a broader review of the economics related to sexual orientation and gender identity, see Badgett et al.
+(2023).
 We exploit the differential roll-out of state and local laws from 2005-2019 in a modern triple difference-in-differences framework to analyze how these anti-discrimination laws impact the wages and labor supply of LGB workers.
-The economics literature on LGB workers consistently finds that gay/bisexual men have a pay and labor supply penalty and lesbian/bisexual women have a pay and labor supply premium over their heterosexual counterparts222(Badgett (1995); Klawitter and Flatt (1998); Allegretto and Arthur (2001); Black et al.
+The economics literature on LGB workers consistently finds that gay/bisexual men have a pay and labor supply penalty and lesbian/bisexual women have a pay and labor supply premium over their heterosexual counterparts333(Badgett (1995); Klawitter (2011); Klawitter and Flatt (1998); Allegretto and Arthur (2001); Black et al.
 (2003); Carpenter (2005); Black et al.
 (2007); Antecol et al.
-(2008); Klawitter (2015); Jepsen and Jepsen (2017); Carpenter and Eppink (2017)).
+(2008); Klawitter (2015); Jepsen and Jepsen (2017); Carpenter and Eppink (2017)), Tilcsik et al.
+(2015).
 The differences in pay are attributed to a host of factors including discrimination and intra-household labor allocation and specialization.
@@
 [^5] The Williams Institute found that sexual orientation discrimination occurs at a similar rate to sex-based discrimination using state-level complaint data following state anti-discrimination laws (Ramos et al., 2008).
-Martell (2013) study demonstrated that such laws have significantly reduced wage disparities for gay men by 20%.555 Since Martell (2013) and Klawitter and Flatt (1998) there has been a booming literature on difference-in-differences.
+Martell (2013) study demonstrated that such laws have significantly reduced wage disparities for gay men by 20%.666 Since Martell (2013) and Klawitter and Flatt (1998) there has been a booming literature on difference-in-differences.
 Accordingly, unlike Martell (2013) and Klawitter and Flatt (1998), we employ modern difference-in-difference estimators to alleviate biases stemming from staggered policy design.
@@
 Furthermore, this study also estimates the effects of such laws on women, a demographic previously under-explored in this context.
-Our study expands the scope of Klawitter and Flatt (1998) by leveraging a larger dataset, utilizing American Community Survey (ACS) data spanning from 2004 to 2019, in contrast to Klawitter and Flatt (1998) which relied on the 1990 U.S.
+Our study expands the scope of Klawitter and Flatt (1998) by leveraging a larger dataset, utilizing American Community Survey (ACS) data spanning from 2005 to 2019, in contrast to Klawitter and Flatt (1998) which relied on the 1990 U.S.
 census data encompassing approximately 13,000 married couples, 14,500 different-sex unmarried couples, and 6,800 same-sex couples.
@@
 Previous studies have not fully addressed potential biases arising from variations in the timing of these laws’ enactment.
-To bridge this gap, our research employs the methodology proposed by Borusyak et al.
-(2021), comparing it with the conventional triple difference-in-differences estimator, finding both approaches yield similar results.
 This paper is the first to examine both local and state-level sexual orientation anti-discrimination laws in a quasi-experimental design.
@@
 We find a significant reduction in differences between LGB workers and heterosexual workers across labor supply and wage measures, due to anti-discrimination laws.
-Anti-discrimination laws significantly reduce the gap in labor force participation and employment of gay men by 2.1 p.p.
-and 1.5 p.p., respectively.
-The laws also significantly increase hourly wages for gay men by 6.2%.
-The results differ for women, with the laws significantly reducing their labor force participation and employment premium over straight women by 1.7 p.p., and 2.3 p.p., respectively.
+Anti-discrimination laws significantly reduce the gap in labor force participation and employment of men in male same-sex couples by 3.8 percentage points (78%) and 3.7 percentage points (77%), respectively.
+The laws improve men in male same-sex couples relative position in the wage distribution, with their percentile rank in the hourly wage and annual earnings distributions improving on average by approximately 2.7 and 3 percentiles, respectively.
+By contrast, the wage distribution for women in same-sex partnerships remains largely unaffected by these laws.
 Using an event study plot, we show that the outcome trends are parallel before the implementation of the anti-discrimination laws suggesting the workers in treatment and control regions are reasonable comparisons.
-[^7] In Appendix A, we replicate the findings of past research that finds gay/bisexual men have lower labor force participation and employment rates and make 8-11% less than their employed straight counterparts using hourly wages and annual earnings.
-We also replicate findings that lesbian/bisexual women have a higher labor force participation and employment rate and earn 5-15% more than their employed straight counterparts.
-We explore theories for the differing effects of anti-discrimination laws on gay men and lesbian women in the Discussion section, using the Becker (1981) model of household specialization.
-In the traditional Becker (1981) model of household specialization, men typically specialize in market production, and women typically specialize in household production, in part due to differences in biology where women birth and care for children, resulting in a one-earner household.
-These differences in household specialization are less pronounced in same-sex partnerships, but they may become more similar for women in same-sex partnerships following the passage of an anti-discrimination law if it gives greater protection to the higher wage earner.
-We show empirically that the difference in hours worked between partners within a lesbian household goes up relative to gay households following anti-discrimination laws, suggesting that lesbian households could become more specialized following anti-discrimination laws with one woman working more hours and the other woman working fewer to focus on household production.
-We also show that lesbian households have significantly more children than gay households after the passage of an anti-discrimination law.
-More children could induce lesbian households to further specialize in the intra-household division of labor, adopting a more traditional household model to help care for additional children.
+[^8] In Appendix B A, we replicate the findings of past research that finds men in male same-sex couples have lower labor force participation and employment rates and make 8-11% less than their employed straight counterparts using hourly wages and annual earnings.
+We also replicate findings that women in female same-sex couples have a higher labor force participation and employment rate and earn 5-15% more than their employed straight counterparts.
+We explore theories for the differing effects of anti-discrimination laws on men in male same-sex couples and women in female same-sex couples in the Discussion section.
+Reduced discrimination improves job security for sexual minorities (Badgett, 2006), which may allow the secondary earner to reduce market participation in favor of increased household production, including child rearing, a pattern consistent with household bargaining models (Chiappori, 1992; Lundberg and Pollak, 1993).
+These differences in household specialization are less pronounced in same-sex partnerships, but they may become more similar for women in female same-sex couples following the passage of an anti-discrimination law if it gives greater protection to the higher wage earner.
+We show empirically that female same-sex households have significantly more children than male same-sex households after the passage of an anti-discrimination law.
 Finally, our last contribution is the use of polling data on the support for same-sex marriage to examine the relationship between anti-discrimination laws and state sentiment toward LGB workers.
@@
 2 Data
-A common issue in analyzing the pay gap/premium between homosexual and heterosexual workers is a lack of high-quality data about sexual orientation, wages, and employment.
+A common issue in analyzing the pay disparities between homosexual and heterosexual workers is a lack of high-quality data about sexual orientation, wages, and employment.
 We follow the literature (Klawitter and Flatt (1998); Gates (2009); Jepsen and Jepsen (2017)) in inferring sexual orientation by looking at household composition.
@@
 [^9] We exclude individuals whose sex or relationship status was allocated following (Gates and Steinberger, 2009)..
-The ACS references a close personal relationship for unmarried partners as well as giving an option for “partner/roommate”, which reduces the likelihood that straight roommates would misclassify as unmarried partners.
-The comparisons that we make are between people in same-sex relationships, defined as being in an unmarried partnership or married with someone of the same sex, and those in different-sex relationships.
+The ACS references a close personal relationship for unmarried partners as well as giving an option for “partner/roommate”, which reduces the likelihood that different-sex roommates would misclassify as unmarried partners.
+The comparisons that we make are between people in same-sex partnerships (SSP), defined as being in an unmarried partnership or married with someone of the same sex, and those in different-sex partnerships (DSP).
 This comparison based on household composition, though standard in the literature, is not equivalent to comparing LGB and heterosexual workers.
 It compares those who are in a same-sex relationship to those who are in a different-sex relationship.
-Notably, this comparison excludes all workers who are not in a cohabitating relationship and classifies bisexual people in different-sex relationships as members of the control group.
-For example, if a person is misclassified as straight instead of bisexual, and bisexual workers have lower pay than their straight counterparts, then the misclassification would bias our pay gap/premium results toward zero.
+Importantly, this comparison excludes all workers who are not in a cohabitating relationship and classifies bisexual individuals in different-sex relationships as members of the control group.
+Because bisexual workers consistently earn less than both heterosexual and gay or lesbian counterparts ((Mize, 2016)), this classification introduces a bias toward underestimating pay disparities.
+Additionally, similar misclassification may occur if gay or lesbian individuals are categorized as heterosexual.
+Thus, the estimated pay gaps presented here likely represent conservative estimates.
 We use the 2005-2019 yearly ACS from (Ruggles et al., 2019).
@@
 The ACS reports wage earnings in annual terms.
-We use the annual figures and convert them to hourly wage figures using variables on average weekly hours worked and weeks worked999“Weeks worked” in the ACS is a categorical variable giving a range of weeks worked.
+We use the annual figures and convert them to hourly wage figures using variables on average weekly hours worked and weeks worked101010“Weeks worked” in the ACS is a categorical variable giving a range of weeks worked.
 We take the median value given in the range to compute weeks worked..
@@
 We completed the dataset for the relevant years using old media reports, correspondence with local officials, and FOIA requests.
-This dataset is the first complete city-level dataset on sexual orientation anti-discrimination laws in the U.S.
+Building on earlier city-level analyses such as (Klawitter, 2011), this dataset provides the first complete city-level dataset covering 15 years of sexual orientation anti-discrimination laws in the U.S.
 We focus on cities reported in the ACS and match those city laws with their corresponding counties to merge with the ACS.
-The ACS only reports county of residence for those in metro areas, so any rural counties are lumped together.
-We make the assumption in our analysis that the impact of these laws is the same across states and localities, estimating an average effect.
+The ACS only reports county of residence for those in metro areas, so any rural counties are lumped together within a state.
 We report the state, city, and county laws and the year they were enacted in Table 1.
@@
 We mitigate this by adopting a saturated triple-difference model, including all two-way fixed effects (county-year, year-SSP, and county-SSP), as proposed by (Burn, 2018).
-The inclusion of county-year fixed effects allows us not to include county controls hence eliminating the need to control for unavailable local levels of support for same-sex marriage.
-We refine the approach of (Burn, 2018) by employing the methodology of (Borusyak et al., 2021), particularly suited for analyzing staggered policy adoptions like ours.
-This method accurately estimates the effects of anti-discrimination laws, considering their varied implementation times.
-For additional validation, we include results from a simpler difference-in-differences estimator in the appendix, providing robustness to our findings.
+The inclusion of county-year fixed effects remove the need for county controls, addressing the lack of local support data for same-sex marriage.
+We refine the approach of (Burn, 2018) by employing the methodology of (Borusyak et al., 2021), designed for analyzing staggered policy adoptions like ours.
+This method accurately estimates the effects of anti-discrimination laws, considering their staggered implementation times.
 In Figure 2, we present the distribution of same-sex partnerships by state from 2005 to 2019, and in Table 2, we present the 10 counties with the largest share of same-sex partnerships.
 Unsurprisingly, the distribution of same-sex partnerships by state and by county is skewed toward more progressive states and counties with large cities in them that are known for having a large LGB population like San Francisco, the District of Columbia, New York City, and Boston.
-The LGB population varies from the straight population in many dimensions including geography.
-We present descriptive statistics showing the differences in the LGB population compared to the straight population broken down by education level and sex in Table 3.
+The LGB population varies from the heterosexual population in many dimensions including geography.
+We present descriptive statistics showing the differences in the same-sex couple population compared to the different-sex couple population broken down by education level and sex in Table 3.
 Descriptively, there is a clear difference in labor market outcomes and characteristics between men and women in same-sex partnerships and those in different-sex partnerships when controlling for education.
@@
 Formally, we estimate this equation on labor supply:
-$\begin{split}Y_{ict}=\alpha+\beta_{0}(SSP_{ict}\times ENDA_{ct}\times I_{\textrm{Sex=Female})}+\beta_{1}(SSP_{ict}\times ENDA_{ct}\times I_{\textrm{Sex=Male})}\\ +(I_{c}\times I_{t}\times I_{\textrm{Sex=Male}}))\gamma_{sct}+(SSP_{ict}\times I_{c}\times I_{\textrm{Sex=Male}}))\theta_{S,s,c}\\ +(SSP_{ict}\times I_{t}I_{\textrm{Sex=Male}}))\theta_{S,s,t}+X^{\prime}_{ict}\delta+\epsilon_{ict}\end{split}$ (1)
-In our regression, equation (1) defines the labor market outcomes $Y_{ict}$ for individual $i$, in county $c$, at time $t$ with sex $s$ and same-sex partnership status $S$.
-These include the labor force participation, employment status, or the real value of wages, in 1999 dollars.
-The binary variable $SSP_{ict}$ indicates if an individual is in a same-sex partnership.
-The model explores the interaction between being in a same-sex partnership and the presence of county-level Employment Non-Discrimination Acts $(SSP_{ict}\times ENDA_{ct})$.
-[^11] Treatment status is computed using the earliest year of ENDAs at the county or state level, with individuals remaining classified as “treated” thereafter throughout the dataset.
-Given the ACS’s 23-month reference period for earnings, as discussed by Hansen et al.
-(2020), there’s a potential for bias in treatment assignment.
-This extended timeframe might not precisely align with the exact implementation dates of ENDAs, potentially biasing the immediate effects of these policies towards zero.
-The fixed effects for county-year-sex interactions $(\gamma_{ct})$, control for unobserved heterogeneity.
-Interactions between the same-sex partnership status, sex and county $(SSP_{ict}\times I_{c}\times I_{\textrm{Sex=Male}})$ and year fixed effects $(SSP_{ict}\times I_{t}\times I_{\textrm{Sex=Male}})$ capture varying impacts over time and across different regions by sex.
-$X^{\prime}_{ict}$ is a vector of person-specific covariates including race, children, age, and education level - each represented up to a quadratic term to capture potential nonlinear effects.
-The error term $\epsilon_{ict}$ is clustered at the county level to account for within-group correlation.
-[^12] Since the ACS only identifies metro counties, the county fixed effect is close to a pseudo-county fixed effect.
-The rural counties are all compared together with a true county fixed effect for metro area counties.
-This specification aims to isolate the effects of anti-discrimination laws on the labor market outcomes for individuals in same-sex partnerships by sex.
-Our sample is limited to people aged 25 to 65, and we estimate this model jointly by sex using the ACS person weights.
-When examining wages, the $\ln(Y_{ict}+1)$ specification is adopted to include both employed and unemployed individuals in the analysis, recognizing employment as endogenous to anti-discrimination laws.
-This approach ensures a comprehensive evaluation of the law’s impact on the entire income spectrum.
-In equation (1), the differential effect of being a cohabiting same-sex couple by sex in the United States over time is captured by $\theta_{S,s,t}$.
-The county fixed effects for cohabiting same-sex couples by sex $\theta_{S,s,c}$ capture how the wage differential varies across counties by sex.
-The parameters of interest is $\beta_{0}$ and $\beta_{1}$, which capture how the differential between homosexuals and heterosexuals in counties with ENDAs changed after the law’s passage by sex.
-Hence, $\beta_{1}$, will give the effect of ENDAs on the labor supply and pay gap/premium for males, and $\beta_{0}$ for females.
-[^13] We recognize the heterogeneous impact of anti-discrimination laws across different regions, yet our analysis aims to estimate an average effect for those directly subject to the laws.
-One concern is that the distribution of LGB workers is not random, and states/counties that pass anti-discrimination laws have a higher concentration of LGB workers.
-However, using county fixed effects will make comparisons within a given county, so San Francisco will be compared to San Francisco.
-A potential concern would be if there is sorting of LGB workers to areas with anti-discrimination laws protecting them, changing the composition of the area.
-We show that there is little evidence to support this in the robustness check section.
-To interpret $\beta_{0}$ and $\beta_{1}$ as being the causal effect of enacting employment protections for sexual orientation, it must be the case that there are no other factors related to changes in the gay/lesbian-heterosexual wage differential in states that do and do not have LGBT employment protections.
-To accurately estimate $\beta_{0}$ and $\beta_{1}$ while considering the staggered implementation of ENDAs, we employ the methodology from (Borusyak et al., 2021).
-This approach uses observations only for units and periods that are not yet treated to infer the never-treated potential outcome for each treated unit.
-Our identifying assumption in this framework is that, in the absence of anti-discrimination laws, the trends in labor supply and wage disparities in counties with such laws would have mirrored those without them.
-This difference-in-differences framework hinges on parallel trends across all groups and periods, assuming no prior anticipation of the laws.
-While directly testing this parallel trend assumption is challenging, examining pre-treatment trends can provide some indication of its validity.
-We examine the outcomes relative to the timing of the laws by estimating an event study model, where we estimate coefficients for each period relative to the treatment time.
-This approach allows us to assess the impact of the laws over time and ensure a thorough understanding of their effects before and after implementation.
-We plot the effect of the law on the labor supply and pay gap/premium in each year relative to the start date.
-The effect of the law in the year before the law is enacted is normalized to zero.
+$Y_{ipct}$ $=\alpha\;+\;\beta_{1}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{ENDA}_{ct}\bigr{]}\;+\;\sum_{s=1}^{S}\lambda_{s}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{SSM}_{st}\times\mathbf{1}_{\text{state}(c)=s}\bigr{]}$ (1)
+$\quad+\theta_{ct}+\mu_{pc}+\kappa_{pt}+X_{ipct}^{\prime}\,\delta+\varepsilon_{ipct}.$
+In our regression, equation (1) defines the labor market outcomes $Y_{ipct}$ for individual $i$, in partnership type $p$, residing in county $c$, at time $t$.
+These outcomes include labor force participation, employment status, and the percentile rank of real wages (in 2019 dollars).
+The binary variable $\mathrm{SSP}_{ipct}$ indicates whether an individual is in a same‐sex partnership (1) or a different‐sex partnership (0).
+$\mathrm{SSM}_{st}$ indicates whether state $s$ recognizes same‐sex marriage in year $t$, and $\mathbf{1}_{\{\mathrm{state}(c)=s\}}=1$ if county $c$ lies in state $s$.
+The error term $\epsilon_{ipct}$ is clustered at the county level to account for within-county correlation.
+[^12] The ACS only identifies metro counties, so the county fixed effect is close to a pseudo-county fixed effect.
+Rural counties within a state are compared together with a true county fixed effect for metro area counties.
+Our key variable of interest is the interaction of same‐sex partnership status with the presence of a county‐level Employment Non‐Discrimination Act, $\mathrm{ENDA}_{ct}$.
+[^13] Treatment status is computed using the earliest year of ENDA implementation at the county or state level, with individuals remaining classified as “treated” thereafter.
+We include three sets of high‐dimensional fixed effects to account for a wide range of unobserved factors.
+First, $\theta_{ct}$ (county×$\times$×year) absorbs any time‐varying shocks or policy changes common to all individuals within a county in a given year.
+Second, $\mu_{pc}$ (partnership×$\times$×county) controls for any unobserved, time‐invariant characteristics that are specific to each partnership type within a county.
+Third, $\kappa_{pt}$ (partnership×$\times$×year) accounts for year‐specific shocks that affect each partnership type differentially.
+The vector $X_{ipct}^{\prime}\delta$ includes additional demographic controls (e.g., race, age, and education).
+The coefficient of interest, $\beta_{1}$, identifies how ENDA coverage $\mathrm{ENDA}_{ct}$ differentially affects same‐sex partners, net of any confounding from same‐sex marriage legalization.
+The summation term $\sum_{s=1}^{S}\lambda_{s}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{SSM}_{st}\bigr{]}$ partials out state‐specific labor‐market impacts of marriage legalization on same‐sex couples.
+The key variation used to estimate $\beta_{1}$ is the differential timing of when counties adopted ENDAs, and how this interacted with the labor market outcomes for individuals in same-sex partnerships relative to those in different-sex partnerships.
+This variation allows us to isolate the effect of anti-discrimination laws on labor market outcomes by comparing same-sex and different-sex couples across counties with and without ENDAs, before and after their passage.
+Thus, $\beta_{1}$ captures the differential impact of ENDA on labor supply and wage outcomes specifically for individuals in same-sex partnerships.
+[^14] We recognize the heterogeneous impact (e.g., Burn (2018) of anti-discrimination laws across different regions, yet our analysis aims to estimate an average effect for those directly subject to the laws.
+We limit our sample to people aged 25 to 65, and we estimate this model separately by sex using the ACS person weights.
+Following Chen and Roth (2024), we compute the percentile rank of annual and hourly real wages.
+[^15] Given the ACS’s 23-month reference period for earnings, as discussed by Hansen et al.
+(2020), this extended timeframe might not precisely align with the implementation dates of ENDAs, potentially biasing the immediate effects of these policies towards zero.
+To interpret $\beta_{1}$ as the causal effect of enacting employment protections for sexual orientation, we assume that no other factors are driving changes in the same-sex versus different-sex wage differential in states that do and do not have LGBT employment protections.
+One concern is that the distribution of workers in same-sex partnerships is not random, and states/counties that pass anti-discrimination laws have a higher concentration of workers in same-sex partnerships.
+However, with county-year fixed effects, comparisons are made among individuals within the same county in the same year, ensuring that workers in San Francisco are compared to others in San Francisco in that same year.
+A potential concern would be if there is sorting of workers in same-sex partnerships to areas with anti-discrimination laws protecting them, changing the composition of the area.
+To account for this, we remove individuals who moved in the past year as a robustness check.
+[^16] While we address potential sorting by excluding recent movers, unobserved migration patterns could still affect the composition of same-sex partnerships in jurisdictions with anti-discrimination laws.
+Another potential concern is that same-sex marriage legalization may confound our estimates.
+We address this by explicitly partialing out any state-specific effects of marriage legalization for same-sex partners.
+In particular, we include the term $\sum_{s=1}^{S}\lambda_{s}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{SSM}_{st}\times\mathbf{1}_{\text{state}(c)=s}\bigr{]}$, which absorbs any differential labor-market impacts associated with the varying timing of same-sex marriage legalizations across states.
+This ensures that our key coefficient, $\beta_{1}$, isolates the independent effect of the local Employment Non-Discrimination Act.
+While these controls mitigate potential bias, we caution that, despite our efforts, overlapping policies may still exert residual confounding influences on the estimated effects.
+To address potential biases from staggered policy adoption, we employ the methodology of Borusyak et al.
+(2021).
+Unlike standard TWFE models, which may misattribute post-treatment effects from early-treated units to later-treated ones, their approach uses only untreated observations to infer counterfactual outcomes for treated units.
+This strategy minimizes contamination and yields a more reliable estimate of $\beta_{1}$.
+Our identification hinges on the parallel trends assumption, that in the absence of anti-discrimination laws, labor market trends in counties with and without such laws would have evolved similarly.
+[^17] The difference-in-differences strategy hinges on the parallel trends assumption across groups and periods, with no anticipation of the laws prior to their implementation.
+We assess this assumption using event studies that estimate the dynamic effects of the policy, allowing us to observe pre-treatment trends and validate the plausibility of our identification strategy.
 We report the coefficients estimated via Borusyak et al.
-(2021) in the main text, and standard TWFE in the Appendix.
-This paper’s main contribution is applying a staggered policy design approach, providing an understanding of the average effects of anti-discrimination laws.
-Our approach highlights the importance of considering local contexts when assessing the efficacy of legislative measures.
+(2021) in the main text using a fully saturated triple-difference (DDD) specification, and standard TWFE along with other difference-in-differences (DiD) estimators in Section 5.
+As we show in Section 5, the estimates for same-sex and different-sex individuals estimated separately are similar across estimators.
 4 Primary Findings
-4.1 Extensive & Intensive Margins, Wages
-In Table 4, we present the regression on the extensive & intensive margins of labor supply with $\beta_{1}$ giving the effect of anti-discrimination laws on gay men’s labor supply and $\beta_{0}$ giving the effect for lesbian women.
-Columns (1-2) show that anti-discrimination laws have a significant effect in reducing the labor force participation and employment gap for gay men.
-The labor force participation rate gap is reduced by 2.1 p.p., and the employment gap is reduced by 1.5 p.p.
-Conversely, the labor force participation and employment premium that lesbian women have over their straight counterparts is reduced by 1.7 p.p.
-and 2.3 p.p., respectively.
-Columns (3-6) present the regression results on the intensive margin of labor supply as measured by weekly hours worked and weeks worked.
-We find contrasting effects of anti-discrimination laws on male and female outcomes.
-We find that the intensive labor supply gap for gay men as measured by weekly hours worked and weeks worked is insignificantly reduced, and the intensive labor supply premium for lesbian women is significantly reduced when measured in weeks worked.
-We later show that the positive impact for lesbian women on weekly hours and the negative effect on weeks worked for gay men are gone once controlled for occupational fixed effects.
-In Table 5, we present the regression results on the wage gap as measured by hourly wages and annual wage earnings.
-These variables received a $\ln(Y_{ict}+1)$ transformation, to handle zero values.
-[^14] We see small differences in effect sizes in a log-only specification on the employed-only population.
-We run the regressions jointly by sex and present the regressions showing how the sexual orientation anti-discrimination laws impact both employed and unemployed people in columns (1) and (2) and limit the regressions to only those who are employed in columns (3) and (4).
-Our preferred specification is to look at both employed and unemployed people in the sample as employment is endogenous.
-Conditioning on employed workers would create an endogeneity problem since labor force participation and employment are significantly impacted by anti-discrimination laws.
-Given Table 4, the sample of employed workers could be changing as a result of the law changes, which would lead to a change in the sample composition.
-We present the regressions limited to employed workers to replicate past work on the wage gap/premium where it is common to only examine employed workers.
-Gay men experience a notable increase in hourly wages of 6.2%, which is statistically significant at the 95% confidence level.
-Additionally, there is a substantial increase of 22.2% concerning their annual earnings, indicating significant progress towards income equality.
-On the other hand, lesbian women see an insignificant change in their annual earnings premium and an insignificant reduction in their hourly wage premium.
-The pay gap/premium is reduced quite significantly once the sample is limited to employed workers in columns (3) and (4), which is unsurprising.
-The effects also turn insignificant, which shows that they come from increased labor force participation and employment rates.
-4.2 Occupation Fixed Effects
-LGB workers tend to sort to different occupations and industries based on sex and sexual orientation (Black et al., 2007).
-They show that gay men tend to sort to more female-dominated fields and lesbian women sort to more male-dominated fields relative to their heterosexual counterparts.
-It is typical to not consider those factors because discrimination affects occupation and industry choice, which would make occupation and industry choice endogenous.
-While controlling for occupation will bias the wage gap, it can still be instructive for seeing the effect of anti-discrimination laws on the wage and labor supply gap.
-We run these regressions with occupation fixed effects using four-digit occupation codes in the ACS to see how the pay gap/premium changes after anti-discrimination laws are passed.
-Using occupation-fixed effects will examine the effect of anti-discrimination laws within an occupation.
-One potential mechanism for anti-discrimination laws is that gay workers sort into more in-demand occupations, lowering their relative labor supply and wage gap, while lesbian workers do the opposite.
-However, we show that even when controlling for occupation there is a significant change in relative labor supply for men and women in same-sex partnerships and a significant change in relative pay for women in same-sex partnerships, indicating there is a significant effect of these laws even within an occupation.
-We present the extensive labor supply and pay regressions with occupation fixed effects in Table 6 and 7.
-In these tables, there is a smaller pay gap for gay men and a smaller pay premium for lesbian women when controlling for occupation or industry.
-However, qualitatively the results are similar to the main results in Table 4 and Table 5.
-Gay men see a significant reduction in the labor supply gap, and lesbian women see a significant reduction in their labor supply and annual wage premium.
-These findings suggest that anti-discrimination laws have a significant effect even within occupations.
-The effect does not come from workers shifting into occupations with less discrimination, but even when conditional on occupation choice, there is a significant reduction in differences in labor supply across sexual orientations for men and women.
-4.3 Impacts Over Time
-We use Equation (1) with the inclusion of occupation-fixed effects and conduct event study regressions via Borusyak et al.
-(2021).
-This approach allows us to analyze how the impact estimates evolve.
-First, we present the event study regressions for male and female extensive labor supply outcomes in Figure 3.
-Panel (a) illustrates the effects on labor force participation rates, whereas Panel (b) shows the effects on employment levels.
-The plots reveal an initial jump in both employment and labor force participation rates among gay men, a trend that appears to wane over subsequent years.
-This pattern hints at an early beneficial impact that seems to moderate with time.
-Conversely, the trajectory for women follows a different pattern, indicating initial adverse effects on employment and labor force participation rates.
-The pre-trends look fairly parallel, which, while not directly confirming the parallel trends assumption for the post-period, lends it greater plausibility.
-These visualizations indicate a temporary increase in labor force participation and employment rates for gay men following the enactment of ENDAs, contrasting with a negative impact on the same metrics for lesbian women.
-We present the event study regressions for male and female pay penalty/premium outcomes in Figure 4.
-The figures suggest a delayed response in wage parity gains for gay employees compared to their heterosexual counterparts, with a notable but transient increase in relative gay wages observed a year post-ENDA implementation.
-This delayed effect may come from implementation lags and the labor market outcomes captured in the ACS, which reflect earnings over a 23-month reference period.
-Conversely, there is a marked decline in relative wages for lesbian women following the enactment of the law, becoming significant one year after its introduction.
+4.1 Extensive & Intensive Margins
+In Table 4, we present regression results that examine the extensive and intensive margins of labor supply for same-sex partnerships.
+The coefficient $\beta_{1}$ captures the effect of anti-discrimination laws on the labor supply outcomes for men in male same-sex partnerships and for women in female same-sex partnerships.
+Panel A focuses on men in male same‐sex partnerships.
+Columns (1) and (2) show that anti-discrimination laws reduce both the labor force participation and employment gaps by approximately 3.8 and 3.7 percentage points, which translates into reductions of about 78% and 77% respectively.
+These coefficients are statistically significant at the 1% level.
+Columns (3) through (6) address the intensive margin, indicating modest increases in weekly hours and weeks worked among all men.
+However, when the sample is restricted to employed men (Columns (5) and (6)), these effects become statistically insignificant.
+This pattern suggests that the primary channel through which these laws operate is by increasing the rate of employment.
+Panel B reports analogous estimates for women in female same‐sex partnerships.
+In contrast, the extensive margin results for women (Columns (1) and (2)) yield statistically insignificant coefficients of $-0.37$ and $-0.50$ percentage points.
+Women in same‐sex partnerships exhibit higher labor force participation rates than women in different‐sex partnerships.
+This higher baseline may partially explain the lack of significant improvement on the extensive margin, as there is less room for gains.
+On the intensive margin, Column (3) shows a modest increase of 0.95 weekly hours (significant at the 10% level) for all women, while Column (4) finds a small, statistically insignificant reduction in weeks worked.
+When focusing on employed women (Columns (5) and (6)), we observe a statistically significant increase of 0.70 weekly hours (at the 5% level) alongside a significant decrease of 0.89 weeks worked (at the 1% level).
+These mixed findings suggest that anti-discrimination laws may lead employed women to work longer hours per week but in fewer weeks overall, reflecting a complex adjustment on the intensive margin.
+4.2 Wages
+As shown in Table 4, anti-discrimination laws significantly affect labor force participation and employment outcomes, thereby potentially altering the composition of the employed sample and raising endogeneity concerns for wage analyses.
+Table 5 shows regression results on the wage gap, using real hourly wages and annual earnings in 2019 dollars, measured by percentile ranks and computed jointly for men and women.
+As Chen and Roth (2024) argue, this robust, scale-invariant metric reduces the impact of outliers and reflects an individual’s relative wage position, allowing us to evaluate how anti-discrimination laws shift wage standings for individuals in same-sex partnerships.
+For men in same-sex partnerships, the increased employment rates reported in Table 4 are accompanied by significant wage gains.
+Specifically, anti-discrimination laws boost hourly wages by roughly 2.7 percentiles and annual earnings by 3 percentiles on average.
+Furthermore, these laws raise the probability of ranking above the 25th and 50th percentiles in the hourly wage distribution by 3.6 and 3.4 percentage points, respectively, although the effect at the 75th percentile is not significant.
+In the annual earnings distribution, anti-discrimination laws improve the likelihood of being in the upper tiers by 3.6, 5.2, and 3.7 percentage points at the 25th, 50th, and 75th percentiles, respectively.
+For women in same-sex partnerships, the wage distribution remains largely unaffected by anti-discrimination laws, as evidenced by statistically insignificant estimates across the board.
+This finding is consistent with the muted effects on employment outcomes observed in Table 4 for women.
+These findings suggest that the wage gains for men in same-sex partnerships operate primarily through increased employment.
+New entrants to the labor force, individuals who were previously at the bottom of the wage distribution, enter at various points, thereby shifting the overall distribution upward.
+4.3 Impact Dynamics
+We estimate Equation (1) using event study regressions following Borusyak et al.
+(2021) to trace the evolution of the impact estimates.
+Figure 3 displays the results for extensive labor supply outcomes: Panel (a) shows labor force participation rates, while Panel (b) presents employment levels.
+For men in male same‐sex partnerships, both panels reveal an immediate increase in labor force participation and employment following the enactment of ENDAs, although this positive effect wanes over time.
+In contrast, we see no impacts on women in same-sex partnerships.
+The pre-treatment trends are largely parallel, supporting the plausibility of the parallel trends assumption, yet the post-treatment dynamics diverge sharply by gender.
+The evidence here suggests that ENDAs lead to a temporary boost in labor market entry for men in same-sex partnerships, while not affecting employment and participation rates for women in same-sex partnerships.
+Figure 4 presents the event study regressions for male and female pay disparities.
+The results mirror the trends observed in Figure 3, with a significant wage jump occurring one year after the law’s enactment for men in same-sex partnerships.
+However, the difference between the one-year and four-year post‑enactment estimates is relatively modest, suggesting that the initial wage improvement for men stabilizes and remains persistent over time.
+We find no statistically significant effects on wage disparities for women in any post‑ENDA period.
+These findings indicate that while anti‑discrimination laws yield immediate and lasting wage benefits for men in same‑sex partnerships, the impact on wage outcomes for women in same-sex partnerships is insignificant.
 5 Robustness Checks
+To ensure the robustness of our findings on the labor market impacts of anti-discrimination laws, we conduct a series of sensitivity analyses addressing potential sources of bias.
+Specifically, we examine (i) the endogeneity of law adoption using state-level support for same-sex marriage as a proxy for public sentiment, (ii) the potential confounding influence of same-sex marriage legalization, (iii) the role of migration patterns, (iv) geographic heterogeneity between metropolitan and non-metropolitan areas, (v) reporting changes associated with ACS redesign, (vi) the impact of applying sample weights, and (vii) the validity of our identification strategy using placebo tests and alternative estimators.
 5.1 Endogenous Adoption of Anti-Discrimination Laws
-One concern in this analysis is the endogenous adoption of anti-discrimination laws.
 Anti-discrimination laws are not randomly distributed.
@@
 The main concern is that there is an unobservable factor like general sentiment toward LGB workers that affects both the passage of laws as well as the labor market outcomes for LGB workers.
-In our main regressions, we alleviate this concern by using county-year-sex fixed effects.
+In our main regressions, we alleviate this concern by using county-year fixed effects.
 Nonetheless, we investigate the endogenous adoption of anti-discrimination laws at the state level.
@@
 Specifically, we estimate this equation:
-$\text{Support}_{st}=\sum_{j=-5,j\neq-1}^{j=5}\psi_{j}\mathbbm{1}(\text{Laws}_{st}=j)+\phi_{s}+\delta_{t}+\epsilon_{ts}$ (2)
-[…]
-The event study supports the idea that anti-discrimination laws at the state level significantly increase the support for same-sex marriage at the state level.
-The trends before the legislation exhibit similar trends, with a noticeable jump in the fraction of individuals favoring same-sex marriage.
-This pattern suggests that states, irrespective of their legal stance on anti-discrimination, shared similar trajectories in support for same-sex marriage before the enactment of such laws, and only began to diverge following the implementation of the legislation.
-It might be argued that the passage of these laws was a consequence of shifting attitudes towards same-sex marriage at the time of enactment.
-Still, if the adoption of these laws were endogenous, one would expect to see a steady climb in support for same-sex marriage leading up to the law’s passage, with the legislation itself having minimal impact on this trend.
-However, the distinct and significant jump in support concurrent with the introduction of anti-discrimination laws lends more weight to the argument that it is the legislative changes that are driving the increase in support for same-sex marriage and, likely, increasing the overall sentiment towards LGB individuals, rather than the other way around.
-The jump seen in state-level sentiment toward LGB workers following state anti-discrimination law passages could suggest that there is a corresponding jump in sentiment at the local level that is unobservable and potentially acting as a mechanism, affecting the labor supply and pay for LGB workers.
-5.2 Sorting and Increased Reporting
-One finding from Klawitter and Flatt (1998) suggests that LGB workers sort to areas with anti-discrimination laws.
-If LGB workers were sorting to areas that recently passed anti-discrimination laws, then it could violate the identifying assumption.
-It would be concerning if high-wage LGB workers migrated from unprotected areas to areas that recently passed anti-discrimination laws since it would lower the average wage in untreated counties and increase the average wage in treated counties.
-Another potential concern would be increased reporting for same-sex partnerships.
-Same-sex partnerships, while less stigmatized in 2005-2019 than in previous years, were still heavily stigmatized.
-Anti-discrimination laws could signal to LGB people that their community is more accepting, and people may be more likely to declare that they are in a same-sex partnership.
-This change in reporting could affect the composition of our sample, which in turn could violate the identifying assumption.
-To address this concern, we run the regression below to see if there is an increase in same-sex partnerships in a given county after an anti-discrimination law is passed:
-$\text{SSP}_{ict}=\beta_{0}\text{Laws}_{ct}+\beta_{1}X_{i}+\beta_{2}\gamma_{st}+\mu_{c}+\delta_{t}+\epsilon_{itcs}$ (3)
-In this regression, $\beta_{0}$ is the coefficient of interest, and it will give the effect of anti-discrimination laws passing on the number of same-sex partnerships in a given county whether that is from in-migration or increased reporting of same-sex partnerships.
-We present these results in Table 8.
-We find no significant effect of anti-discrimination laws on the number of same-sex partnerships at the county level, suggesting there is minimal sorting or change in reporting following the passage of anti-discrimination laws.
+$\text{Support}_{st}$ $=\sum_{\begin{subarray}{c}j=-4\\ j\neq-4\end{subarray}}^{4}\psi_{j}\,\mathbbm{1}\Bigl{(}\text{ENDA}_{st}=j\Bigr{)}+\phi_{s}+\delta_{t}+\epsilon_{ts}.$ (2)
+Supportst denotes the level of support for same-sex marriage in state $s$ during year $t$, while $\mathbbm{1}(\text{ENDA}_{st}=j)$ is an indicator variable equal to 1 if state $s$ has had a sexual orientation anti-discrimination law for $j$ years by time $t$.
+We estimate the coefficients $\psi_{j}$, which capture the effect of anti-discrimination laws on support for same-sex marriage, using an event study framework within a difference-in-differences design.
+This strategy relies on the identifying assumption that, in the absence of anti-discrimination laws, support for same-sex marriage in states that passed these laws would have followed parallel trends to those in states that did not.
+Figure 5 presents the event study plot.
+The event study indicates that state-level anti-discrimination laws significantly boost support for same-sex marriage.
+Prior to the legislation, trends in support moved in parallel across states regardless of their legal stance, suggesting that, absent the laws, these trajectories would have remained similar.
+Although one might argue that the passage of such laws was driven by a pre-existing increase in support for same-sex marriage, an endogenous explanation would predict a gradual rise in support leading up to the law’s enactment, with little change at the moment of passage.
+Instead, we observe a distinct, significant jump in support coinciding with the implementation of the legislation.
+This pattern implies that the legislative changes themselves are likely driving the increase in support for same-sex marriage, and possibly improving overall attitudes toward LGB individuals.
+Moreover, the observed jump in state-level sentiment toward LGB workers following the passage of anti-discrimination laws may reflect an unobserved corresponding increase in local sentiment.
+This local-level change could act as a mechanism through which these laws impact the labor supply decisions of LGB workers.
+5.2 Marriage Laws
+Same-sex marriage legalization may confound our identification strategy.
+For example, Utah legalized same-sex marriage in 2015, the same year it passed its statewide sexual orientation anti-discrimination bill, while Iowa and Arizona experienced similar overlaps (with Iowa enacting its anti-discrimination law in 2007 and legalizing same-sex marriage in 2009, and Arizona in 2014).
+Prior studies have shown that same-sex marriage legalization can differentially affect employment outcomes for same-sex couples.
+To address this issue, we included fixed effects that interact same-sex couple status with state-level same-sex marriage legalization.
+Nevertheless, to further isolate the impact of anti-discrimination laws, we restrict our sample to periods before same-sex marriage became legal in each state (e.g., excluding observations from 2008 onward for states that legalized in 2008).
+Table 6 presents results restricting the sample to periods before same-sex marriage legalization.
+Although this approach removes nearly half of the same-sex partnership observations and reduces statistical power, the qualitative patterns remain robust.
+Specifically, the point estimates for both labor force participation and employment remain comparable, exceeding 3 percentage points, with labor force participation retaining statistical significance.
+However, the employment coefficient for men declines from 0.037 to 0.0319, becoming borderline insignificant (p-value = 0.102), and similar shrinkages are observed on the intensive margin.
+In contrast, for women in same-sex partnerships, the intensive margin estimates are slightly larger.
+Overall, this sensitivity analysis indicates that our identification strategy is not driven by concurrent same-sex marriage policies, as the observed effects of anti-discrimination laws on labor market outcomes for same-sex couples persist despite the reduced sample size.
+5.3 Migration
+Next, we account for the possibility that migration patterns could influence our results.
+If LGB individuals were disproportionately moving to areas with recently enacted anti‑discrimination laws, it could bias the treatment effects.
+To address this concern, we restrict our sample to individuals who had not moved in the past year and re‑estimate the models presented in Table 4.
+By focusing on a more stable population, we can better isolate the effects of anti‑discrimination laws on labor outcomes.
+The corresponding estimates, reported in Table 7, remain consistent with the full sample results, providing further robustness and suggesting that migration does not significantly alter our conclusions.
+Interestingly, the estimates for men in same‑sex partnerships are slightly larger for this subset, which may indicate that the positive employment effects are even more pronounced among those who are less mobile.
+5.4 Metropolitan Areas
+We estimate our model specifically for residents in metropolitan areas to address concerns about potential differences between urban and rural populations.
+By focusing on metro areas, where labor markets may function differently due to higher population density and greater economic diversity, we aim to test whether the effects of anti-discrimination laws vary by geographic location.
+Table 8 shows that when the sample is restricted to metropolitan areas, the overall patterns for men in same‑sex partnerships remain consistent, though the estimated effects of anti‑discrimination laws on labor supply are slightly smaller than in the full sample.
+This may suggest that the urban labor market environment, with its higher baseline employment levels and different institutional dynamics, somewhat tempers the positive impact seen elsewhere.
+In contrast, for women in same‑sex partnerships, the estimates become significantly negative across several outcomes, labor force participation, employment, and weeks worked.
+One plausible explanation is that ENDAs may improve employment stability in urban labor markets, and with high childcare costs in metro areas, a stable job by one partner might encourage women to opt out of the workforce.
+With a secure income from the other partner, some women may choose to stay home to manage childcare responsibilities rather than work additional hours or remain continuously employed.
+This pattern highlights clear geographic heterogeneity, as the dynamics in urban settings differ markedly from those in the broader sample for women in same‑sex partnerships.
+5.5 Reporting Changes
+Prior research has highlighted higher rates of misclassification errors before 2008, where different-sex couples were sometimes incorrectly identified as same-sex couples due to typos in the sex variable.
+In response, and considering the 2008 ACS redesign aimed at reducing such errors (O’Connell and Lofquist (2009)), we restrict our sample to 2008–2019 to test the robustness of our findings reported in Table 4.
+Table 9 presents the estimates based on this restricted sample.
+For men in same-sex partnerships, the results remain qualitatively similar to our full sample analysis.
+The estimated impacts on both labor force participation and employment continue to be positive and statistically significant, with coefficients that are comparable in magnitude to those in Table 4.
+Among women, the extensive margin estimates remain small and statistically insignificant, while the intensive margin results are largely consistent with the full sample, albeit with some variations in magnitude.
+These findings suggest that potential misclassification errors in earlier years do not drive our main conclusions.
+The consistency of the results between the full sample (2005–2019) and the restricted sample (2008–2019) reinforces the robustness of our empirical strategy and the validity of our conclusions regarding the impact of anti-discrimination laws on labor market outcomes for same-sex couples.
+5.6 Weighting
+Although the ACS recommends applying sample weights to ensure representativeness, we conduct a robustness check by re‐estimating our models without these weights, following Solon et al.
+(2015).
+This approach allows us to verify that our main results are not driven solely by the weighting scheme.
+The results in Table 10 indicate that the overall impacts of anti‑discrimination laws on labor supply remain robust even when ACS weights are removed.
+For men in same‑sex partnerships, the magnitude of the estimated effects on both labor force participation and employment is slightly smaller compared to the weighted estimates, but they continue to be statistically significant.
+This suggests that the positive impact of these laws for men is not driven solely by the weighting procedure.
+For women, the pattern of the results is similar to the weighted sample, with some variations in magnitude across different labor outcomes.
+5.7 Placebo Tests and Alternative Estimators
+We next conduct a series of placebo tests and apply alternative estimators to assess the impact of anti‑discrimination laws on labor market outcomes among both same‑sex and different‑sex couples.
+Specifically, we utilize four distinct methods to assess the impact of anti‑discrimination laws on labor market outcomes among same‑sex and different‑sex couples: TWFE, Callaway and Sant’Anna (2021), Sun and Abraham (2021), and Borusyak et al.
+(2021).
+As a placebo test, we apply the DiD framework to different‑sex couples, who are not the target of these laws, and thus we expect no significant changes in their labor market outcomes.
+Panel (a) of Tables 11 and 12 report the DiD estimates for different‑sex couples, while Panels (a) and (c) of Figures 6 and 7 display the corresponding event study plots.
+Our placebo tests show that the DiD estimates yield no statistically significant impacts on labor force participation or employment rates for different‑sex couples.
+The pre‑ENDA coefficients are insignificant, confirming the absence of differential trends prior to the law’s enactment.
+Most post‑ENDA coefficients remain statistically insignificant, with only a few minor exceptions of small magnitude, largely consistent with a 5% false positive rate.
+Next, Panel (b) of Tables 11 and 12 presents the impact of anti‑discrimination laws on same‑sex couples, while Panels (b) and (d) of Figures 6 and 7 display the corresponding event studies.
+In these analyses, we compare labor market outcomes for same‑sex couples in regions that have enacted the laws with those in regions that have not.
+The results reveal that male same‑sex couples in areas with anti‑discrimination laws experience significant improvements in labor market outcomes, particularly through increased labor force participation and higher employment rates.
+For female same‑sex couples, however, we observe no significant effects on the extensive margin; instead, the positive and significant impact is confined to the intensive margin, as evidenced by an increase in average weekly hours worked.
+Hence, our alternative estimators yield similar point estimates and comparable confidence intervals, reinforcing the robustness of our findings.
 6 Discussion
 6.1 Differences in Response by Sex
-The results paint an interesting picture of the effect of anti-discrimination laws on the workforce, in that they have differing effects on lesbian and gay workers.
-Lesbian workers see a reduction in the labor supply and pay premium while gay workers see a reduction in their labor supply and pay gap.
-The anti-discrimination laws appear to push the labor market outcome to a greater level of equality across sexual orientation, where lesbian women move toward the labor market outcomes of straight women and gay men move toward the labor market outcomes of straight men.
-One potential explanation for this convergence of labor market outcomes is that employers stop differentiating between workers based on sexual orientation after anti-discrimination laws are passed and that results in greater equality.
-However, much of the labor market differences arise from differences in human capital accumulation and decisions that come before many of the anti-discrimination laws such as occupation/industry choice and desired number of children (Black et al., 2007).
-Another explanation for different responses to anti-discrimination laws for same-sex couples lies in the canonical Becker (1981) theory of the family and division of labor, which is typically used to explain differences in LGB labor market outcomes from their straight counterparts (Black et al., 2007).
-In a male-female household with a “traditional” division of labor, the man specializes in market production, selling his labor for money to buy market goods, and the woman specializes in household production, using her labor to rear children, cook, and take care of the home.
-However, LGB people are less likely to end up in a traditional Beckerian household.
-This expectation of future household composition will induce gay men to be less specialized in market production and lesbian women to be more specialized in market production than their straight counterparts.
-This theory of household specialization is typically used to explain why gay men make less than straight men and lesbian women make more than straight women (Black et al., 2007).
-This household model can also be used to think about responses to anti-discrimination laws.
-One potential explanation for lesbian women becoming more likely to leave the workforce and earn less is that a lesbian household may be more likely to shift to a more traditional Beckerian household where one parent specializes in household production and the other shifts to a more market orientation.
-If anti-discrimination laws decreased the risk of the primary earner getting fired for their sexual orientation, it could induce the secondary earner to switch to more part-time work or exit the workforce to specialize in household production or focus on child-rearing.
-Alternatively, both women in a female same-sex partnership could switch to reducing hours to spend more time in household production as a result of increased job security.
-Households consisting of lesbian women typically have more children than households consisting of gay men.
-Given lesbian women have more children than gay men, it would be unsurprising to see lesbian households switch more to a specialization of their household labor than gay men would, which would help explain why anti-discrimination laws have different effects on lesbian women and gay men.
-To test the different responses by men and women in same-sex partnerships to anti-discrimination laws and the implications of Becker’s household model, we implement an alternative difference-in-difference model.
+The results reveal that anti-discrimination laws affect household labor allocation differently for same-sex couples by gender.
+Men in male same-sex partnerships experience a reduction in their labor supply gap relative to men in different-sex partnerships, while there are no significant average effects for women in female same-sex partnerships.
+One potential explanation is that changes in job security, stemming from reduced discrimination, alter the household’s allocation of labor between market and non‑market activities.
+As documented by Badgett (2006), reduced discrimination in the labor market improves employment prospects and job security for sexual minorities.
+With greater job security for the primary earner, the secondary earner may reduce market participation in favor of increased household production, including child rearing.
+This pattern of adjustment is consistent with models of household bargaining and collective decision‑making (see, e.g., Chiappori (1992) and Lundberg and Pollak (1993)), which predict that improvements in labor market conditions for one partner can induce a reallocation of tasks within the household.
+In this framework, the observed increase in intra‑household differences in hours worked may reflect a shift toward greater specialization, with one partner focusing more on market production while the other devotes more time to household responsibilities.
+Thus, the differential responses between male and female same‑sex partnerships could arise from distinct household preferences and constraints.
+Households consisting of women in same-sex partnerships typically have more children than those consisting of men in same-sex partnerships.
+Given this difference, it would be unsurprising to see female same-sex households shift more toward a specialization of household labor compared to male same-sex households.
+This difference in specialization could help explain why anti-discrimination laws have different effects on women in same-sex partnerships and men in same-sex partnerships.
+To test the different responses by men and women in same-sex partnerships to anti-discrimination laws, we implement an alternative difference-in-difference model.
 Instead of comparing those in same-sex partnerships to those in different-sex partnerships, we compare men and women in same-sex partnerships to examine their differential responses at the household level.
@@
 Specifically, we estimate this model:
-$y_{jct}=\alpha_{0}Law_{ct}+\alpha_{1}FemSSP_{j}+\alpha_{2}Law_{ct}*FemSSP_{j}+\mu_{c}+\delta_{t}+\epsilon_{jct}$
+$Y_{jct}$ $=\alpha_{0}\,ENDA_{ct}+\alpha_{1}\,FemSSP_{j}+\alpha_{2}\,\bigl{(}ENDA_{ct}\times FemSSP_{j}\bigr{)}+\mu_{c}+\delta_{t}+\epsilon_{jct}$ (3)
 The notation is the same as before with two main differences.
-Our interaction term uses an indicator for if household $j$ is a female same-sex partnership, denoted by $FemSSP_{j}$.
-$y_{jct}$ is a variable for outcomes at the household level.
-$\alpha_{2}$ is our coefficient of interest and will show how lesbian households differ in their response to anti-discrimination laws compared to gay households.
-The first outcome examined is an indicator for if the household is a one-earner family.
-We also examine the difference in absolute value between the two partners in terms of weekly hours worked.
-These labor supply measures inform how intra-household labor is divided between partners along the extensive and intensive margin of labor supply.
-Finally, we examine how the anti-discrimination laws may affect child rearing by examining if the households have any children and the number of children they have.
+First, our interaction term uses an indicator for whether household $j$ is a female same‐sex partnership, denoted by $FemSSP_{j}$.
+Second, $Y_{jct}$ now represents outcomes measured at the household level.
+The coefficient $\alpha_{2}$ is our parameter of interest and captures how female same‐sex households differ in their response to anti‑discrimination laws relative to male same‑sex households.
+The first outcome examined is an indicator for whether the household is a one‑earner family.
+We also analyze the intra‑household division of labor by measuring the absolute difference in weekly hours worked between partners, thereby capturing both the extensive and intensive margins of labor supply.
+Finally, we explore the potential effects on child rearing by examining both the presence of children in the household and the number of children present.
 We present the results in Table 13.
 Anti-discrimination laws do not affect the likelihood of same-sex partnerships becoming a one-earner household, but there is an effect on the intensive margin.
-There is a significant effect of anti-discrimination laws in the intra-household difference in hours worked for gay and lesbian households.
+There is a significant effect of anti-discrimination laws in the intra-household difference in hours worked for male same-sex and female same-sex households.
 An increase in the difference of hours worked within the household could suggest a greater specialization in market production for one person and a greater specialization in household production and working fewer hours for the other person in the partnership.
-Lesbian households do see a significant increase in the likelihood of having any child and the number of children they have relative to gay households.
-This evidence tied with Becker’s theoretical work suggests that men and women in same-sex partnerships are responding differently to these anti-discrimination laws.
-Lesbian households begin to have more children than gay households.
-It may appear in the main results that anti-discrimination laws hurt lesbian women because their labor supply and pay premium over straight women shrinks.
+Female same-sex households do see a significant increase in the likelihood of having any child and the number of children relative to male same-sex households.
+This evidence suggests that men and women in same-sex partnerships are responding differently to these anti-discrimination laws.
+Female same-sex households begin to have more children than male same-sex households.
+It may appear in the main results that anti-discrimination laws do not help women in female same-sex partnerships.
 However, their response in the labor market could be driven by a greater desire for children and subsequent changes in household labor allocation due to an increase in job security.
-The corresponding changes in the labor market could represent an increase in welfare for lesbian households if they value their new children more than their lost hours worked and subsequently lost wages.
-Consistent with the theory and empirical findings, lesbian households could become more secure with one woman being the primary earner following anti-discrimination laws.
+The corresponding changes in the labor market could represent an increase in welfare for female same-sex households if they value their new children more than their lost hours worked and subsequently lost wages.
+Consistent with the theory and empirical findings, female same-sex households could become more secure with one woman being the primary earner following anti-discrimination laws.
 6.2 Mechanisms
-The results of this research indicate that sexual orientation anti-discrimination laws could significantly impact the labor market gaps for gay/bisexual men as well as the labor market premiums for lesbian/bisexual women.
+The results of this research indicate that sexual orientation anti-discrimination laws could significantly impact the labor market gaps for men in male same-sex partnerships.
 It is possible that these laws are binding and eliminate significant discrimination in the labor market.
@@
 Our regression results already condition on the change in the county-level support for same-sex marriage via fixed effects.
-In Appendix A, we show that local laws matter significantly, so it seems plausible that local law changes could significantly change local sentiment and impact the labor supply and pay of LGB workers.
+It seems plausible that local law changes could significantly change local sentiment and impact the labor supply and pay of LGB workers.
 Changing public sentiment in conjunction with greater protection for LGB workers seems more plausible as the mechanism than simply greater protection in the workplace.
@@
 Another pitfall to using partnerships to infer sexual orientation is that one could incorrectly infer someone’s sexual orientation from a partnership.
-Bisexual people are a larger group than lesbian or gay people, and this specification could erroneously assign bisexual people to be “straight” through being in a different-sex partnership (Gates, 2011).
-If bisexual workers appear more similar to their lesbian and gay counterparts then the results of quantifying the labor supply and pay gap/premium would likely be biased toward zero and the change in those labor market gaps/premiums would also be biased toward zero.
+Bisexual individuals represent a larger group than lesbian or gay individuals ((Gates, 2011)), and our specification erroneously classifies bisexual people in different-sex relationships as heterosexual.
+This misclassification biases estimates toward zero, as bisexual workers face significant wage penalties ((Mize, 2016)) and their inclusion in the reference group depresses the heterosexual average.
+Future work must disentangle these groups to avoid obscuring inequities.
 Another potential concern in using partnerships to infer sexual orientation is that people in a same-sex partnership may be the group that is most at risk of discrimination.
@@
 It is dependent on several factors that are unknowable in this analysis such as the wages and labor supply for single LGB workers and bisexual workers in different-sex partnerships and how discrimination affects LGB people for those in same-sex partnerships.
+Furthermore, it is important to emphasize that this study cannot identify transgender individuals in the ACS, nor can it estimate the impact of anti-discrimination laws specifically protecting gender minorities.
+This restricts the ability to generalize findings to the entire LGBTQ+ community.
 7 Conclusion
-This analysis is the first quasi-experimental research examining how both local and state anti-discrimination laws on sexual orientation affect the labor supply and pay gap/premium between LGB and straight workers.
-We construct a novel panel dataset on local anti-discrimination laws to capture the granular nature of anti-discrimination laws.
-We find that anti-discrimination laws appear to decrease sexual orientation inequality in the labor market.
-Specifically, we find that anti-discrimination laws significantly reduce the gay labor force participation gap by 2.1 p.p., the employment gap by 1.5 p.p., and increase hourly wages by 6.2%.
-Anti-discrimination laws also reduce the lesbian labor force participation premium by 1.7 p.p., and the employment premium by 2.3 p.p.
-We also show that one potential mechanism that anti-discrimination laws work through is increasing positive sentiments toward LGB Americans as measured by the support of same-sex marriage.
-Finally, we explain the differential response to these laws between gay and lesbian households through Becker’s household specialization model and support the explanation with empirical evidence.
-Before 2020, it was legal in some states to fire someone solely based on their sexual orientation, and states have been less inclined to pass anti-discrimination laws recently.
-Arkansas and Tennessee even passed anti-anti-discrimination laws preventing cities from enacting anti-discrimination laws.
-From 2009 to 2019, more states have barred cities from protecting their residents than the number of states extending protection.
-Our research gives a comprehensive look at a policy that Congress is considering in “The Equality Act” to extend federal protection to sexual orientation and gender identity.
-Our research gives the most comprehensive look at policy and can meaningfully inform the policy discussion around anti-discrimination laws.
-Overall, this research broadly informs the policy decision of extending employment protections to the LGB community by the Supreme Court of the United States.
-Appendix A TWFE Estimates
+This analysis is the first quasi-experimental study examining the impact of both local and state anti-discrimination laws on labor supply and wage disparities between workers in same-sex and different-sex partnerships.
+Utilizing a novel city-level panel dataset capturing the granular detail of anti-discrimination laws, we find that these laws significantly reduce sexual orientation-based inequalities in labor market outcomes.
+Specifically, anti-discrimination laws decrease the labor force participation difference by 3.8 percentage points (78%) and employment difference by 3.7 percentage points (77%) for men in male same-sex partnerships, with insignificant effect for women in female same-sex partnerships.
+We also document important wage effects.
+Sexual orientation anti-discrimination laws lead to significant wage gains for men in same-sex partnerships, with their percentile rank in the hourly wage and annual earnings distribution improving on average by approximately 2.7 and 3 percentiles, respectively.
+Moreover, these laws increase the probability of ranking above the 25th percentile in hourly wages by 3.6 percentage points and above the 50th percentile by 3.4 percentage points, with similar improvements observed in the annual earnings distribution.
+By contrast, the wage distribution for women in same-sex partnerships remains largely unaffected by these laws.
+We identify increased positive societal sentiments toward LGB Americans, reflected in growing support for same-sex marriage, as a potential mechanism through which these laws improve labor market outcomes.
+Furthermore, we explain differential responses to these laws between male and female same-sex households using Becker’s household specialization model and provide empirical support for this explanation.
+Prior to the landmark Supreme Court decision in Bostock v.
+Clayton County (2020), many states permitted employers to fire employees based solely on sexual orientation.
+While Bostock extended federal protections under Title VII to LGBTQI+ workers, legislative momentum for anti-discrimination laws at the state level has slowed significantly since then.
+The Equality Act, reintroduced in the 118th Congress on June 21, 2023, seeks to expand federal civil rights laws to explicitly include sexual orientation and gender identity.
+As of March 2025, the bill is pending in the Senate, facing significant opposition from some senators, particularly concerning its implications for religious freedoms and women’s rights in certain contexts.
+At the state level, legislative momentum for anti-discrimination laws has slowed.
+Between 2009 and 2019, more states barred cities from protecting their residents from discrimination than those that expanded protections.
+Additionally, recent developments have seen states like Arkansas and Tennessee actively prohibiting localities from enacting their own anti-discrimination measures.
+Our research provides critical and timely insights into the impact of anti-discrimination laws on labor market outcomes for individuals in same-sex partnerships.
+By offering a comprehensive evaluation of these laws’ effects, our findings meaningfully inform ongoing policy discussions and judicial decisions, including deliberations by the Supreme Court of the United States, about extending employment protections to the LGB community.
+Appendix A Additional Robustness Checks
+A.1 Event Studies
+Figures A1, A2, A3, A4, A5 present the event study regressions corresponding to the robustness checks discussed in Sections 5.2, 5.3, 5.4, 5.5, 5.6.
+In each figure, we trace the evolution of labor supply impacts on the external margin relative to the enactment of anti-discrimination laws.
+Across all specifications, the pre-treatment coefficients are statistically insignificant, and the post-treatment dynamics closely mirror those observed in our main analyses.
+Specifically, estimates for women in metropolitan areas are smaller on average, unweighted models and those estimated before the legalization of same-sex marriage typically produce slightly lower coefficients, and both the post-2008 sample and non-movers yield somewhat larger effects.
+A.2 Same-Sex Marriage Fixed Effects
+The results in Table A1 reinforce the robustness of our primary findings for men.
+Although the extensive margin effects for men in same-sex partnerships are slightly lower (0.0267 for labor force participation and 0.0270 for employment) than in our primary specification, they remain statistically significant, with generally smaller intensive margin estimates.
+For women, the point estimates on the extensive margin are quite similar in Table A1 and Table 6.
+However, the employment effect for women turns negative and is marginally significant (–0.0212, p $<$ 0.10), which contrasts with the mostly null effects observed elsewhere, aside from the negative effects noted in metropolitan areas.
+This conflicting evidence suggests that further research is needed to better understand the impacts on women.
+Our hypothesis remains that high child-care costs in metropolitan areas may be an important contributing factor.
+A.3 Occupation Fixed Effects
+LGB workers tend to sort to different occupations and industries based on sex and sexual orientation [Black et al., 2007].
+They show that gay men tend to sort to more female-dominated fields and lesbian women sort to more male-dominated fields relative to their heterosexual counterparts.
+It is typical to not consider those factors because discrimination affects occupation and industry choice, which would make occupation and industry choice endogenous.
+Although controlling for occupation can pose selection concerns, because individuals not in the labor force typically lack occupation codes, we nonetheless examine within-occupation effects of anti-discrimination laws on wages and labor supply.
+Specifically, we employ regressions with four-digit occupation fixed effects to assess how wage and labor supply disparities shift following the enactment of these policies.
+[^18] In order to estimate within-occupation effects while still capturing individuals who are not in the labor force, we include a separate “no occupation” category for those without valid occupational codes.
+This approach allows us to retain non-employed respondents in the analysis, rather than excluding them outright due to missing occupation information.
+We acknowledge that categorizing all non-employed individuals into a single group can mask heterogeneity in reasons for not working (e.g., unemployment, home production, or retirement).
+Nonetheless, by creating a distinct “no occupation” category, we preserve a measure of occupational affiliation for those who are employed while ensuring that non-employed individuals remain part of the sample.
+One potential mechanism is that workers in male same-sex partnerships may transition into more in-demand occupations, thereby mitigating relative labor supply and wage gaps.
+However, as shown in Tables A2 and A3, even after accounting for occupation, men in same-sex partnerships exhibit significant changes in both labor supply and wages.
+These findings suggest that the impact of anti-discrimination laws is not merely a by-product of occupational sorting.
+Nonetheless, caution is warranted when interpreting these results, given the inherent limitations of conditioning on occupation in a broader labor force context.
+Appendix Tables
+Table A1: Effects of Anti-Discrimination Laws on Labor Supply: Analysis of Same-Sex Marriage Fixed Effects
+(1) (2) (3) (4) (5) (6)
+Extensive Margin Intensive Margin
+Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
+Panel A: Men
+ENDA $\times$ SSP 0.0267∗∗ 0.0270∗∗ 0.5377 0.8852 -0.4480 -0.3068
+(0.0129) (0.0133) (0.4982) (0.6173) (0.3629) (0.2818)
+Observations 3,798,599 3,798,599 3,798,599 3,798,599 3,073,844 3,073,844
+Mean of DV 0.8254 0.7903 35.5714 39.9865 42.4118 47.6751
+Employed Only X X
+Panel B: Women
+ENDA $\times$ SSP -0.0130 -0.0212 0.8477 -0.7081 0.9893∗∗∗ -0.7657∗∗
+(0.0108) (0.0115) (0.5163) (0.5760) (0.3081) (0.3166)
+Observations 3,982,894 3,982,894 3,982,894 3,982,894 2,691,849 2,691,849
+Mean of DV 0.8200 0.7846 34.5069 39.5937 41.2556 47.2540
+Employed Only X X
+  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
+We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment.
+The regressions are estimated separately by sex: Panel A for men, while Panel B for women.
+The coefficients indicate the impact of anti-discrimination laws.
+The analysis is divided into two parts: Columns (1-2) focus on the extensive margin, examining changes in labor force participation and employment status as a result of the laws.
+Columns (3-6), on the other hand, explore the intensive margin by leveraging variations in the number of weeks worked per year and average weekly working hours.
+We exclude $\sum_{s=1}^{S}\lambda_{s}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{SSM}_{st}\times\mathbf{1}_{\text{state}(c)=s}\bigr{]}$ from Equation (1) for this table.
+Standard errors are clustered at the county level.
+$p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
+Table A2: Anti-Discrimination Laws on Labor Supply with Occupation FEs
+(1) (2) (3) (4) (5) (6)
+Extensive Margin Intensive Margin
+Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
+Panel A: Men
+ENDA $\times$ SSP 0.0228∗∗ 0.0246∗∗ 0.5170 0.8743∗∗ -0.1462 -0.0550
+(0.0093) (0.0100) (0.3720) (0.4334) (0.3307) (0.2797)
+Observations 3,798,542 3,798,542 3,798,542 3,798,542 3,073,797 3,073,797
+Mean of DV 0.8254 0.7903 35.5714 39.9865 42.4118 47.6751
+Employed Only X X
+Panel B: Women
+ENDA $\times$ SSP -0.0064 -0.0132 0.5614 -0.8344 0.5854∗∗ -0.9524∗∗∗
+(0.0098) (0.0111) (0.4008) (0.4630) (0.2939) (0.3219)
+Observations 3,982,816 3,982,816 3,982,816 3,982,816 2,691,783 2,691,783
+Mean of DV 0.8200 0.7846 34.5069 39.5937 41.2556 47.2540
+Employed Only X X
+  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
+We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment.
+The regressions are estimated separately by sex: Panel A for men, while Panel B for women, and include four-digit occupation fixed effects.
+The coefficients indicate the impact of anti-discrimination laws.
+The analysis is divided into two parts: Columns (1-2) focus on the extensive margin, examining changes in labor force participation and employment status as a result of the laws.
+Columns (3-6), on the other hand, explore the intensive margin by leveraging variations in the number of weeks worked per year and average weekly working hours.
+Standard errors are clustered at the county level.
+$p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
+Table A3: Anti-Discrimination Laws on Pay with Occupation FEs
+Hourly Wage Annual Earnings
+(1) (2) (3) (4) (5) (6) (7) (8)
+Percentile $>$25th $>$50th $>$75th Percentile $>$25th $>$50th $>$75th
+Panel A: Men
+ENDA $\times$ SSP 1.9105∗∗ 0.0241∗∗ 0.0262 0.0239 2.2973∗∗∗ 0.0241∗∗ 0.0445∗∗∗ 0.0340∗∗
+(0.9218) (0.0113) (0.0147) (0.0147) (0.8353) (0.0113) (0.0148) (0.0132)
+Observations 3,798,542 3,798,542 3,798,542 3,798,542 3,798,542 3,798,542 3,798,542 3,798,542
+Mean of DV 52.0273 0.7917 0.5653 0.3121 52.4128 0.7917 0.5746 0.3163
+Panel B: Women
+ENDA $\times$ SSP -0.9428 -0.0023 -0.0232∗∗ -0.0096 -1.1544 -0.0023 -0.0217 -0.0106
+(0.6914) (0.0097) (0.0116) (0.0148) (0.6491) (0.0097) (0.0118) (0.0154)
+Observations 3,982,816 3,982,816 3,982,816 3,982,816 3,982,816 3,982,816 3,982,816 3,982,816
+Mean of DV 49.3832 0.8002 0.5246 0.2564 49.5588 0.8002 0.5324 0.2551
+  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
+We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment.
+The regressions are estimated separately by sex and include four-digit occupation fixed effects.
+The coefficients in the first row indicate the impact of anti-discrimination laws.
+Columns (1–4) report estimates for hourly real wage ranks, while Columns (5–8) report estimates for annual real wage ranks.
+In Columns 2–4 and 6–8, the estimates are indicator variables for being above the 25th, 50th, and 75th percentiles, respectively.
+All income variables are measured in 2019 dollars and expressed as percentile ranks.
+Standard errors are clustered at the county level.
+$p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
+Appendix Figures
+((a)) Labor Force Participation
+((b)) Employment
+Figure A1: Labor Supply Impacts on the Extensive Margin: Same-Sex Marriage Legalization
+Notes: Event study plots on the difference in labor supply between people in same-sex partnerships and different-sex partnerships broken down by sex following the county-level and state-level anti-discrimination laws.
+Sample is limited to observations before same-sex marriage legalization.
+Coefficients with 95% confidence intervals.
+Standard errors are clustered at the county level.
+((a)) Labor Force Participation
+((b)) Employment
+Figure A2: Labor Supply Impacts on the Extensive Margin: Non-Movers
+Notes: Event study plots on the difference in labor supply between people in same-sex partnerships and different-sex partnerships broken down by sex following the county-level and state-level anti-discrimination laws.
+Sample is limited to individuals who have not moved in the past year.
+Coefficients with 95% confidence intervals.
+Standard errors are clustered at the county level.
+((a)) Labor Force Participation
+((b)) Employment
+Figure A3: Labor Supply Impacts on the Extensive Margin: Metropolitan Areas
+Notes: Event study plots on the difference in labor supply between people in same-sex partnerships and different-sex partnerships broken down by sex following the county-level and state-level anti-discrimination laws.
+Sample is limited to individuals who live in metropolitan areas.
+Coefficients with 95% confidence intervals.
+Standard errors are clustered at the county level.
+((a)) Labor Force Participation
+((b)) Employment
+Figure A4: Labor Supply Impacts on the Extensive Margin: 2008-2019
+Notes: Event study plots on the difference in labor supply between people in same-sex partnerships and different-sex partnerships broken down by sex following the county-level and state-level anti-discrimination laws.
+Sample is limited to cover 2008 through 2019.
+Coefficients with 95% confidence intervals.
+Standard errors are clustered at the county level.
+((a)) Labor Force Participation
+((b)) Employment
+Figure A5: Labor Supply Impacts on the Extensive Margin: Weighting Schemes
+Notes: Event study plots on the difference in labor supply between people in same-sex partnerships and different-sex partnerships broken down by sex following the county-level and state-level anti-discrimination laws.
+We compare estimates weighted by ACS weights and without.
+Coefficients with 95% confidence intervals.
+Standard errors are clustered at the county level.
+Online Appendix*: Effects of State and Local Sexual Orientation Anti-Discrimination Laws on Labor Market Differentials
 In the original version of our paper, we employed a traditional difference-in-differences (DiD) setup, incorporating a range of controls to assess the impact of anti-discrimination legislation on labor market outcomes for LGB workers.
@@
 Furthermore, in response to recent advancements in DiD methodologies, we use Borusyak et al.
-[2021] to estimate the impact of anti-discrimination laws.
+[2021] in the main text to estimate the impact of anti-discrimination laws.
 Our analysis reveals that the application of these modern DiD techniques does not substantially alter the point estimates we initially obtained.
@@
 The polling information contains data on every state for every year, except for Alaska and Hawaii, which are missing polling information for 2005-2008.
-A.1 Identification Strategy & Estimation
-The differential roll-out of anti-discrimination laws by state and locality over time lends itself to a difference-in-differences framework with the main outcomes of interest as the change in labor supply and pay between two groups, LGB workers and heterosexual workers.
+O.4 Identification Strategy & Estimation
+The differential roll-out of anti-discrimination laws by state and locality over time lends itself to a difference-in-differences framework with the main outcomes of interest as the change in labor supply and pay between two groups, workers in same-sex versus different-sex partnerships.
 This strategy uses the variation presented in Figure 1, which shows how the laws changed over time by state and county.
 Formally, we estimate this equation on labor supply:
-$LS_{ict}=\alpha_{0}Law_{ct}+\alpha_{1}SSP_{i}+\alpha_{2}Law_{ct}\cdot SSP_{i}+\alpha_{3}X_{i}+\alpha_{4}\gamma_{st}+\mu_{c}+\delta_{t}+\epsilon_{itcs}$ (4)
+$LS_{ict}=\alpha_{0}ENDA_{ct}+\alpha_{1}SSP_{i}+\alpha_{2}ENDA_{ct}\cdot SSP_{i}+\alpha_{3}X_{i}+\alpha_{4}\gamma_{st}+\mu_{c}+\delta_{t}+\epsilon_{itcs}$ (4)
 Where $LS_{ict}$ is the labor force participation or employment status for person $i$ in county $c$ in year $t$.
-$Law_{ct}$ is an indicator for if county $c$ has an employment anti-discrimination law on sexual orientation or is covered by a state-wide anti-discrimination law in year $t$, and $SSP_{it}$ is an indicator for if person $i$ is in a same-sex partnership.
+$ENDA_{ct}$ is an indicator for if county $c$ has an employment anti-discrimination law on sexual orientation or is covered by a state-wide anti-discrimination law in year $t$, and $SSP_{it}$ is an indicator for if person $i$ is in a same-sex partnership.
 $X_{i}$ is a vector of person-specific covariates including race, children, along with age and education level - each represented up to a quadratic term to capture potential nonlinear effects.
@@
 We estimate a similar equation for wages:
-$\ln(Y_{ict}+1)=\beta_{0}\text{Law}_{ct}+\beta_{1}\text{SSP}_{i}+\beta_{2}\text{Law}_{ct}\cdot\text{SSP}_{i}+\beta_{3}X_{i}+\beta_{4}\gamma_{st}+\mu_{c}+\delta_{t}+\epsilon_{itcs}$ (5)
-Where $Y_{ict}$ is the real value, in 1999 dollars, of person $i$’s wages in county $c$ year $t$.
+$\ln(Y_{ict}+1)=\beta_{0}\text{ENDA}_{ct}+\beta_{1}\text{SSP}_{i}+\beta_{2}\text{ENDA}_{ct}\cdot\text{SSP}_{i}+\beta_{3}X_{i}+\beta_{4}\gamma_{st}+\mu_{c}+\delta_{t}+\epsilon_{itcs}$ (5)
+Where $Y_{ict}$ is the real value, in 2005 dollars, of person $i$’s wages in county $c$ year $t$.
 The rest of the notation is the same as above.
@@
 This approach ensures a comprehensive evaluation of the law’s impact on the entire income spectrum, capturing nuances missed by focusing solely on employed workers.
-We report the coefficients estimated via Borusyak et al.
-[2021] in the main text, and standard TWFE in the Appendix.
-A.2 TWFE Results
-In Table A1, we present the regression on the extensive margin of labor supply with Panel A giving the effect of anti-discrimination laws on gay men’s labor supply and Panel B giving the effect for lesbian women.
+O.5 TWFE Results
+In Table O1, we present the regression on the extensive margin of labor supply, with Panel A showing the effect of anti-discrimination laws on the labor supply of men in male same-sex partnerships, and Panel B showing the effect for women in female same-sex partnerships.
 We present the full specification in columns (1) and (2) and include a specification without accounting for local anti-discrimination laws in columns (3) and (4) to show how failing to account for local laws would lead to incorrect inference.
-Consistent with the previous literature, we find that lesbian women have a labor supply premium and gay men have a labor supply penalty on the extensive margin.
-Lesbian women are 7.3% more likely to be employed and gay men are 7.9% less likely to be employed than their straight peers.
-Anti-discrimination laws have a significant effect in reducing the labor force participation and employment gap for gay men.
+Consistent with the previous literature, we find that women in female same-sex partnerships have a labor supply premium and men in male same-sex partnerships have a labor supply penalty on the extensive margin.
+Women in female same-sex partnerships are 7.3% more likely to be employed and men in male same-sex partnerships are 7.9% less likely to be employed than their different-sex counterparts.
+Anti-discrimination laws have a significant effect in reducing the labor force participation and employment gap for men in male same-sex partnerships.
 The labor force participation rate gap is reduced by 1.3 p.p.
@@
 (18%).
-Conversely, the labor force participation and employment premium that lesbian women have over their straight counterparts is reduced by 1.3 p.p.
+Conversely, the labor force participation and employment premium that women in female same-sex partnerships have over their different-sex counterparts is reduced by 1.3 p.p.
 (16%) and 1 p.p.
 (14%), respectively.
-As seen in columns (3) and (4) of Panel A, estimating the effect of anti-discrimination laws only using state laws will lead one to erroneously conclude that these laws have no significant effect on the gay labor force and employment gap.
-However, the effect on the extensive margin of labor supply for lesbian women looks similar when estimating the effect of just state anti-discrimination laws.
-In Table A2, we present the regression results on the intensive margin of labor supply as measured by weekly hours worked and weeks worked.
+As seen in columns (3) and (4) of Panel A, estimating the effect of anti-discrimination laws using only state laws may lead to the erroneous conclusion that these laws have no significant effect on the labor force participation and employment gaps for men in male same-sex partnerships.
+However, the effect on the extensive margin of labor supply for women in female same-sex partnerships looks similar when estimating the effect of just state anti-discrimination laws.
+In Table O2, we present the regression results on the intensive margin of labor supply as measured by weekly hours worked and weeks worked.
 Similar to Table 4, we include the results when only estimating the effects of the state laws in columns (3) and (4) to show the contribution of including local laws in estimating the effect of anti-discrimination laws.
-We replicate past findings where gay men supply significantly less labor and lesbian women supply significantly more labor on the intensive margin.
-As with Table A1, we find contrasting effects of anti-discrimination laws on male and female outcomes.
-We find that the intensive labor supply gap for gay men as measured by weekly hours worked and weeks worked is significantly reduced, and the intensive labor supply premium for lesbian women is significantly reduced as well.
-The effect of anti-discrimination laws on weekly hours worked becomes insignificant for gay men and lesbian women when only estimating off of state-wide anti-discrimination laws.
-In Table A3, we present the regression results on the wage gap as measured by hourly wages and annual wage earnings.
+We replicate past findings where men in male same-sex partnerships supply significantly less labor and women in female same-sex partnerships supply significantly more labor on the intensive margin.
+As with Table O1, we find contrasting effects of anti-discrimination laws on male and female outcomes.
+We find that the intensive labor supply gap for men in male same-sex partnerships as measured by weekly hours worked and weeks worked is significantly reduced, and the intensive labor supply premium for women in female same-sex partnerships is significantly reduced as well.
+The effect of anti-discrimination laws on weekly hours worked becomes insignificant for men in male same-sex partnerships and women in female same-sex partnerships when only estimating off of state-wide anti-discrimination laws.
+In Table O3, we present the regression results on the wage gap as measured by hourly wages and annual wage earnings.
 These variables received a log-transformation, to handle zero values so the changes can be interpreted in percentage terms.
@@
 Conditioning on employed workers would create an endogeneity problem since labor force participation and employment are significantly impacted by anti-discrimination laws.
-Given Table A1, the sample of employed workers could be changing as a result of the law changes, which would lead to a change in the sample composition.
+Given Table O1, the sample of employed workers could be changing as a result of the law changes, which would lead to a change in the sample composition.
 We present the regressions limited to employed workers to replicate past work on the wage gap/premium where it is common to only examine employed workers.
@@
 We present those results in columns (5) - (8).
-Gay men see a significant decline in the hourly wage gap by 2.8 p.p.
+Men in male same-sex partnerships see a significant decline in the hourly wage gap by 2.8 p.p.
 (11%) at the 90% significance level and see an insignificant decline in the annual earnings gap by 5.8 p.p.
-On the other hand, lesbian women see a significant change in their annual earnings premium by 13 p.p.
+On the other hand, women in female same-sex partnerships see a significant change in their annual earnings premium by 13 p.p.
 (16%) and an insignificant reduction in their hourly wage premium.
-Columns (3) and (4) replicate the past literature where employed gay men make about 8% to 11% less than their straight counterparts and employed lesbian women make about 5% to 15% more than their straight counterparts.
+Columns (3) and (4) replicate the past literature where employed men in male same-sex partnerships make about 8% to 11% less than their different-sex counterparts and employed women in female same-sex partnerships make about 5% to 15% more than their different-sex counterparts.
 The pay gap/premium is reduced quite significantly once the sample is limited to employed workers, which is unsurprising.
-As shown in Table A1 and Table A2, there are significant differences in labor supply for LGB and straight workers.
-Gay men and straight women are far more likely to be unemployed, out of the labor force, and work fewer hours and weeks, so including their zero wages in the regression will increase the gap/premium.
+As shown in Table O1 and Table O2, there are significant differences in labor supply for same-sex and different-sex workers.
+Men in male same-sex partnerships and women in different-sex partnerships are far more likely to be unemployed, out of the labor force, and work fewer hours and weeks, so including their zero wages in the regression will increase the gap/premium.
 There is a significant decrease in the hourly wage gap now and an insignificant decrease in the annual wage gap.
-There is also a significant decrease in the annual earnings premium for lesbian women but no change in their hourly wage premium.
-This result is consistent with the labor supply results in Table A1 and A2 where their intensive and extensive margin of labor supply premium are declining, so the annual earnings decline, but hourly wage does not.
-A.3 Occupation Fixed Effects
+There is also a significant decrease in the annual earnings premium for women in female same-sex partnerships but no change in their hourly wage premium.
+This result is consistent with the labor supply results in Table O1 and O2 where their intensive and extensive margin of labor supply premium are declining, so the annual earnings decline, but hourly wage does not.
+O.6 Occupation Fixed Effects
 We show that even when controlling for occupation there is a significant change in relative labor supply for men and women in same-sex partnerships and a significant change in relative pay for women in same-sex partnerships, indicating there is a significant effect of these laws even within occupation.
-We present the extensive labor supply and pay regressions with occupation fixed effects in Table A4 and A5.
-In these tables, there is a smaller pay gap for gay men and a smaller pay premium for lesbian women when controlling for occupation or industry.
-However, qualitatively the results seem to be similar to the main results in Table A1 and Table A3 .
-Gay men see a significant reduction in the labor supply gap, and lesbian women see a significant reduction in their labor supply and annual wage premium.
-There is no longer a significant effect on the wage gap for gay men.
+We present the extensive labor supply and pay regressions with occupation fixed effects in Table O4 and O5.
+In these tables, there is a smaller pay gap for men in male same-sex partnerships and a smaller pay premium for women in female same-sex partnerships when controlling for occupation or industry.
+However, qualitatively the results seem to be similar to the main results in Table O1 and Table O3.
+Men in male same-sex partnerships see a significant reduction in the labor supply gap, and women in female same-sex partnerships see a significant reduction in their labor supply and annual wage premium.
+There is no longer a significant effect on the wage gap for men in male same-sex partnerships.
 These findings suggest that anti-discrimination laws have a significant effect even within occupations.
 The effect does not come from workers shifting into occupations with less discrimination, but even when conditional on occupation choice, there is a significant reduction in differences in labor supply across sexual orientations for men and women.
-Appendix Tables
-Table A1: Effect of Anti-Discrimination Laws: Extensive Margin of Labor Supply
+O.7 Sorting and Increased Reporting
+One finding from Klawitter and Flatt [1998] suggests that LGB workers sort to areas with anti-discrimination laws.
+If LGB workers were sorting to areas that recently passed anti-discrimination laws, then it could violate the identifying assumption.
+It would be concerning if high-wage LGB workers migrated from unprotected areas to areas that recently passed anti-discrimination laws since it would lower the average wage in untreated counties and increase the average wage in treated counties.
+Another potential concern would be increased reporting for same-sex partnerships.
+Same-sex partnerships, while less stigmatized in 2005-2019 than in previous years, were still heavily stigmatized.
+Anti-discrimination laws could signal to LGB people that their community is more accepting, and people may be more likely to declare that they are in a same-sex partnership.
+This change in reporting could affect the composition of our sample, which in turn could violate the identifying assumption.
+To address this concern, we run the regression below to see if there is an increase in same-sex partnerships in a given county after an anti-discrimination law is passed:
+$\text{SSP}_{ict}=\beta_{0}\text{Laws}_{ct}+\beta_{1}X_{i}+\beta_{2}\gamma_{st}+\mu_{c}+\delta_{t}+\epsilon_{itcs}$ (6)
+In this regression, $\beta_{0}$ is the coefficient of interest, and it will give the effect of anti-discrimination laws passing on the number of same-sex partnerships in a given county whether that is from in-migration or increased reporting of same-sex partnerships.
+We present these results in Table O6.
+We find no significant effect of anti-discrimination laws on the number of same-sex partnerships at the county level, suggesting there is minimal sorting or change in reporting following the passage of anti-discrimination laws.
+Online Appendix Tables
+Table O1: Effect of Anti-Discrimination Laws: Extensive Margin of Labor Supply
 Panel A: Men
@@
 Standard errors are clustered at the county level.
-$*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
-Table A2: Effect of Anti-Discrimination Laws: Intensive Margin of Labor Supply
+$p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
+Table O2: Effect of Anti-Discrimination Laws: Intensive Margin of Labor Supply
 Panel A: Men
@@
 Standard errors are clustered at the county level.
-$*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
-Table A3: Effect of Anti-Discrimination Laws: Wages and Earnings
+$p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
+Table O3: Effect of Anti-Discrimination Laws: Wages and Earnings
 Panel A: Men
@@
 Columns (3) and (4) are limited to those who are employed.
-All of the are in terms of 1999 dollars and received the log(x+1) transformation to have the interpretation of the coefficient be in terms of percent.
+All of the are in terms of 2019 dollars and received the log(x+1) transformation to have the interpretation of the coefficient be in terms of percent.
 The first row of coefficients show the effect of anti-discrimination laws on the pay gap or premium, and the second row of coefficients give the pay gap or premium.
@@
 Standard errors are clustered at the county level.
-$*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
-Table A4: Anti-Discrimination Laws on Labor Supply with Occupation FEs
+$p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
+Table O4: Anti-Discrimination Laws on Labor Supply with Occupation FEs
 Men Women
@@
 Standard errors are clustered at the county level.
-$*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
-Table A5: Anti-Discrimination Laws on Pay with Occupation FEs
+$p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
+Table O5: Anti-Discrimination Laws on Pay with Occupation FEs
 Panel A: Men
@@
   • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex and include four-digit occupation fixed effects.
-All of the are in terms of 1999 dollars and received the log(x+1) transformation to have the interpretation of the coefficient be in terms of percent.
+All of the are in terms of 2019 dollars and received the log(x+1) transformation to have the interpretation of the coefficient be in terms of percent.
 The first row of coefficients show the effect of anti-discrimination on the pay gap or premium, and the second row of coefficients give the pay gap or premium.
 Standard errors are clustered at the county level.
-$*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
-Table A6: Sorting and Reporting
-(1) (2) (3)
-All SSP Male SSP Female SSP
-Laws 0.000170 0.000633 -0.000304
-(0.000489) (0.000572) (0.000573)
-Observations 12,872,572 6,295,028 6,577,544
-R-squared 0.068 0.072 0.065
-  • Notes: Data comes from the 2005-2016 yearly ACS seeing how the number of same-sex partnerships change in a county after the passage of an anti-discrimination law.
+$p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
+Table O6: Sorting and Reporting
+(1) (2) (3) (4) (5)
+All SSP Male SSP Female SSP Age Years of Education
+ENDA -0.0002 0.0004 -0.0008 0.0257 -0.0212
+(0.0005) (0.0005) (0.0005) (0.0648) (0.0182)
+Observations 16,071,420 7,850,821 8,220,599 16,071,420 16,071,420
+Mean of DV 0.0122 0.0123 0.0122 45.5220 13.8539
+$R^{2}$ 0.0043 0.0081 0.0024 0.0101 0.0467
+  • Notes: Data comes from the 2005-2019 yearly ACS seeing how the number of same-sex partnerships change in a county after the passage of an anti-discrimination law.
 The first column looks at both men and women with the next two columns separating the sexes.
+Column (4) investigates age and (5) looks at years of education.
 Standard errors are clustered at the county level.
-$*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
-Table A7: Differential Responses between Male and Female Same-Sex Partnerships
+$p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
+Table O7: Differential Responses between Male and Female Same-Sex Partnerships
 (1) (2) (3) (4)
@@
 Standard errors are clustered at the county level.
-$*{}^{*}$ $p<0.10$, ${}^{**}$ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
+$p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.

### v3

@@
 Scott Delhommer University of Texas at Austin; sdelhommer@utexas.edu    Domonkos F. Vamossy University of Pittsburgh, d.vamossy@pitt.edu
-(April 1, 2025)
-This paper presents the first quasi-experimental research examining the effect of both local and state anti-discrimination laws on sexual orientation on the labor supply and wages of lesbian, gay, and bisexual (LGB) workers.
+(November 19, 2025)
+This paper presents quasi-experimental research examining the effect of both local and state anti-discrimination laws on sexual orientation on the labor supply and wages of lesbian, gay, and bisexual (LGB) workers.
 To do so, we use the American Community Survey data on household composition to infer sexual orientation and combine this with a unique panel dataset on state and local anti-discrimination laws.
-Leveraging variation in law implementation across localities over time and between same-sex and different-sex couples, we find that anti-discrimination laws not only significantly narrow gaps in labor force participation and employment for male same-sex couples relative to men in different-sex couples, but also boost their percentile rank in the wage distribution.
+Leveraging variation in law implementation across localities over time and between same-sex and different-sex couples, we find that anti-discrimination laws significantly narrow gaps in labor force participation and employment for men in same-sex couples relative to men in different-sex couples, and also increase their percentile rank in the wage distribution.
 Our analysis reveals mostly null effects for female same-sex couples; however, in metropolitan areas these laws significantly reduce their employment compared to women in different-sex couples.
@@
 1 Introduction
-On April 15, 2020, the Supreme Court of the United States ruled that federal employment discrimination law applied to lesbian, gay, bisexual, and transgender (LGBT) individuals.
-Before this ruling, the legislation in 25 states made it legal for private businesses to fire an employee for being LGBT.
-Unlike race, sex, age, and religion, federal law did not include sexual orientation as a protected class for federal anti-discrimination laws until 2020.
-As a result, laws extending protected class status based on sexual orientation varied between states and localities.
-The first state to pass an anti-discrimination law based on sexual orientation was Wisconsin in 1982, and the most recent state to pass such a law was Pennsylvania and Michigan in 2023.111Since August 2018, Pennsylvania has interpreted discrimination based on sexual orientation and gender identity as being covered under the “sex” category of the Pennsylvania Human Relations Act.
-This allowed LGBT individuals to file complaints for discrimination in employment, housing, education, and public accommodations.
-Following Michigan’s lead in May 2018, Pennsylvania became the second state to implement statewide LGBT protections in this manner.
-There has been a renewed push to enact sexual orientation protections at the legislative level as Democrats in the House of Representatives introduced “The Equality Act” in early 2019, which would add sexual orientation and gender identity as federally protected characteristics.
-The focus of this paper is the effect of local and state sexual orientation anti-discrimination legislation on labor market differentials for lesbian, gay, and bisexual (LGB) workers.
-[^2] For a review of the impacts of legal access to same-sex marriage, see Badgett et al.
-(2024).
-For a broader review of the economics related to sexual orientation and gender identity, see Badgett et al.
-(2023).
-We exploit the differential roll-out of state and local laws from 2005-2019 in a modern triple difference-in-differences framework to analyze how these anti-discrimination laws impact the wages and labor supply of LGB workers.
-The economics literature on LGB workers consistently finds that gay/bisexual men have a pay and labor supply penalty and lesbian/bisexual women have a pay and labor supply premium over their heterosexual counterparts333(Badgett (1995); Klawitter (2011); Klawitter and Flatt (1998); Allegretto and Arthur (2001); Black et al.
-(2003); Carpenter (2005); Black et al.
-(2007); Antecol et al.
-(2008); Klawitter (2015); Jepsen and Jepsen (2017); Carpenter and Eppink (2017)), Tilcsik et al.
-(2015).
+On June 15, 2020, the U.S.
+Supreme Court held in Bostock v.
+Clayton County, 590 U.S.
+644 (2020), that Title VII’s prohibition of sex discrimination in employment includes discrimination based on sexual orientation and gender identity.
+Prior to this ruling, more than half of U.S.
+states lacked explicit statewide employment protections based on sexual orientation, though many localities had adopted their own ordinances.
+As a result, protections varied substantially across and within states throughout the period studied.
+Wisconsin became the first state to enact a statewide sexual orientation employment non-discrimination law in 1982.
+There has been continued legislative interest in expanding federal protections, most notably the introduction of the “Equality Act” in early 2019, which would add sexual orientation and gender identity to federal civil rights statutes.
+This variation in state and local protections provides the basis for our empirical analysis.
+In this paper, we examine how sexual orientation employment non-discrimination laws affect labor market differentials for lesbian and gay workers.
+We exploit the differential roll-out of state and local laws from 2005–2019 in a modern difference-in-differences-in-differences (DDD) framework to analyze how these anti-discrimination laws impact the wages and labor supply of LGB workers.
+The economics literature consistently finds that gay men (or men in male same-sex couples) face earnings and labor-supply penalties, while lesbian women (or women in female same-sex couples) exhibit earnings and labor-supply premia relative to comparable heterosexual counterparts.
+[^1] For a review of the impacts of legal access to same-sex marriage, see badgett2024review.
+For a broader review of the economics related to sexual orientation and gender identity, see badgett2023review.
+In addition, see badgett1995wage; klawitter1998effects; allegretto2001empirical; black2003earnings; carpenter2005self; antecol2008sexual; klawitter2011multilevel; klawitter2015meta; jepsen2017self; carpenter2017does; Burn2018.
 The differences in pay are attributed to a host of factors including discrimination and intra-household labor allocation and specialization.
+[^2] Occupational sorting and concealment may also mediate gaps, with sexual minorities selecting into jobs with lower customer discrimination or greater autonomy (tilcsik2015concealable; black2007economics).
+Consistent with administrative complaints, sexual-orientation discrimination occurs at rates comparable to sex-based discrimination (ramos2008evidence).
 We test the role of discrimination by examining how the passage of anti-discrimination laws affects the labor supply and wage differences between LGB and heterosexual workers.
-There have been previous studies examining how these state and local laws affect the pay gap, but these studies have been limited in their ability to identify causal effects (Klawitter and Flatt (1998); Gates (2009)).
-Klawitter and Flatt (1998) compared people in same-sex partnerships to those in different-sex partnerships in the 1990 Census and found that same-sex couples have no significant difference in wages in places with anti-discrimination laws.
-Gates (2009) uses the 2000 Census and conducts a similar analysis.
-Gates (2009) finds that gay/bisexual men in places with anti-discrimination laws have a 3% wage premium over gay/bisexual men in places without these laws, and lesbian/bisexual women have a 2% wage premium over lesbian/bisexual women in places without these laws [^4] Tilcsik (2011) found that resumes signaling LGB status received significantly fewer callbacks in localities without sexual orientation anti-discrimination laws.
-These past studies on anti-discrimination laws, while informative, fail to fully account for changes over time.
-Previous research has leveraged state-level sexual orientation anti-discrimination laws within a difference-in-differences framework.
-[^5] The Williams Institute found that sexual orientation discrimination occurs at a similar rate to sex-based discrimination using state-level complaint data following state anti-discrimination laws (Ramos et al., 2008).
-Martell (2013) study demonstrated that such laws have significantly reduced wage disparities for gay men by 20%.666 Since Martell (2013) and Klawitter and Flatt (1998) there has been a booming literature on difference-in-differences.
-Accordingly, unlike Martell (2013) and Klawitter and Flatt (1998), we employ modern difference-in-difference estimators to alleviate biases stemming from staggered policy design.
-Our research extends beyond the scope of previous studies by incorporating an examination of local laws, thereby offering a more comprehensive understanding of the legislative landscape.
-Furthermore, this study also estimates the effects of such laws on women, a demographic previously under-explored in this context.
-Our study expands the scope of Klawitter and Flatt (1998) by leveraging a larger dataset, utilizing American Community Survey (ACS) data spanning from 2005 to 2019, in contrast to Klawitter and Flatt (1998) which relied on the 1990 U.S.
-census data encompassing approximately 13,000 married couples, 14,500 different-sex unmarried couples, and 6,800 same-sex couples.
-The ACS data allows for a robust analysis across a broader temporal spectrum, capturing a wide array of demographic and economic changes over 15 years.
-This extensive dataset, combined with the application of modern difference-in-difference estimators, enables a more detailed and dynamic examination of the effects of anti-discrimination policies on the earnings and household incomes of same-sex couples.
-Additionally, Burn (2018)’s more recent examination in a triple difference-in-differences framework confirms these findings but also underscores a gap in the literature: the lack of consideration for local anti-discrimination ordinances.
-Previous studies have not fully addressed potential biases arising from variations in the timing of these laws’ enactment.
-This paper is the first to examine both local and state-level sexual orientation anti-discrimination laws in a quasi-experimental design.
-This paper is also the first to analyze the effects of any local anti-discrimination laws on sex, race, or sexual orientation in a modern difference-in-difference framework [^7] Using state variation in anti-discrimination laws has also been analyzed to understand racial- and sex-based discrimination (Neumark and Stock (2006); (Donohue III and Heckman, 1991);((Margo, 1995); Goldin and Margo (1992)).
-However, these studies on sex and race anti-discrimination laws also omit local laws, potentially skewing their estimates..
+This paper provides quasi-experimental evidence on both local and state sexual orientation anti-discrimination laws.
+It is also, to our knowledge, the first study to evaluate the effects of local anti-discrimination ordinances, whether based on sex, race, or sexual orientation, within a modern difference-in-differences framework.
+[^3] Legal changes and enforcement have been used to study race- and sex-based discrimination (e.g., neumark2006labor; see also donohue1991continuous).
+These analyses do not incorporate local ordinances, which can bias estimates that rely on state-only policy variation.
 We use the 2005-2019 American Community Survey (ACS) and household composition to infer sexual orientation and create a unique and novel panel dataset on the passage of local and state anti-discrimination laws.
@@
 We find a significant reduction in differences between LGB workers and heterosexual workers across labor supply and wage measures, due to anti-discrimination laws.
-Anti-discrimination laws significantly reduce the gap in labor force participation and employment of men in male same-sex couples by 3.8 percentage points (78%) and 3.7 percentage points (77%), respectively.
-The laws improve men in male same-sex couples relative position in the wage distribution, with their percentile rank in the hourly wage and annual earnings distributions improving on average by approximately 2.7 and 3 percentiles, respectively.
-By contrast, the wage distribution for women in same-sex partnerships remains largely unaffected by these laws.
-Using an event study plot, we show that the outcome trends are parallel before the implementation of the anti-discrimination laws suggesting the workers in treatment and control regions are reasonable comparisons.
-[^8] In Appendix B A, we replicate the findings of past research that finds men in male same-sex couples have lower labor force participation and employment rates and make 8-11% less than their employed straight counterparts using hourly wages and annual earnings.
-We also replicate findings that women in female same-sex couples have a higher labor force participation and employment rate and earn 5-15% more than their employed straight counterparts.
-We explore theories for the differing effects of anti-discrimination laws on men in male same-sex couples and women in female same-sex couples in the Discussion section.
-Reduced discrimination improves job security for sexual minorities (Badgett, 2006), which may allow the secondary earner to reduce market participation in favor of increased household production, including child rearing, a pattern consistent with household bargaining models (Chiappori, 1992; Lundberg and Pollak, 1993).
-These differences in household specialization are less pronounced in same-sex partnerships, but they may become more similar for women in female same-sex couples following the passage of an anti-discrimination law if it gives greater protection to the higher wage earner.
+Anti-discrimination laws reduce the labor force participation and employment gaps for men in male same-sex couples by 3.8 percentage points (78% of the baseline gap) and 3.7 percentage points (77%), respectively.
+The laws improve men in male same-sex couples’ relative position in the wage distribution: their percentile ranks in the hourly wage and annual earnings distributions rise by about 2.6 and 2.9 percentile points, respectively.
+By contrast, the wage distribution for women in same-sex couples remains largely unaffected by these laws.
+Event-study estimates show no pre-trends, supporting the parallel-trends assumption and the comparability of treated and control regions.
+We explore potential mechanisms for the differing effects on men and women in the Discussion section, without imposing a single unified explanation; as in the broader literature, some differences remain an empirical puzzle.
+Reduced discrimination improves job security for sexual minorities (badgett2006discrimination), which may allow the secondary earner to reduce market participation in favor of increased household production, including child rearing, a pattern consistent with household bargaining models (chiappori1992collective; lundberg1993separate).
+These differences in household specialization are less pronounced in same-sex couples, but they may become more similar for women in female same-sex couples following the passage of an anti-discrimination law if it gives greater protection to the higher wage earner.
 We show empirically that female same-sex households have significantly more children than male same-sex households after the passage of an anti-discrimination law.
@@
 Alternatively, the anti-discrimination laws may normalize being a sexual minority, and improve sentiments toward LGB workers.
-We collected every poll on same-sex marriage for each state and year from the Pew Research Center.
-We show that the passage of state-wide sexual orientation anti-discrimination laws persistently increases favorability toward LGB people through increased support for same-sex marriage.
+We compiled state–year polls on same-sex marriage from the Pew Research Center.
+We show that the passage of statewide sexual orientation anti-discrimination laws persistently increases favorability toward LGB people through increased support for same-sex marriage.
 The states that passed anti-discrimination laws had parallel pre-trends in support of same-sex marriage before the passage of the laws.
@@
 The observed increase in support following the enactment of the law, rather than before, implies that policy changes can indeed influence public opinion.
-This notion aligns with the findings of Sansone (2019), suggesting that legislative actions can lead to shifts in societal attitudes.
-This is in contrast to the thermostatic model proposed by Wlezien (1995), which posits that public sentiment drives policy changes rather than the other way around.
+This matches the findings of sansone2019pink, suggesting that legislative actions can lead to shifts in societal attitudes.
+This is in contrast to the thermostatic model proposed by wlezien1995public, which posits that public sentiment drives policy changes rather than the other way around.
 Increased favorability toward LGB people may act in conjunction with greater job security to improve labor market outcomes for LGB Americans.
+Our study complements and extends prior state-focused analyses (e.g., klawitter2011multilevel; martell2013endas; Burn2018) by introducing a novel city–county ENDA panel and applying modern staggered-adoption estimators.
+Unlike work centered on marriage laws (e.g., hansen2020labor; hansen2022tolerance), we isolate employment protections and link them to labor supply, wage ranks, fertility, and opinion dynamics.
+1.1 Related Literature
+Research on sexual-orientation anti-discrimination laws (ENDAs) spans cross-sectional, causal state-level, and related policy literatures.
+Early cross-sectional work linked protections to narrower wage gaps but could not establish causality: klawitter1998effects use the 1990 Census and find no significant wage differences in protected places, while gates2009impact use the 2000 Census and report small premia.
+tilcsik2011pride found that resumes signaling LGB status received significantly fewer callbacks in localities without sexual orientation anti-discrimination laws.
+klawitter2011multilevel estimate multi-level earnings models with 2000 Census data, documenting systematic associations between state/local policies and earnings but not exploiting within-place policy changes over time.
+These past studies on anti-discrimination laws, while informative, fail to fully account for changes over time.
+Quasi-experimental state-level studies strengthen identification.
+martell2013endas implement a state DiD design and find sizable wage convergence for gay men post-ENDAs, and Burn2018 adopt a triple-difference framework, still at the state level.
+A limitation of these state-only designs is that local ordinances, the likely first binding margin for many workers, are omitted; moreover, conventional TWFE estimators can be biased under staggered adoption and heterogeneous effects.
+To highlight these issues, Appendix B replicates a state-only specification and shows how excluding local ordinances can mask meaningful within-state variation, underscoring the value of incorporating local laws into the analysis.
+Two adjacent strands inform mechanisms.
+First, policy shocks around marriage legalization affect household specialization and fertility: hansen2020labor show changes in labor supply following marriage legalization; hansen2022tolerance link tolerance to paid work among cohabiting gays and lesbians.
+These patterns match our household-reallocation results for women and the fertility margin we study.
+Second, mann2025antidiscrimination document mental-health benefits from anti-discrimination laws, suggesting channels beyond pecuniary outcomes and complementing our evidence on opinion dynamics.
+Broader anti-discrimination evidence underscores that legal protections and enforcement can shift hiring and employment margins across protected classes: disability (audit evidence in ArmourButtonHollands2018; policy evidence in Button2018Expanding) and age (field-experiment evidence in NeumarkBurnButtonChehras2019).
+While these are not sexual-orientation ENDAs, they corroborate that statutory protections and their salience can meaningfully change labor-market opportunities.
 2 Data
-A common issue in analyzing the pay disparities between homosexual and heterosexual workers is a lack of high-quality data about sexual orientation, wages, and employment.
-We follow the literature (Klawitter and Flatt (1998); Gates (2009); Jepsen and Jepsen (2017)) in inferring sexual orientation by looking at household composition.
-Specifically, we infer a person’s sexual orientation to be homosexual/bisexual if they have an “unmarried partner” or a spouse that is the same sex as themselves.
-[^9] We exclude individuals whose sex or relationship status was allocated following (Gates and Steinberger, 2009)..
-The ACS references a close personal relationship for unmarried partners as well as giving an option for “partner/roommate”, which reduces the likelihood that different-sex roommates would misclassify as unmarried partners.
-The comparisons that we make are between people in same-sex partnerships (SSP), defined as being in an unmarried partnership or married with someone of the same sex, and those in different-sex partnerships (DSP).
-This comparison based on household composition, though standard in the literature, is not equivalent to comparing LGB and heterosexual workers.
-It compares those who are in a same-sex relationship to those who are in a different-sex relationship.
-Importantly, this comparison excludes all workers who are not in a cohabitating relationship and classifies bisexual individuals in different-sex relationships as members of the control group.
-Because bisexual workers consistently earn less than both heterosexual and gay or lesbian counterparts ((Mize, 2016)), this classification introduces a bias toward underestimating pay disparities.
-Additionally, similar misclassification may occur if gay or lesbian individuals are categorized as heterosexual.
-Thus, the estimated pay gaps presented here likely represent conservative estimates.
-We use the 2005-2019 yearly ACS from (Ruggles et al., 2019).
+A common challenge in analyzing pay disparities between different types of same-sex and different-sex couples is the limited availability of high-quality data on sexual orientation, wages, and employment.
+We follow the literature in inferring sexual orientation by examining household relationship structure in the ACS (klawitter1998effects; gates2009impact; jepsen2017self).
+Specifically, we classify a person as being in a same-sex couple if they report a spouse or “unmarried partner” who is the same sex as themselves.
+[^4] The ACS measures self-reported “sex” rather than gender identity.
+Following the existing literature, we interpret this variable as a proxy for gender in couple-type inference, but this may misclassify some transgender or gender-nonconforming respondents.
+See, for example, gates2009impact and jepsen2017self.
+We exclude individuals whose sex or relationship status was allocated following gates2009same.
+To identify “unmarried partners,” the ACS asks each household member to report their relationship to the householder from a fixed list of categories.
+The category “unmarried partner” is defined by the Census Bureau as a person who has a close personal relationship with the householder.
+Importantly, the questionnaire provides a separate category for “housemate/roommate,” which reduces the likelihood that unrelated adults, such as different-sex roommates, are erroneously classified as unmarried partners.
+We use the 2005-2019 yearly ACS from ipumsusa.
 We use wages, defined as salaried wages from an employer.
 The ACS reports wage earnings in annual terms.
-We use the annual figures and convert them to hourly wage figures using variables on average weekly hours worked and weeks worked101010“Weeks worked” in the ACS is a categorical variable giving a range of weeks worked.
-We take the median value given in the range to compute weeks worked..
+We use the annual figures and convert them to hourly wage figures using variables on average weekly hours worked and weeks worked.555“Weeks worked” in the ACS is a categorical variable giving a range of weeks worked.
+We take the median value given in the range to compute weeks worked.
 Finally, we limit our sample to prime-age working adults and only examine those who are ages 25-65.
-We use 25 as the lower age cutoff to allow workers to complete college and have more fully developed human capital, which is typically thought of as an important factor for the differences in wages between LGB and straight workers (Black et al., 2007).
+We use 25 as the lower age cutoff to allow workers to complete college and have more fully developed human capital, which is typically thought of as an important factor for the differences in wages between LGB and straight workers (black2007economics).
 We obtained data on the passage of state anti-discrimination laws from LGBTMap.org, an LGBT advocacy group.
@@
 We completed the dataset for the relevant years using old media reports, correspondence with local officials, and FOIA requests.
-Building on earlier city-level analyses such as (Klawitter, 2011), this dataset provides the first complete city-level dataset covering 15 years of sexual orientation anti-discrimination laws in the U.S.
+Building on earlier city-level analyses such as klawitter2011multilevel, this dataset provides the first complete city-level dataset covering 15 years of sexual orientation anti-discrimination laws in the U.S.
 We focus on cities reported in the ACS and match those city laws with their corresponding counties to merge with the ACS.
 The ACS only reports county of residence for those in metro areas, so any rural counties are lumped together within a state.
-We report the state, city, and county laws and the year they were enacted in Table 1.
+We report the state, city, and county laws and the year they were enacted in Table 1.666Since August 2018, Pennsylvania has interpreted discrimination based on sexual orientation and gender identity as covered under the “sex” category of the Pennsylvania Human Relations Act, allowing LGBT individuals to file complaints for discrimination in employment, housing, education, and public accommodations.
 In Figure 2, we show which counties had sexual orientation anti-discrimination laws in 2005 and 2019, respectively.
-In our analysis, laws are coded based on their initial application to private employment and housing, focusing on broad protections rather than sector-specific ones.
-For instance, while Delaware extended protections to state employees in 1999, our coding reflects 2009, the year broader protections were enacted.
-This approach could overlook the incremental impact of earlier, narrower protections on societal attitudes.
-The same principle applies to local laws, where initial protections for public employees in some cities might precede more comprehensive legislation, potentially influencing early shifts in public sentiment towards sexual minorities.
-The enforcement mechanisms of anti-discrimination laws, such as fines, mandatory training, and public rebukes, serve as significant deterrents, encouraging the adoption of more inclusive employment policies.
-The rise in successful legal challenges against companies concerning discrimination highlights the real legal and financial consequences they encounter, reinforcing the laws’ effectiveness in promoting workplace equality (for Human Resource Management (2023)).
 The enactment of anti-discrimination legislation is not random.
-The states that have these protections are generally considered friendlier to LGB workers than those without these laws and have a larger share of same-sex partnerships than those states without these laws.
+The states that have these protections are generally considered friendlier to LGB workers than those without these laws and have a larger share of same-sex couples than those states without these laws.
 Also, many counties with anti-discrimination laws have large cities with a larger concentration of LGB workers than rural counties.
 However, this is not universal.
-For instance, Utah extended protection to LGB workers despite being a relatively conservative state, and certain liberal cities like Houston [^11] Houston passed a sexual orientation and gender identity anti-discrimination ordinance, but it was only in effect for 3 months before being challenged.
+For instance, Utah extended protection to LGB workers despite being a relatively conservative state, and certain liberal cities like Houston are noticeably absent from extending protection to LGB workers.
+[^7] Houston passed a sexual orientation and gender identity anti-discrimination ordinance, but it was only in effect for 3 months before being challenged.
 The law was put up to a public vote and lost, repealing the law.
-are noticeably absent from extending protection to LGB workers.
-One potential concern in this analysis is the endogenous adoption of laws.
-The areas that are friendly and less hostile to LGB workers may be the areas that are most likely to adopt sexual orientation anti-discrimination laws.
-The main threat to identification is that the timing of anti-discrimination laws is correlated with another factor causing an omitted variable bias.
-It could be that positive sentiment toward LGB workers causes both the law change and any change in labor market outcomes.
-We mitigate this by adopting a saturated triple-difference model, including all two-way fixed effects (county-year, year-SSP, and county-SSP), as proposed by (Burn, 2018).
-The inclusion of county-year fixed effects remove the need for county controls, addressing the lack of local support data for same-sex marriage.
-We refine the approach of (Burn, 2018) by employing the methodology of (Borusyak et al., 2021), designed for analyzing staggered policy adoptions like ours.
-This method accurately estimates the effects of anti-discrimination laws, considering their staggered implementation times.
-In Figure 2, we present the distribution of same-sex partnerships by state from 2005 to 2019, and in Table 2, we present the 10 counties with the largest share of same-sex partnerships.
-Unsurprisingly, the distribution of same-sex partnerships by state and by county is skewed toward more progressive states and counties with large cities in them that are known for having a large LGB population like San Francisco, the District of Columbia, New York City, and Boston.
-The LGB population varies from the heterosexual population in many dimensions including geography.
+In Figure 2, we present the distribution of same-sex couples by state from 2005 to 2019, and in Table 2, we present the 10 counties with the largest share of same-sex couples.
+Unsurprisingly, the distribution of same-sex couples by state and by county is skewed toward more progressive states and counties with large cities in them that are known for having a large LGB population like San Francisco, the District of Columbia, New York City, and Boston.
 We present descriptive statistics showing the differences in the same-sex couple population compared to the different-sex couple population broken down by education level and sex in Table 3.
-Descriptively, there is a clear difference in labor market outcomes and characteristics between men and women in same-sex partnerships and those in different-sex partnerships when controlling for education.
+Descriptively, there is a clear difference in labor market outcomes and characteristics between men and women in same-sex couples and those in different-sex couples when controlling for education.
+Table 1: Timing of Sexual Orientation Anti-discrimination Laws
+Year State City or County
+2005 and Before CA, CT, Boulder, CO; Denver, CO; Fort Collins, CO; Gainsville, FL; Hialeah, FL; Hollywood, FL;
+DC, HI, Key West, FL; Miami, FL; Orlando, FL; Pembroke, FL; Saint Petersburg, FL;
+ME, MD, Tampa, FL; West Palm, FL; Atlanta, GA; Ames, IA; Cedar Rapids, IA; Davenport, IA;
+MA, MN, Des Moines, IA; Iowa City, IA; Campaigne, IL; Chicago, IL; Peoria, IL; Urbana, IL;
+NV, NH, Bloomington, IN; Fort Wayne, IN; Michigan City, IN; Terre Haute, IN; Lawerence, KS;
+NJ, NM, Covington, KY; Lexington, KY; Louisville, KY; New Orleans, LA; Ann Arbor, MI;
+NY, RI, Detroit, MI; Grand Rapids, MI; Ypsilanti, MI; Columbia, MO; Kansas City, MO;
+VT, WI Saint Louis, MO; Cleveland, OH; Columbus, OH; Toledo, OH; Eugene, OR;
+Portland, OR; Benton County, OR; Salem, OR; Allentown, PA; Erie, PA; Harrisburg, PA;
+Lancaster, PA; Philadelphia, PA; Pittsburgh; Scranton, PA; Austin, TX; Dallas, TX;
+Fort Worth, TX; Alexandria, VA; Arlington, VA; Seattle, WA; Spokane, WA; Tacoma, WA
+2006 IL, WA Dubuque, IA; Indianapolis, IN, Ferndale, MI; Lansing, MI;
+Cincinnati, OH; Easton, PA; West Chester, PA; Charleston, SC
+2007 CO, IA, OR Waterloo, IA; Coshocton, OH; Dayton, OH; Newark, OH; Charleston, WV
+2008 Columbia, SC
+2009 DE Allegheny, PA; Reading, PA; Salt Lake City, UT
+2010 Tallahassee, FL; Traverse City, MI; Missoula, MT;
+Lower Merion, PA; Grand County, UT; Summit County, UT
+2011 Volusia County, FL; Evansville, IN; University City, MO; East Cleveland, OH,
+Bethlehem, PA; Conshohocken, PA; Haverford, PA; Ogden, UT
+2012 St. Augustine, FL; Boise, ID; New Albany, IN; South Bend, IN; Flint, MI;
+Muskegon, MI; Maplewood, MO; Helena, MT; Omaha, NE; Canton, OH;
+Abington, PA; Cheltenham, PA; Morgantown, WV
+2013 Phoenix, AZ; Pocatello, ID; Frankfort, KY; Shreveport, LA; Battle Creek, MI;
+Bristol, PA; Pittston, PA; San Antonio, TX; Charlottesville, VA; Huntington, WV
+2014 Tempe, AZ; Adrian, MI; Macomb County, MI; Butte, MT; Atlantic Beach, FL
+2015 UT Anchorage, AK; Osceola County, FL; Anderson, IN; Clinton, IN; Hammond, IN; Muncie, IN
+2016 Kokomo, IN; Manahattan, KS; St. Charles, MO; Jackson, MS; Lakewood, OH;
+Carlisle, PA; Dickson City, PA; Wilkes-Barre, PA; Martinsburg, WV; Wheeling, WV
+2017 Sitka, AK; Glendale, AZ;
+2018
+2019 Fort Lauderdale, FL; Fernandina Beach, FL; Decatur, GA
+2020 VA
+  • Notes: List of states, cities, and counties with sexual orientation anti-discrimination laws pulled from LGBTMap.org, an advocacy group, as well as through media reports and local FOIA requests.
+We only list city or county laws if there is no state law.
+Table 2: Counties with Largest LGB Populations
+County State Percent of SSPs
+San Francisco County CA 7.37
+District of Columbia DC 7.20
+New York County NY 5.43
+Suffolk County MA 4.37
+St. Louis city MO 4.01
+Multnomah County OR 3.99
+Alexandria city VA 3.70
+Santa Fe County NM 3.43
+Baltimore city MD 3.37
+  • Notes: Using ACS and person weights to recover the percentage of partnerships that are same-sex partnerships by county over 2005-2019.
+Table 3: Summary Statistics
+Panel A: Men
+High School Grad or Lower Some College or Higher
+SSP DSP SSP DSP
+n = 21,737 n = 2,805,327 n = 80,456 n = 4,943,301
+Variable Mean Mean Difference Mean Mean Difference
+In Labor Force 0.738 0.829 -0.091∗∗∗ 0.863 0.898 -0.035∗∗∗
+(0.440) (0.377) (0.344) (0.303)
+Employed 0.687 0.783 -0.096∗∗∗ 0.831 0.871 -0.041∗∗∗
+(0.464) (0.412) (0.375) (0.335)
+Real Wage 28,991.2 36,602.8 -7,611.6∗∗∗ 71,867.9 79,338.7 -7470.8∗∗∗
+(40,363.5) (39,963.0) (89,316.6) (90,257.4)
+Real Hourly Wage 16.735 18.661 -1.925∗∗∗ 35.718 37.691 -1.974∗∗∗
+(58.837) (44.692) (76.987) (115.923)
+Age 46.074 47.440 -1.367∗∗∗ 45.609 46.786 -1.177∗∗∗
+(10.932) (10.912) (10.645) (11.071)
+Number of Children 0.484 1.187 -0.702∗∗∗ 0.234 1.124 -0.890∗∗∗
+(1.014) (1.280) (0.692) (1.187)
+Asian 0.029 0.029 -0.000 0.046 0.062 -0.017∗∗∗
+(0.169) (0.169) (0.209) (0.242)
+Black 0.079 0.080 -0.001 0.041 0.057 -0.016∗∗∗
+(0.269) (0.271) (0.198) (0.232)
+Hispanic 0.115 0.118 -0.003 0.069 0.050 0.020∗∗∗
+(0.319) (0.323) (0.254) (0.218)
+White 0.789 0.798 -0.009∗∗ 0.863 0.842 0.021∗∗∗
+(0.408) (0.401) (0.344) (0.365)
+Panel B: Women
+High School Grad or Lower Some College or Higher
+SSP DSP SSP DSP
+n = 22,238 n = 2,654,209 n = 79,414 n = 5,464,738
+Variable Mean Mean Difference Mean Mean Difference
+In Labor Force 0.724 0.596 0.128∗∗∗ 0.862 0.749 0.114∗∗∗
+(0.447) (0.491) (0.345) (0.434)
+Employed 0.673 0.558 0.115∗∗∗ 0.833 0.722 0.111∗∗∗
+(0.469) (0.497) (0.373) (0.448)
+Real Wage 25,032.0 16,555.5 8,476.5∗∗∗ 56,374.5 39,254.0 17,120.6∗∗∗
+(33,315.4) (24,328.6) (64,852.8) (50,889.8)
+Real Hourly Wage 14.331 10.497 3.834∗∗∗ 29.083 22.721 6.363∗∗∗
+(37.390) (39.943) (62.691) (86.710)
+Age 45.437 48.151 -2.714∗∗∗ 44.860 45.382 -0.522∗∗∗
+(11.356) (10.973) (11.001) (11.134)
+Number of Children 0.716 1.088 -0.371∗∗∗ 0.497 1.103 -0.606∗∗∗
+(1.093) (1.287) (0.902) (1.179)
+Asian 0.024 0.045 -0.021∗∗∗ 0.029 0.067 -0.038∗∗∗
+(0.152) (0.206) (0.168) (0.250)
+Black 0.107 0.065 0.042∗∗∗ 0.057 0.056 0.001
+(0.309) (0.247) (0.231) (0.230)
+Hispanic 0.099 0.120 -0.021∗∗∗ 0.058 0.054 0.004∗∗∗
+(0.299) (0.325) (0.234) (0.226)
+White 0.778 0.799 -0.022∗∗∗ 0.861 0.836 0.025∗∗∗
+(0.416) (0.400) (0.346) (0.370)
+  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
+Summary statistics are presented by education level.
+T-tests were conducted to determine significant differences between those in same-sex partnerships and different-sex partnerships.
+$p<0.10$, ∗∗ $p<0.05$, $^{{}^{***}}$ $p<0.01$.
+((a)) 2005
+((b)) 2019
+Figure 1: Sexual Orientation and Anti-Discrimination Laws
+Notes: State and local sexual orientation employment anti-discrimination laws in 2005 (a) and 2019 (b).
+Data on laws obtained from LGBTMap.org and authors’ own investigation using media reports and FOIA requests.
+Figure 2: Percentage of Same Sex Couples: 2005–2019
+Notes: Percentage of couples that are same-sex for each state and DC over 2005–2019 using the American Community Survey.
+Author’s calculations.
+District of Columbia’s share (7.14%) is a significant outlier relative to the other states.
 3 Identification Strategy & Estimation
-The differential roll-out of anti-discrimination laws by state and locality over time and between same-sex and different-sex couples lends itself to a saturated triple difference-in-differences framework with the main outcomes of interest as the change in labor supply and pay.
+The differential roll-out of anti-discrimination laws by state and locality over time, and between same-sex and different-sex couples, naturally lends itself to a saturated triple-differences framework, where the primary outcomes of interest are changes in labor supply and pay.
 This strategy uses the variation presented in Figure 2, which shows how the laws changed over time by state and county.
 Formally, we estimate this equation on labor supply:
-$Y_{ipct}$ $=\alpha\;+\;\beta_{1}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{ENDA}_{ct}\bigr{]}\;+\;\sum_{s=1}^{S}\lambda_{s}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{SSM}_{st}\times\mathbf{1}_{\text{state}(c)=s}\bigr{]}$ (1)
-$\quad+\theta_{ct}+\mu_{pc}+\kappa_{pt}+X_{ipct}^{\prime}\,\delta+\varepsilon_{ipct}.$
-In our regression, equation (1) defines the labor market outcomes $Y_{ipct}$ for individual $i$, in partnership type $p$, residing in county $c$, at time $t$.
+$Y_{ipct}$ $=\alpha_{g}\;+\;\beta_{1g}\bigl[\mathrm{SSP}_{ipct}\times\mathrm{ENDA}_{ct}\bigr]\;+\;\lambda_{s(c),g}\bigl[\mathrm{SSP}_{ipct}\times\mathrm{SSM}_{s(c),t}\bigr]$ (1)
+$\quad+\theta_{gct}+\mu_{gpc}+\kappa_{gpt}+X_{ipct}^{\prime}\,\delta_{g}+\epsilon_{ipct}.$
+In our regression, equation (1) defines the labor market outcomes $Y_{ipct}$ for individual $i$ of sex $g$, in partnership type $p$, residing in county $c$ in state $s$, at time $t$.
 These outcomes include labor force participation, employment status, and the percentile rank of real wages (in 2019 dollars).
 The binary variable $\mathrm{SSP}_{ipct}$ indicates whether an individual is in a same‐sex partnership (1) or a different‐sex partnership (0).
-$\mathrm{SSM}_{st}$ indicates whether state $s$ recognizes same‐sex marriage in year $t$, and $\mathbf{1}_{\{\mathrm{state}(c)=s\}}=1$ if county $c$ lies in state $s$.
+$\mathrm{SSM}_{s(c),t}$ indicates whether the state $s(c)$ recognizes same‐sex marriage in year $t$, where $s(c)$ denotes the state corresponding to county $c$.
+Our key coefficient of interest is $\beta_{1g}$, which measures the sex-specific effect of the interaction of same‐sex partnership status with the presence of a county‐level Employment Non‐Discrimination Act, $\mathrm{ENDA}_{ct}$.
+Treatment status is computed using the earliest year of ENDA implementation at the county or state level, with individuals remaining classified as “treated” thereafter.
+All coefficients ($\alpha_{g},\beta_{1g},\lambda_{s(c),g},\delta_{g}$) and fixed effects ($\theta_{gct},\mu_{gpc},\kappa_{gpt}$) are fully interacted with sex $g$, so that the model is estimated jointly but allows each sex to have its own coefficients.
 The error term $\epsilon_{ipct}$ is clustered at the county level to account for within-county correlation.
@@
 Rural counties within a state are compared together with a true county fixed effect for metro area counties.
-Our key variable of interest is the interaction of same‐sex partnership status with the presence of a county‐level Employment Non‐Discrimination Act, $\mathrm{ENDA}_{ct}$.
-[^13] Treatment status is computed using the earliest year of ENDA implementation at the county or state level, with individuals remaining classified as “treated” thereafter.
-We include three sets of high‐dimensional fixed effects to account for a wide range of unobserved factors.
-First, $\theta_{ct}$ (county×$\times$×year) absorbs any time‐varying shocks or policy changes common to all individuals within a county in a given year.
-Second, $\mu_{pc}$ (partnership×$\times$×county) controls for any unobserved, time‐invariant characteristics that are specific to each partnership type within a county.
-Third, $\kappa_{pt}$ (partnership×$\times$×year) accounts for year‐specific shocks that affect each partnership type differentially.
-The vector $X_{ipct}^{\prime}\delta$ includes additional demographic controls (e.g., race, age, and education).
-The coefficient of interest, $\beta_{1}$, identifies how ENDA coverage $\mathrm{ENDA}_{ct}$ differentially affects same‐sex partners, net of any confounding from same‐sex marriage legalization.
-The summation term $\sum_{s=1}^{S}\lambda_{s}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{SSM}_{st}\bigr{]}$ partials out state‐specific labor‐market impacts of marriage legalization on same‐sex couples.
-The key variation used to estimate $\beta_{1}$ is the differential timing of when counties adopted ENDAs, and how this interacted with the labor market outcomes for individuals in same-sex partnerships relative to those in different-sex partnerships.
-This variation allows us to isolate the effect of anti-discrimination laws on labor market outcomes by comparing same-sex and different-sex couples across counties with and without ENDAs, before and after their passage.
-Thus, $\beta_{1}$ captures the differential impact of ENDA on labor supply and wage outcomes specifically for individuals in same-sex partnerships.
-[^14] We recognize the heterogeneous impact (e.g., Burn (2018) of anti-discrimination laws across different regions, yet our analysis aims to estimate an average effect for those directly subject to the laws.
-We limit our sample to people aged 25 to 65, and we estimate this model separately by sex using the ACS person weights.
-Following Chen and Roth (2024), we compute the percentile rank of annual and hourly real wages.
-[^15] Given the ACS’s 23-month reference period for earnings, as discussed by Hansen et al.
-(2020), this extended timeframe might not precisely align with the implementation dates of ENDAs, potentially biasing the immediate effects of these policies towards zero.
-To interpret $\beta_{1}$ as the causal effect of enacting employment protections for sexual orientation, we assume that no other factors are driving changes in the same-sex versus different-sex wage differential in states that do and do not have LGBT employment protections.
-One concern is that the distribution of workers in same-sex partnerships is not random, and states/counties that pass anti-discrimination laws have a higher concentration of workers in same-sex partnerships.
-However, with county-year fixed effects, comparisons are made among individuals within the same county in the same year, ensuring that workers in San Francisco are compared to others in San Francisco in that same year.
-A potential concern would be if there is sorting of workers in same-sex partnerships to areas with anti-discrimination laws protecting them, changing the composition of the area.
-To account for this, we remove individuals who moved in the past year as a robustness check.
-[^16] While we address potential sorting by excluding recent movers, unobserved migration patterns could still affect the composition of same-sex partnerships in jurisdictions with anti-discrimination laws.
-Another potential concern is that same-sex marriage legalization may confound our estimates.
-We address this by explicitly partialing out any state-specific effects of marriage legalization for same-sex partners.
-In particular, we include the term $\sum_{s=1}^{S}\lambda_{s}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{SSM}_{st}\times\mathbf{1}_{\text{state}(c)=s}\bigr{]}$, which absorbs any differential labor-market impacts associated with the varying timing of same-sex marriage legalizations across states.
-This ensures that our key coefficient, $\beta_{1}$, isolates the independent effect of the local Employment Non-Discrimination Act.
-While these controls mitigate potential bias, we caution that, despite our efforts, overlapping policies may still exert residual confounding influences on the estimated effects.
-To address potential biases from staggered policy adoption, we employ the methodology of Borusyak et al.
-(2021).
-Unlike standard TWFE models, which may misattribute post-treatment effects from early-treated units to later-treated ones, their approach uses only untreated observations to infer counterfactual outcomes for treated units.
-This strategy minimizes contamination and yields a more reliable estimate of $\beta_{1}$.
-Our identification hinges on the parallel trends assumption, that in the absence of anti-discrimination laws, labor market trends in counties with and without such laws would have evolved similarly.
-[^17] The difference-in-differences strategy hinges on the parallel trends assumption across groups and periods, with no anticipation of the laws prior to their implementation.
-We assess this assumption using event studies that estimate the dynamic effects of the policy, allowing us to observe pre-treatment trends and validate the plausibility of our identification strategy.
-We report the coefficients estimated via Borusyak et al.
-(2021) in the main text using a fully saturated triple-difference (DDD) specification, and standard TWFE along with other difference-in-differences (DiD) estimators in Section 5.
-As we show in Section 5, the estimates for same-sex and different-sex individuals estimated separately are similar across estimators.
+We restrict the sample to individuals aged 25–65 and estimate all models jointly by sex using ACS person weights.
+Following chen2024logs, we compute weighted within-year percentile ranks for real annual earnings and real hourly wages (expressed in 2019 dollars) by assigning individuals to weighted percentiles each year.
+We then construct three distributional indicators, at the 25th, 50th, and 75th percentiles, set to one when an individual’s percentile rank meets or exceeds each threshold.
+Because ACS earnings reflect a 23-month reference period (hansen2020labor), measured earnings may not align exactly with ENDA adoption dates, likely attenuating treatment effects.
+We include three sets of high-dimensional fixed effects, each fully interacted with sex to reflect the joint estimation.
+First, the county-by-year fixed effects $\theta_{gct}$ absorb all time-varying shocks or policy changes common to individuals of sex $g$ within county $c$ and year $t$.
+Second, the partnership-by-county fixed effects $\mu_{gpc}$ capture time-invariant characteristics specific to partnership type $p$ within county $c$ for each sex $g$.
+Third, the partnership-by-year fixed effects $\kappa_{gpt}$ absorb year-specific shocks that differentially affect partnership types for each sex.
+The vector $X_{ipct}^{\prime}\delta_{g}$ includes demographic controls (race, age, and education), also allowed to vary by sex.
+The coefficient of interest, $\beta_{1g}$, identifies how ENDA coverage $\mathrm{ENDA}_{ct}$ differentially affects same-sex partners relative to different-sex partners for each sex $g$, net of any confounding from same-sex marriage legalization.
+The interacted term $\lambda_{s(c),g}\bigl[\mathrm{SSP}_{ipct}\times\mathrm{SSM}_{s(c),t}\bigr]$ partials out state-specific labor-market impacts of same-sex marriage legalization on same-sex partners for each sex.
+The identifying variation for $\beta_{1g}$ comes from differences in the timing of county-level ENDA adoption, combined with how these policy changes differentially affect individuals in same-sex versus different-sex couples within counties, holding constant sex-specific fixed effects.
+This allows us to isolate how employment protections change labor supply and wage outcomes for same-sex partners relative to different-sex partners of the same sex.
+Thus, $\beta_{1g}$ captures the sex-specific differential impact of ENDA on labor market outcomes for individuals in same-sex couples.
+[^9] We recognize that the effects of anti-discrimination laws may vary across regions (e.g., Burn2018), but our objective is to estimate an average effect for those directly subject to the laws.
+To interpret $\beta_{1g}$ as the causal effect of enacting employment protections for sexual orientation, we assume that no other factors differentially shift outcomes for same-sex versus different-sex partners in counties that do and do not adopt ENDAs.
+Below, we outline the threats to this assumption and describe how our empirical design addresses each one.
+3.1 Threats to Identification
+We group the main identification concerns into three categories: (i) threats that can be addressed directly through sample restrictions, (ii) threats mitigated through design choices and high-dimensional fixed effects, and (iii) residual concerns that cannot be fully ruled out but primarily generate attenuation rather than spurious effects.
+3.1.1 Threats Addressed
+Selective Migration.
+Individuals in same-sex couples may relocate to jurisdictions with stronger LGBT protections, altering the composition of treated counties.
+To assess this possibility, we augment the specification with an additional layer of interactions distinguishing non-mover females, non-mover males, mover females, and mover males.
+This four-way structure lets ENDA effects differ flexibly by migration status.
+Comparing coefficients across these groups provides a direct test for migration-induced bias, and the results show no meaningful differences.
+[^10] Long-run unobserved mobility may still shift the composition of same-sex couples in treated areas, though our mover–non-mover comparisons indicate limited bias.
+Household Formation and Sample Composition.
+A potential concern is that ENDAs may affect the likelihood that individuals enter or remain in a same-sex partnership, thereby altering who appears in our coupled-household sample.
+If anti-discrimination protections increase relationship stability or the willingness to publicly identify a partner, then treatment could shift the composition of the same-sex population in ways correlated with labor market outcomes.
+To assess this concern, we estimate whether the share of adults living in same-sex couples changes following ENDA adoption at the county level.
+We find no evidence of differential changes for men or women.
+This suggests that ENDAs do not meaningfully affect partnership formation or household structure, and thus do not induce endogenous entry into the coupled sample.
+As a result, the identifying variation in our main specification is unlikely to reflect compositional changes in who is observed as part of a same-sex household.
+Residence–Work Geography Mismatch.
+ENDAs apply to the county of employment, whereas ACS geography reflects the county of residence.
+As with migration, this creates the possibility of misclassifying some individuals with respect to treatment if they commute across county (or state) lines.
+We evaluate this concern using the same strategy as our mover–non-mover analysis: we allow ENDA effects to differ flexibly across groups likely to experience geographic mismatch.
+The estimates are stable across these specifications, indicating that any residence–work misalignment does not materially influence the results.
+As with other sources of geographic measurement error, this type of misclassification primarily generates attenuation rather than spurious treatment effects.
+3.1.2 Threats Mitigated
+Staggered Adoption.
+To address potential biases from staggered policy adoption, we employ the methodology of borusyak2021revisiting.
+Unlike standard TWFE models, which may misattribute post-treatment effects from early-treated units to later-treated ones, their estimator uses only untreated observations to form counterfactuals for treated units.
+Because we estimate the model jointly by sex, this approach is applied separately within each sex group, yielding the sex-specific treatment effects $\beta_{1g}$ without contamination from negative weighting.
+Our identification hinges on the parallel trends assumption: in the absence of ENDAs, labor market trends for same-sex and different-sex partners would have evolved similarly within each sex group across treated and untreated counties.
+[^11] This requires no anticipatory responses and equal pre-treatment trends within sex-by-partnership cells across treatment timing groups.
+We assess this assumption with event-study analyses (Figures 2–3), which show no evidence of differential pre-trends.
+[^12] We conduct a sensitivity analysis following RambachanRoth2023.
+As detailed in Appendix Table C.8 and discussed in Appendix C.7, we find that our extensive margin estimates for men are robust to post-treatment trend violations up to 40% of the maximal pre-trend deviation, while the null results for women persist even under relaxed assumptions.
+Same-Sex Marriage Legalization.
+Same-sex marriage (SSM) legalization may confound the effects of ENDAs by shifting household specialization or labor supply among same-sex partners.
+We directly partial out these effects by including sex-specific interactions of same-sex partnership status with SSM legalization status.
+This absorbs any differential impacts of state-level marriage reforms on same-sex couples.
+[^13] While these controls mitigate potential bias, overlapping policies may still exert residual confounding influences on the estimated effects.
+As an additional check, Section 5.2 restricts the sample to pre-SSM years within each state, again yielding similar estimates.
+Differential Local Trends and Endogenous Adoption.
+Jurisdictions adopting ENDAs tend to be more supportive toward sexual minorities.
+If positive sentiment drives both ENDA passage and improvements in labor market outcomes, estimates may reflect omitted variables.
+Our saturated triple-difference specification addresses this by including county-year, partnership-by-year, and partnership-by-county fixed effects, all interacted with sex.
+These fixed effects ensure comparisons are made only within county-year cells and within sex-by-partnership cells, substantially reducing concerns about correlated local shocks or endogenous policy timing.
+3.1.3 Residual Threats
+Misclassification of Sexual Orientation.
+Our comparison relies on household composition rather than self-identified sexual orientation.
+Bisexual individuals in different-sex relationships enter the control group, and some gay or lesbian workers may also be misclassified into that group.
+Because lesbian women tend to earn more than heterosexual women, gay men tend to earn less than heterosexual men, and bisexual workers earn less than both (mize2016sexual), such misclassification does not shift group earnings in a uniform direction.
+In addition, bisexual workers may respond differently to ENDAs than gay or lesbian workers, so the sign of any resulting bias in the estimated ENDA coefficients is theoretically ambiguous.
+At the same time, misclassification that places gay or lesbian individuals into the control group reduces the contrast between partnership types and therefore operates through the usual attenuation mechanism.
+Overall, these forms of misclassification complicate interpretation but are unlikely to generate spurious positive treatment effects.
+Overlap and Heterogeneity in Local and State ENDA Laws.
+Local and state ENDAs vary in scope and enforcement.
+Our coding focuses on the first year in which a county receives broad protections in private employment and housing.
+For example, although Delaware extended protections to state employees in 1999, we code treatment beginning in 2009 when comprehensive protections were enacted.
+Similarly, some local ordinances initially covered only public employees before broader laws passed.
+These earlier, narrower protections may still have influenced social attitudes or employer behavior prior to full coverage.
+In addition, the strength of protections differs across jurisdictions, some laws exempt small firms or specific sectors, and many large employers adopt nondiscrimination policies independent of legal requirements (e.g., firms scoring highly on the Corporate Equality Index).
+Collapsing this heterogeneity into a single binary ENDA indicator necessarily compresses meaningful variation.
+This generates classical attenuation bias, pushing estimated effects toward zero and making our findings conservative.
+Geographic Misclassification From PUMA-to-County Assignment.
+ACS microdata identify geography at the PUMA level, which sometimes maps imperfectly to counties.
+A small share of individuals may thus be assigned to counties with classification error.
+We do not exclude ambiguous PUMAs; however, such misclassification behaves like classical measurement error in the treatment indicator and therefore attenuates estimated treatment effects rather than creating false positives.
+To summarize, our empirical strategy exploits a fully saturated triple-differences design estimated jointly by sex.
+We control for confounding policy changes such as same-sex marriage legalization using sex-specific marriage interactions, and we implement the estimator of borusyak2021revisiting to mitigate biases arising from staggered ENDA adoption.
+Together, these features allow us to identify how local ENDAs affect the labor market outcomes of individuals in same-sex couples relative to their different-sex counterparts within each sex group.
+As we show in Appendix C.1, the estimates for males and females are similar in sign and magnitude across alternative estimators, supporting the robustness of our findings.
 4 Primary Findings
 4.1 Extensive & Intensive Margins
-In Table 4, we present regression results that examine the extensive and intensive margins of labor supply for same-sex partnerships.
-The coefficient $\beta_{1}$ captures the effect of anti-discrimination laws on the labor supply outcomes for men in male same-sex partnerships and for women in female same-sex partnerships.
-Panel A focuses on men in male same‐sex partnerships.
-Columns (1) and (2) show that anti-discrimination laws reduce both the labor force participation and employment gaps by approximately 3.8 and 3.7 percentage points, which translates into reductions of about 78% and 77% respectively.
+In Table 4, we present regression results that examine the extensive and intensive margins of labor supply for same-sex couples.
+The coefficient $\beta_{1g}$ captures the effect of anti-discrimination laws on the labor supply outcomes for men in male same-sex couples and for women in female same-sex couples.
+Table 4: Effect of Anti-Discrimination Laws: Extensive & Intensive Margin of Labor Supply
+(1) (2) (3) (4) (5) (6)
+Extensive Margin Intensive Margin
+Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
+ENDA $\times$ SSP: Female -0.0037 -0.0050 0.9519 -0.4143 0.7017∗∗ -0.8915∗∗∗
+(0.0116) (0.0135) (0.5332) (0.6235) (0.3080) (0.3349)
+ENDA $\times$ SSP: Male 0.0379∗∗∗ 0.0370∗∗∗ 1.1289∗∗ 1.5089∗∗ -0.0920 -0.0719
+(0.0130) (0.0135) (0.4984) (0.6276) (0.3727) (0.2821)
+$p$-value: M - F 0.0032 0.0177 0.7908 0.0082 0.1160 0.0465
+Employed Only X X
+Observations 7,781,358 7,781,358 7,781,358 7,781,358 5,765,580 5,765,580
+Mean (F) 0.818 0.782 34.42 39.51 41.25 47.27
+Mean (M) 0.825 0.790 35.53 39.96 42.41 47.69
+  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
+We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment.
+The regressions are estimated jointly, allowing the treatment effect to differ for men and women through interaction terms.
+Columns (1–2) focus on the extensive margin, examining changes in labor force participation and employment.
+Columns (3–6) explore the intensive margin using variation in weeks worked per year and usual weekly hours; Columns (5–6) restrict the sample to employed individuals.
+The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples, respectively.
+Reported $p$-values test whether the ENDA effect differs significantly between men and women.
+Standard errors are clustered at the county level.
+$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
+The joint estimation allows the effects of anti-discrimination laws to differ for men and women through interaction terms.
+Columns (1) and (2) show that, for men in male same-sex couples, anti-discrimination laws reduce both the labor force participation and employment gaps by approximately 3.8 and 3.7 percentage points, which translates into reductions of about 78% and 77%, respectively.
 These coefficients are statistically significant at the 1% level.
@@
 However, when the sample is restricted to employed men (Columns (5) and (6)), these effects become statistically insignificant.
-This pattern suggests that the primary channel through which these laws operate is by increasing the rate of employment.
-Panel B reports analogous estimates for women in female same‐sex partnerships.
-In contrast, the extensive margin results for women (Columns (1) and (2)) yield statistically insignificant coefficients of $-0.37$ and $-0.50$ percentage points.
-Women in same‐sex partnerships exhibit higher labor force participation rates than women in different‐sex partnerships.
-This higher baseline may partially explain the lack of significant improvement on the extensive margin, as there is less room for gains.
-On the intensive margin, Column (3) shows a modest increase of 0.95 weekly hours (significant at the 10% level) for all women, while Column (4) finds a small, statistically insignificant reduction in weeks worked.
-When focusing on employed women (Columns (5) and (6)), we observe a statistically significant increase of 0.70 weekly hours (at the 5% level) alongside a significant decrease of 0.89 weeks worked (at the 1% level).
-These mixed findings suggest that anti-discrimination laws may lead employed women to work longer hours per week but in fewer weeks overall, reflecting a complex adjustment on the intensive margin.
+This pattern suggests that the primary channel through which these laws operate is by increasing the rate of employment rather than by altering hours worked among those already employed.
+The estimates for women in same-sex couples follow a different pattern.
+Along the extensive margin (Columns (1) and (2)), the coefficients are small, negative, and statistically insignificant, on the order of 0.3–0.5 percentage points, which are not only imprecise but also economically small relative to mean participation rates.
+On the intensive margin, Column (3) shows a modest increase of 0.95 weekly hours (significant at the 10% level), while Column (4) shows a small, statistically insignificant reduction in weeks worked.
+When focusing on employed women (Columns (5) and (6)), we observe a statistically significant increase of 0.70 weekly hours and a significant decrease of 0.89 weeks worked.
+These mixed but generally modest effects underscore that, for women, the evidence for labor-market responses is weaker and less systematic than for men, consistent with the possibility that adjustments for women occur outside the labor-market margins examined here.
+Finally, the joint estimation allows for explicit comparison across sexes.
+The $p$-values reported in the table indicate that the extensive margin effects differ significantly between men and women, with men experiencing large, positive, and statistically significant gains while women show no measurable changes.
+Additionally, the intensive margin effect on weeks worked is also significantly larger for men than for women.
+These differences underscore meaningful heterogeneity in how anti-discrimination laws affect male and female same-sex couples in the labor market.
 4.2 Wages
 As shown in Table 4, anti-discrimination laws significantly affect labor force participation and employment outcomes, thereby potentially altering the composition of the employed sample and raising endogeneity concerns for wage analyses.
-Table 5 shows regression results on the wage gap, using real hourly wages and annual earnings in 2019 dollars, measured by percentile ranks and computed jointly for men and women.
-As Chen and Roth (2024) argue, this robust, scale-invariant metric reduces the impact of outliers and reflects an individual’s relative wage position, allowing us to evaluate how anti-discrimination laws shift wage standings for individuals in same-sex partnerships.
-For men in same-sex partnerships, the increased employment rates reported in Table 4 are accompanied by significant wage gains.
-Specifically, anti-discrimination laws boost hourly wages by roughly 2.7 percentiles and annual earnings by 3 percentiles on average.
-Furthermore, these laws raise the probability of ranking above the 25th and 50th percentiles in the hourly wage distribution by 3.6 and 3.4 percentage points, respectively, although the effect at the 75th percentile is not significant.
-In the annual earnings distribution, anti-discrimination laws improve the likelihood of being in the upper tiers by 3.6, 5.2, and 3.7 percentage points at the 25th, 50th, and 75th percentiles, respectively.
-For women in same-sex partnerships, the wage distribution remains largely unaffected by anti-discrimination laws, as evidenced by statistically insignificant estimates across the board.
-This finding is consistent with the muted effects on employment outcomes observed in Table 4 for women.
-These findings suggest that the wage gains for men in same-sex partnerships operate primarily through increased employment.
-New entrants to the labor force, individuals who were previously at the bottom of the wage distribution, enter at various points, thereby shifting the overall distribution upward.
+Table 5 shows regression results on the wage gap, using real hourly wages and annual earnings in 2019 dollars, measured by percentile ranks and computed jointly for men and women with heterogeneous ENDA effects by sex.
+[^14] We first compute weighted within-year percentile ranks for real annual earnings and real hourly wages by assigning individuals to weighted percentiles each survey year; we then construct three distributional indicators, set to one when an individual’s percentile rank meets or exceeds the 25th, 50th, and 75th percentiles, respectively.
+This robust, scale-invariant metric reduces the impact of outliers and reflects an individual’s relative wage position, allowing us to evaluate how anti-discrimination laws shift wage standings for individuals in same-sex couples.
+[^15] Following chen2024logs, including non-employed individuals in the wage-rank distribution improves transparency by combining extensive- and intensive-margin responses into a single metric.
+However, this approach does not recover the causal effect of ENDAs on potential wages for all individuals, since ranks for non-employed individuals are mechanically assigned rather than observed wage offers.
+Table 5: Effect of Anti-Discrimination Laws: Wages and Earnings
+Hourly Wage Annual Earnings
+(1) (2) (3) (4) (5) (6) (7) (8)
+Percentile $\geq$25th $\geq$50th $\geq$75th Percentile $\geq$25th $\geq$50th $\geq$75th
+ENDA $\times$ SSP: Female -0.3333 0.0045 -0.0117 -0.0061 -0.5994 0.0051 -0.0204 -0.0016
+(0.8450) (0.0118) (0.0143) (0.0154) (0.8315) (0.0118) (0.0134) (0.0161)
+ENDA $\times$ SSP: Male 2.6310∗∗ 0.0366∗∗∗ 0.0326 0.0232 2.9315∗∗∗ 0.0365∗∗∗ 0.0467∗∗ 0.0334∗∗
+(1.1771) (0.0139) (0.0182) (0.0178) (1.1019) (0.0139) (0.0185) (0.0154)
+$p$-value: M - F 0.0379 0.0340 0.0776 0.2093 0.0066 0.0382 0.0049 0.0713
+Observations 7,781,358 7,781,358 7,781,358 7,781,358 7,781,358 7,781,358 7,781,358 7,781,358
+Mean (F) 48.73 0.799 0.513 0.247 48.68 0.798 0.515 0.241
+Mean (M) 51.51 0.791 0.557 0.305 51.69 0.791 0.561 0.305
+  • Notes: Data are derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
+We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment.
+The regressions are estimated jointly, allowing the treatment effect to differ for men and women through interaction terms.
+Columns (1–4) report effects on hourly real wage ranks, while Columns (5–8) report effects on annual real earnings ranks.
+Columns (2–4) and (6–8) estimate the probability of being above the 25th, 50th, and 75th percentiles of the respective income distributions.
+All earnings measures are expressed in 2019 dollars.
+The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples, respectively.
+Reported $p$-values test whether the ENDA effect differs significantly between men and women.
+Standard errors are clustered at the county level.
+  • $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
+For men in same-sex couples, the increased employment rates reported in Table 4 are accompanied by significant wage gains.
+Specifically, anti-discrimination laws boost hourly wages by roughly 2.6 percentiles and annual earnings by 2.9 percentiles on average.
+Furthermore, these laws raise the probability of ranking above the 25th and 50th percentiles in the hourly wage distribution by 3.7 and 3.3 percentage points, respectively, although the effect at the 75th percentile remains statistically insignificant.
+In the annual earnings distribution, anti-discrimination laws improve the likelihood of being in the upper tiers by 3.6, 4.7, and 3.3 percentage points at the 25th, 50th, and 75th percentiles, respectively.
+For women in same-sex couples, the wage distribution remains largely unaffected by anti-discrimination laws.
+The estimates are statistically insignificant and also small in magnitude—typically well under one percentile point, suggesting limited scope for economically meaningful wage effects.
+This pattern is consistent with the muted employment responses for women reported in Table 4.
+Joint estimation allows us to formally compare effects across sexes.
+The $p$-values reported in Table 5 show that the wage effects for men are significantly larger than those for women at most points in the wage and earnings distributions, reinforcing that anti-discrimination laws generate substantial improvements in labor market standing for men in same-sex couples but have comparatively limited effects for women.
+These findings suggest that the wage gains for men in same-sex couples operate primarily through increased employment on the extensive margin.
+New entrants to the labor force, individuals who were previously non-employed or concentrated at the bottom of the distribution, enter at various points in the wage distribution, shifting the overall distribution upward, particularly around the lower and middle percentiles where such workers are most likely to sort.
+Specialization models among couples imply that primary earners are typically already located high in the earnings distribution; thus, if ENDAs primarily induce secondary earners to enter work or increase their labor supply, we should expect relatively muted effects at the upper end of the hourly wage distribution.
+This is precisely what we observe: the 75th percentile of the hourly wage distribution shows no significant response to ENDAs.
+By contrast, the 75th percentile of the annual earnings distribution rises significantly.
+While this may appear surprising at first glance, it likely reflects a combination of wage gains, increased hours or weeks worked, and the fact that some new entrants can move into higher annual earnings ranks through expanded labor supply.
+Overall, the evidence points to selection through secondary earners entering employment and adjusting labor supply, rather than upward movements among already high-earning primary earners.
 4.3 Impact Dynamics
-We estimate Equation (1) using event study regressions following Borusyak et al.
-(2021) to trace the evolution of the impact estimates.
+We estimate Equation (1) using event study regressions following borusyak2021revisiting to trace the evolution of the impact estimates.
 Figure 3 displays the results for extensive labor supply outcomes: Panel (a) shows labor force participation rates, while Panel (b) presents employment levels.
-For men in male same‐sex partnerships, both panels reveal an immediate increase in labor force participation and employment following the enactment of ENDAs, although this positive effect wanes over time.
-In contrast, we see no impacts on women in same-sex partnerships.
-The pre-treatment trends are largely parallel, supporting the plausibility of the parallel trends assumption, yet the post-treatment dynamics diverge sharply by gender.
-The evidence here suggests that ENDAs lead to a temporary boost in labor market entry for men in same-sex partnerships, while not affecting employment and participation rates for women in same-sex partnerships.
+For men in male same-sex couples, both panels show an immediate rise in labor force participation and employment following ENDA adoption, with the largest effects appearing in the early post-treatment years.
+While the point estimates decline somewhat over time, our data do not extend far enough to determine whether these effects are temporary or simply leveling off.
+In contrast, the estimates for women in same-sex couples are small and statistically insignificant, indicating little evidence of a labor-market response.
+The pre-treatment trends are largely parallel, supporting the plausibility of the identifying assumption, and the post-treatment dynamics again highlight pronounced gender differences in responses to ENDAs.
+((a)) Labor Force Participation
+((b)) Employment
+Figure 3: Labor Supply Impacts on the Extensive Margin
+Notes: Event study plots on the difference in labor supply between people in same-sex couples and different-sex couples broken down by sex following the county-level and state-level anti-discrimination laws.
+Coefficients with 95% confidence intervals.
+Standard errors are clustered at the county level.
+Reference year is -4.
+((a)) Percentile Rank of Real Hourly Wage
+((b)) Percentile Rank of Annual Wage
+Figure 4: Impacts on Wages
+Notes: Event study plot on the difference in pay between people in same-sex couples and different-sex couples broken down by sex following the county-level and state-level anti-discrimination laws.
+Coefficients with 95% confidence intervals.
+Standard errors are clustered at the county level.
+Reference year is -4.
 Figure 4 presents the event study regressions for male and female pay disparities.
-The results mirror the trends observed in Figure 3, with a significant wage jump occurring one year after the law’s enactment for men in same-sex partnerships.
+The results mirror the trends observed in Figure 3, with a significant wage jump occurring one year after the law’s enactment for men in same-sex couples.
 However, the difference between the one-year and four-year post‑enactment estimates is relatively modest, suggesting that the initial wage improvement for men stabilizes and remains persistent over time.
 We find no statistically significant effects on wage disparities for women in any post‑ENDA period.
-These findings indicate that while anti‑discrimination laws yield immediate and lasting wage benefits for men in same‑sex partnerships, the impact on wage outcomes for women in same-sex partnerships is insignificant.
+These findings indicate that while anti‑discrimination laws yield immediate and lasting wage benefits for men in same‑sex couples, the impact on wage outcomes for women in same-sex couples is insignificant.
 5 Robustness Checks
-To ensure the robustness of our findings on the labor market impacts of anti-discrimination laws, we conduct a series of sensitivity analyses addressing potential sources of bias.
-Specifically, we examine (i) the endogeneity of law adoption using state-level support for same-sex marriage as a proxy for public sentiment, (ii) the potential confounding influence of same-sex marriage legalization, (iii) the role of migration patterns, (iv) geographic heterogeneity between metropolitan and non-metropolitan areas, (v) reporting changes associated with ACS redesign, (vi) the impact of applying sample weights, and (vii) the validity of our identification strategy using placebo tests and alternative estimators.
+To assess the robustness of our findings and address key threats to identification, we conduct several sensitivity analyses.
+These tests examine (i) the endogeneity of law adoption using state-level support for same-sex marriage, (ii) potential confounding from same-sex marriage legalization, (iii) migration responses, (iv) metro–non-metro heterogeneity, (v) sensitivity to commuting, and (vi) reporting changes from the ACS redesign.
+For exercises (ii)–(vi), we estimate a fully interacted framework that splits men and women into subsample and complementary groups, allowing us to recover ENDA effects for each subgroup and test differences across sex and across samples.
 5.1 Endogenous Adoption of Anti-Discrimination Laws
@@
 Figure 5 presents the event study plot.
+Figure 5: Impacts on Same-Sex Marriage Support
+Notes: Event study plot showing how support for same-sex marriage changes following state-level anti-discrimination laws.
+Polling data comes from Pew Polling and encompasses 2005-2017 for every state except Hawaii and Alaska, which are missing for 2005-2008.
+Standard errors are clustered at the state level.
 The event study indicates that state-level anti-discrimination laws significantly boost support for same-sex marriage.
@@
 Nevertheless, to further isolate the impact of anti-discrimination laws, we restrict our sample to periods before same-sex marriage became legal in each state (e.g., excluding observations from 2008 onward for states that legalized in 2008).
-Table 6 presents results restricting the sample to periods before same-sex marriage legalization.
+Table 6: Effects of Anti-Discrimination Laws on Labor Supply: Pre–Same-Sex Marriage Legalization
+(1) (2) (3) (4) (5) (6)
+Extensive Margin Intensive Margin
+Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
+Pre–SSM Legalization
+ENDA $\times$ SSP: Female -0.0157 -0.0215 1.1714 -0.6654 1.0376∗∗∗ -1.2330∗∗∗
+(0.0153) (0.0158) (0.6726) (0.7928) (0.3682) (0.3997)
+ENDA $\times$ SSP: Male 0.0283 0.0311 0.8892 0.4945 -0.1655 -0.7909∗∗
+(0.0186) (0.0193) (0.6512) (0.8727) (0.4310) (0.3950)
+$p$: M – F (Pre–SSM) 0.0186 0.0202 0.7304 0.2214 0.0192 0.4193
+$p$: M (Pre–SSM – Post-SSM) 0.1196 0.7317 0.0293 0.0048 0.0000 0.0004
+$p$: F (Pre–SSM – Post-SSM) 0.0702 0.0039 0.0024 0.0270 0.0299 0.0125
+Observations 7,779,108 7,779,108 7,779,108 7,779,108 5,763,735 5,763,735
+Mean (F) 0.817 0.774 34.44 39.13 41.28 46.85
+Mean (M) 0.818 0.778 35.39 39.44 42.40 47.27
+  • Notes: Data are derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
+We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment.
+All regressions in this table are estimated jointly with a fully interacted model that allows ENDA effects to vary across four groups: men and women, each in the pre–same-sex-marriage and post–same-sex-marriage periods.
+This framework nests the separate-by-sex estimations and enables formal tests of equality across groups.
+We report the ENDA effects estimated on the Pre–SSM Legalization sample and include $p$-values testing whether these coefficients differ from those obtained in the complementary Post–SSM sample.
+Columns (1–2) focus on the extensive margin, examining changes in labor force participation and employment.
+Columns (3–6) analyze intensive-margin outcomes, usual weekly hours and weeks worked, where Columns (5–6) restrict the sample to employed individuals.
+The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples in the Pre-SSM sample, respectively.
+Reported $p$-values test differences across sex and across robustness conditions.
+Standard errors are clustered at the county level.
+$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
+Table 6 presents results from the jointly estimated model, reporting ENDA effects separately for the pre–same-sex-marriage subsample while also providing $p$-values testing whether these effects differ significantly from those in the complementary post-legalization period.
 Although this approach removes nearly half of the same-sex partnership observations and reduces statistical power, the qualitative patterns remain robust.
-Specifically, the point estimates for both labor force participation and employment remain comparable, exceeding 3 percentage points, with labor force participation retaining statistical significance.
-However, the employment coefficient for men declines from 0.037 to 0.0319, becoming borderline insignificant (p-value = 0.102), and similar shrinkages are observed on the intensive margin.
-In contrast, for women in same-sex partnerships, the intensive margin estimates are slightly larger.
-Overall, this sensitivity analysis indicates that our identification strategy is not driven by concurrent same-sex marriage policies, as the observed effects of anti-discrimination laws on labor market outcomes for same-sex couples persist despite the reduced sample size.
+For men in same-sex couples, the point estimates for both labor force participation and employment remain comparable to the baseline, exceeding 3 percentage points, with labor force participation retaining statistical significance.
+The employment coefficient declines from 0.037 to 0.0311 and becomes borderline insignificant, consistent with the loss of power, and similar attenuations appear on the intensive margin.
+For women, the intensive-margin estimates are slightly larger in the restricted sample, though still small in magnitude and statistically weak.
+The pre–post $p$-value tests reported in Table 6 indicate that several coefficients differ across the two periods, particularly on the intensive margin.
+Still, the qualitative pattern of the ENDA effects is preserved: the signs of the coefficients remain stable, and the largest impacts continue to appear on the extensive margin for men and on the intensive margin for women.
+Moreover, even within the pre–SSM subsample, the male–female comparison on the extensive margin remains statistically significant and positive, reinforcing that men in same-sex couples experience systematically larger gains in labor force participation and employment.
+Thus, while same-sex marriage legalization may shift the magnitudes of some estimates, it does not overturn the core conclusion that anti-discrimination laws improve labor market outcomes for individuals in same-sex couples.
+This confirms that our identification strategy is not driven solely by concurrent same-sex marriage policies.
 5.3 Migration
 Next, we account for the possibility that migration patterns could influence our results.
-If LGB individuals were disproportionately moving to areas with recently enacted anti‑discrimination laws, it could bias the treatment effects.
-To address this concern, we restrict our sample to individuals who had not moved in the past year and re‑estimate the models presented in Table 4.
-By focusing on a more stable population, we can better isolate the effects of anti‑discrimination laws on labor outcomes.
+If LGB individuals were disproportionately moving to areas with recently enacted anti-discrimination laws, it could bias the treatment effects.
+To address this concern, we restrict our sample to individuals who had not moved in the past year and re-estimate the models presented in Table 4.
+By focusing on a more stable population, we can better isolate the effects of anti-discrimination laws on labor outcomes.
 The corresponding estimates, reported in Table 7, remain consistent with the full sample results, providing further robustness and suggesting that migration does not significantly alter our conclusions.
-Interestingly, the estimates for men in same‑sex partnerships are slightly larger for this subset, which may indicate that the positive employment effects are even more pronounced among those who are less mobile.
+Interestingly, the estimates for men in same-sex couples are slightly larger for this subset, which may indicate that the positive employment effects are even more pronounced among those who are less mobile.
+Table 7: Effects of Anti-Discrimination Laws on Labor Supply: Analysis of Non-Movers
+(1) (2) (3) (4) (5) (6)
+Extensive Margin Intensive Margin
+Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
+Non-Movers
+ENDA $\times$ SSP: Female 0.0050 0.0048 1.1290 0.1140 0.7276∗∗ -0.5031
+(0.0142) (0.0147) (0.6189) (0.7471) (0.3084) (0.3944)
+ENDA $\times$ SSP: Male 0.0421∗∗∗ 0.0424∗∗∗ 1.2112∗∗ 1.7705∗∗ -0.0583 0.2447
+(0.0145) (0.0153) (0.5690) (0.7291) (0.4024) (0.2941)
+$p$: M – F (Non-Movers) 0.0209 0.0412 0.9143 0.0503 0.1284 0.1084
+$p$: M (Non-Movers – Movers) 0.3954 0.2811 0.6663 0.5295 0.9676 0.4559
+$p$: F (Non-Movers – Movers) 0.0504 0.0763 0.2704 0.0013 0.2691 0.0146
+Observations 7,781,169 7,781,169 7,781,169 7,781,169 5,765,391 5,765,391
+Mean (F) 0.815 0.783 34.34 39.57 41.32 47.55
+Mean (M) 0.823 0.791 35.44 39.99 42.45 47.89
+  • Notes: Data are derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
+We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment.
+All regressions in this table are estimated jointly with a fully interacted model that allows ENDA effects to vary across four groups: men and women, each in the mover subsample and in the non-mover subsample.
+This framework nests the separate-by-sex estimations and enables formal tests of equality across groups.
+We report the ENDA effects estimated for non-movers and include $p$-values testing whether these coefficients differ from those obtained in the complementary sample of recent movers.
+Columns (1–2) focus on the extensive margin, examining changes in labor force participation and employment.
+Columns (3–6) analyze intensive-margin outcomes, usual weekly hours and weeks worked, where Columns (5–6) restrict the sample to employed individuals.
+The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples in the non-mover sample, respectively.
+Reported $p$-values test differences across sex and across the mover versus non-mover samples.
+Standard errors are clustered at the county level.
+$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
+The joint-estimation framework, which interacts ENDA exposure with both gender and migration status (male–mover, male–non-mover, female–mover, female–non-mover), also allows us to test whether the ENDA effects differ between movers and non-movers.
+For men, the $p$-value tests show no statistically significant differences across the two samples, reinforcing that migration does not meaningfully drive the male results.
+Among women, several coefficients do differ between movers and non-movers, but the overall qualitative pattern of effects is preserved and most estimates remain small or statistically insignificant.
+Importantly, the male–female comparison on the extensive margin remains statistically significant within the non-mover subsample, indicating that the gender differences highlighted in the main specification are not driven by migration behavior.
 5.4 Metropolitan Areas
-We estimate our model specifically for residents in metropolitan areas to address concerns about potential differences between urban and rural populations.
-By focusing on metro areas, where labor markets may function differently due to higher population density and greater economic diversity, we aim to test whether the effects of anti-discrimination laws vary by geographic location.
-Table 8 shows that when the sample is restricted to metropolitan areas, the overall patterns for men in same‑sex partnerships remain consistent, though the estimated effects of anti‑discrimination laws on labor supply are slightly smaller than in the full sample.
-This may suggest that the urban labor market environment, with its higher baseline employment levels and different institutional dynamics, somewhat tempers the positive impact seen elsewhere.
-In contrast, for women in same‑sex partnerships, the estimates become significantly negative across several outcomes, labor force participation, employment, and weeks worked.
-One plausible explanation is that ENDAs may improve employment stability in urban labor markets, and with high childcare costs in metro areas, a stable job by one partner might encourage women to opt out of the workforce.
-With a secure income from the other partner, some women may choose to stay home to manage childcare responsibilities rather than work additional hours or remain continuously employed.
-This pattern highlights clear geographic heterogeneity, as the dynamics in urban settings differ markedly from those in the broader sample for women in same‑sex partnerships.
+We re-estimate our model specifically for residents in metropolitan areas to assess whether the effects of anti-discrimination laws vary by geographic context.
+Table 8 shows that, within metro areas, the qualitative patterns for men in same-sex couples remain consistent with the main results: extensive-margin effects are positive and statistically significant, and the joint-estimation framework again shows that these effects are significantly larger for men than for women.
+Table 8: Effects of Anti-Discrimination Laws on Labor Supply: Analysis of Metro Areas
+(1) (2) (3) (4) (5) (6)
+Extensive Margin Intensive Margin
+Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
+Metro Areas
+ENDA $\times$ SSP: Female -0.0262∗∗ -0.0284∗∗ -0.4601 -1.3804∗∗ -0.4479 -1.3355∗∗∗
+(0.0115) (0.0129) (0.7244) (0.6511) (0.6693) (0.3471)
+ENDA $\times$ SSP: Male 0.0352∗∗ 0.0398∗∗ 0.8139 1.4503 -0.4661 -0.2430
+(0.0159) (0.0162) (0.5418) (0.7403) (0.3735) (0.3455)
+$p$: M – F (Metro) 0.0005 0.0007 0.1692 0.0013 0.9819 0.0248
+$p$: M (Metro – Non-Metro) 0.6882 0.5923 0.6851 0.2116 0.6688 0.0463
+$p$: F (Metro – Non-Metro) 0.0263 0.0621 0.0065 0.1878 0.0019 0.1576
+Observations 7,781,334 7,781,334 7,781,334 7,781,334 5,765,559 5,765,559
+Mean (F) 0.834 0.801 35.20 40.42 41.43 47.52
+Mean (M) 0.847 0.812 36.50 41.09 42.62 47.96
+  • Notes: Data are derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
+We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment.
+All regressions in this table are estimated jointly with a fully interacted model that allows ENDA effects to vary across four groups: men and women, each in the metro subsample and in the non-metro subsample.
+This framework nests the separate-by-sex estimations and enables formal tests of equality across groups.
+We report the ENDA effects estimated for the metro subsample and include $p$-values testing whether these coefficients differ from those in the non-metro population.
+Columns (1–2) focus on the extensive margin, examining labor force participation and employment.
+Columns (3–6) analyze intensive-margin outcomes, usual weekly hours and weeks worked, with Columns (5–6) restricted to employed individuals.
+The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples in the metro-area sample, respectively.
+Standard errors are clustered at the county level.
+$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
+For women in same-sex couples, the metro-area estimates become negative across several outcomes, including labor force participation, employment, and weeks worked.
+The joint model further indicates that the coefficients for women differ significantly between metro and non-metro samples, whereas for men no such differences are detected.
+One interpretation is that urban labor markets, combined with higher childcare costs, may amplify intra-household specialization, leading some women to reduce labor supply when partners experience greater employment stability.
+The metro-area analysis reinforces the gender differences observed in the main results: male extensive-margin gains remain robust, while female responses vary substantially by geography.
+5.5 Residence - Work Mismatch
+We next examine whether commuting patterns moderate the effects of anti-discrimination laws.
+In particular, we distinguish individuals who both live and work in the same county (“non-commuters”) from those who commute across county lines.
+Table 9 reports the estimates for non-commuters based on the fully interacted joint-estimation framework.
+Table 9: Effects of Anti-Discrimination Laws on Labor Supply: Analysis of Work & Home Residence
+(1) (2) (3) (4) (5) (6)
+Extensive Margin Intensive Margin
+Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
+Work–Residence County Matched (Same County)
+ENDA $\times$ SSP: Female -0.0020 -0.0041 1.3169∗∗ -0.2497 1.1835∗∗∗ -0.6671
+(0.0133) (0.0158) (0.5689) (0.7315) (0.3187) (0.4222)
+ENDA $\times$ SSP: Male 0.0262∗∗ 0.0231 0.5135 1.0258 -0.4456 -0.1132
+(0.0133) (0.0139) (0.5081) (0.6102) (0.3420) (0.3172)
+$p$: M – F (Matched) 0.0858 0.2049 0.2584 0.1563 0.0011 0.2927
+$p$: M (Matched – Mismatched) 0.0490 0.0960 0.0809 0.0271 0.2188 0.3734
+$p$: F (Matched – Mismatched) 0.8780 0.7938 0.0155 0.3999 0.0291 0.5921
+Observations 7,781,313 7,781,313 7,781,313 7,781,313 5,765,525 5,765,525
+Mean (F) 0.779 0.736 32.70 37.46 40.87 46.77
+Mean (M) 0.785 0.741 33.69 37.83 42.07 47.23
+  • Notes: Data are derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
+We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment.
+All regressions in this table are estimated jointly with a fully interacted model that allows ENDA effects to vary across four groups: men and women, each in the commuter subsample and in the non-commuter subsample.
+This framework nests the separate-by-sex estimations and enables formal tests of equality across groups.
+We report ENDA effects for the matched sample (work county = residence county) and include p-values comparing these estimates to those obtained for individuals with a work–home county mismatch.
+Columns (1–2) focus on the extensive margin, examining labor force participation and employment.
+Columns (3–6) analyze intensive-margin outcomes, usual weekly hours and weeks worked, where Columns (5–6) restrict the sample to employed individuals.
+The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples in the matched sample, respectively.
+Standard errors are clustered at the county level.
+$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
+Among non–commuters, the qualitative patterns for men in same–sex couples are broadly consistent with the main results: ENDA generates positive and statistically significant effects on labor force participation and employment (Columns 1–2), while intensive–margin effects are more modest.
+For women, extensive–margin coefficients remain close to zero, but we find sizable and statistically significant increases in usual weekly hours.
+The joint model shows that, within the non–commuter subsample, gender differences on the extensive margin are at most marginally significant, whereas women exhibit significantly larger ENDA effects than men on weekly hours among the employed.
+These marginal or insignificant gender differences likely reflect reduced precision due to the substantial sample restriction rather than substantive shifts in underlying patterns.
+The pp–values comparing non–commuters to the commuter sample indicate that men’s ENDA effects change somewhat across commuting regimes on both the extensive and intensive margins, but the signs and overall pattern remain stable.
+For women, by contrast, the intensive–margin responses in hours differ more sharply between commuters and non–commuters.
 5.6 Reporting Changes
 Prior research has highlighted higher rates of misclassification errors before 2008, where different-sex couples were sometimes incorrectly identified as same-sex couples due to typos in the sex variable.
-In response, and considering the 2008 ACS redesign aimed at reducing such errors (O’Connell and Lofquist (2009)), we restrict our sample to 2008–2019 to test the robustness of our findings reported in Table 4.
-Table 9 presents the estimates based on this restricted sample.
-For men in same-sex partnerships, the results remain qualitatively similar to our full sample analysis.
-The estimated impacts on both labor force participation and employment continue to be positive and statistically significant, with coefficients that are comparable in magnitude to those in Table 4.
-Among women, the extensive margin estimates remain small and statistically insignificant, while the intensive margin results are largely consistent with the full sample, albeit with some variations in magnitude.
-These findings suggest that potential misclassification errors in earlier years do not drive our main conclusions.
-The consistency of the results between the full sample (2005–2019) and the restricted sample (2008–2019) reinforces the robustness of our empirical strategy and the validity of our conclusions regarding the impact of anti-discrimination laws on labor market outcomes for same-sex couples.
-5.6 Weighting
-Although the ACS recommends applying sample weights to ensure representativeness, we conduct a robustness check by re‐estimating our models without these weights, following Solon et al.
-(2015).
-This approach allows us to verify that our main results are not driven solely by the weighting scheme.
-The results in Table 10 indicate that the overall impacts of anti‑discrimination laws on labor supply remain robust even when ACS weights are removed.
-For men in same‑sex partnerships, the magnitude of the estimated effects on both labor force participation and employment is slightly smaller compared to the weighted estimates, but they continue to be statistically significant.
-This suggests that the positive impact of these laws for men is not driven solely by the weighting procedure.
-For women, the pattern of the results is similar to the weighted sample, with some variations in magnitude across different labor outcomes.
-5.7 Placebo Tests and Alternative Estimators
-We next conduct a series of placebo tests and apply alternative estimators to assess the impact of anti‑discrimination laws on labor market outcomes among both same‑sex and different‑sex couples.
-Specifically, we utilize four distinct methods to assess the impact of anti‑discrimination laws on labor market outcomes among same‑sex and different‑sex couples: TWFE, Callaway and Sant’Anna (2021), Sun and Abraham (2021), and Borusyak et al.
-(2021).
-As a placebo test, we apply the DiD framework to different‑sex couples, who are not the target of these laws, and thus we expect no significant changes in their labor market outcomes.
-Panel (a) of Tables 11 and 12 report the DiD estimates for different‑sex couples, while Panels (a) and (c) of Figures 6 and 7 display the corresponding event study plots.
-Our placebo tests show that the DiD estimates yield no statistically significant impacts on labor force participation or employment rates for different‑sex couples.
-The pre‑ENDA coefficients are insignificant, confirming the absence of differential trends prior to the law’s enactment.
-Most post‑ENDA coefficients remain statistically insignificant, with only a few minor exceptions of small magnitude, largely consistent with a 5% false positive rate.
-Next, Panel (b) of Tables 11 and 12 presents the impact of anti‑discrimination laws on same‑sex couples, while Panels (b) and (d) of Figures 6 and 7 display the corresponding event studies.
-In these analyses, we compare labor market outcomes for same‑sex couples in regions that have enacted the laws with those in regions that have not.
-The results reveal that male same‑sex couples in areas with anti‑discrimination laws experience significant improvements in labor market outcomes, particularly through increased labor force participation and higher employment rates.
-For female same‑sex couples, however, we observe no significant effects on the extensive margin; instead, the positive and significant impact is confined to the intensive margin, as evidenced by an increase in average weekly hours worked.
-Hence, our alternative estimators yield similar point estimates and comparable confidence intervals, reinforcing the robustness of our findings.
+In response, and considering the 2008 ACS redesign aimed at reducing such errors (o2009changes), we restrict our sample to 2008–2019 to test the robustness of our findings reported in Table 4.
+Table 10 presents the estimates based on this restricted sample, using the joint estimation framework.
+Table 10: Effects of Anti-Discrimination Laws on Labor Supply: ACS Redesign (2008–2019)
+(1) (2) (3) (4) (5) (6)
+Extensive Margin Intensive Margin
+Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
+ACS Redesign Sample (2008–2019)
+ENDA $\times$ SSP: Female 0.0217 0.0272 1.3482∗∗ 0.6313 0.5888 -0.1525
+(0.0129) (0.0195) (0.6583) (0.7446) (0.4430) (0.4733)
+ENDA $\times$ SSP: Male 0.0459∗∗∗ 0.0492∗∗∗ 1.5759∗∗ 3.1075∗∗∗ -0.2548 0.8835∗∗∗
+(0.0140) (0.0156) (0.7054) (0.8309) (0.7122) (0.3363)
+$p$: M – F (ACS Redesign) 0.1635 0.3301 0.8272 0.0116 0.3877 0.0347
+$p$: M (ACS Redesign – Rest) 0.7712 0.6221 0.5605 0.0837 0.8041 0.0165
+$p$: F (ACS Redesign – Rest) 0.1277 0.0633 0.7380 0.1901 0.7515 0.0070
+Observations 7,778,320 7,778,320 7,778,320 7,778,320 5,763,114 5,763,114
+Mean (F) 0.824 0.787 34.27 39.72 41.09 47.51
+Mean (M) 0.837 0.802 35.66 40.48 42.41 48.11
+  • Notes: Data are derived from the American Community Survey (ACS) conducted between 2005 and 2019, following the major ACS redesign that altered survey implementation and imputation procedures.
+All regressions are estimated jointly using a fully interacted model that allows ENDA effects to vary across men and women within the redesign subsample.
+We report the ENDA effects estimated for the 2008–2019 sample and include $p$-values testing whether (i) male and female coefficients differ within the redesign sample, and (ii) redesign coefficients differ from those estimated in the 2005–2007 sample.
+Columns (1–2) report extensive-margin outcomes (labor force participation and employment).
+Columns (3–6) report intensive-margin outcomes (usual weekly hours and weeks worked), with Columns (5–6) restricted to employed individuals.
+The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples in the ACS redesign sample (2008-2019), respectively.
+Standard errors are clustered at the county level.
+$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
+For men in same-sex couples, the results remain qualitatively similar to our full-sample analysis: ENDA continues to generate positive and statistically significant effects on labor force participation and employment, with magnitudes close to those in Table 4.
+Among women, extensive-margin coefficients remain small, with the labor-force estimate marginally significant, and the employment estimate statistically insignificant.
+The intensive-margin results are broadly consistent with the full sample, though with some variation in magnitude.
+The joint model also allows us to test whether the 2008–2019 coefficients differ from those estimated in the 2005–2007 sample.
+For men, most outcomes do not differ significantly across samples, though the effect on weeks worked does.
+For women, we detect significant differences in employment and in weeks worked among the employed.
+The male–female comparison on the extensive margin is not statistically significant in the ACS Redesign subsample, although the differences remain qualitatively similar to those in the full sample.
+Taken together, these results indicate that potential misclassification errors in earlier ACS waves do not drive our main conclusions.
+The consistency between the full and redesign-restricted samples reinforces the robustness of our empirical strategy and the validity of our conclusions regarding the impact of anti-discrimination laws on labor market outcomes for same-sex couples.
 6 Discussion
 6.1 Differences in Response by Sex
-The results reveal that anti-discrimination laws affect household labor allocation differently for same-sex couples by gender.
-Men in male same-sex partnerships experience a reduction in their labor supply gap relative to men in different-sex partnerships, while there are no significant average effects for women in female same-sex partnerships.
-One potential explanation is that changes in job security, stemming from reduced discrimination, alter the household’s allocation of labor between market and non‑market activities.
-As documented by Badgett (2006), reduced discrimination in the labor market improves employment prospects and job security for sexual minorities.
-With greater job security for the primary earner, the secondary earner may reduce market participation in favor of increased household production, including child rearing.
-This pattern of adjustment is consistent with models of household bargaining and collective decision‑making (see, e.g., Chiappori (1992) and Lundberg and Pollak (1993)), which predict that improvements in labor market conditions for one partner can induce a reallocation of tasks within the household.
-In this framework, the observed increase in intra‑household differences in hours worked may reflect a shift toward greater specialization, with one partner focusing more on market production while the other devotes more time to household responsibilities.
-Thus, the differential responses between male and female same‑sex partnerships could arise from distinct household preferences and constraints.
-Households consisting of women in same-sex partnerships typically have more children than those consisting of men in same-sex partnerships.
-Given this difference, it would be unsurprising to see female same-sex households shift more toward a specialization of household labor compared to male same-sex households.
-This difference in specialization could help explain why anti-discrimination laws have different effects on women in same-sex partnerships and men in same-sex partnerships.
-To test the different responses by men and women in same-sex partnerships to anti-discrimination laws, we implement an alternative difference-in-difference model.
-Instead of comparing those in same-sex partnerships to those in different-sex partnerships, we compare men and women in same-sex partnerships to examine their differential responses at the household level.
-Do women in same-sex partnerships change their household division of labor relative to men in same-sex partnerships following anti-discrimination laws?
-To answer this question, we collapse the data to the household level and examine households instead of individuals.
-Specifically, we estimate this model:
-$Y_{jct}$ $=\alpha_{0}\,ENDA_{ct}+\alpha_{1}\,FemSSP_{j}+\alpha_{2}\,\bigl{(}ENDA_{ct}\times FemSSP_{j}\bigr{)}+\mu_{c}+\delta_{t}+\epsilon_{jct}$ (3)
-The notation is the same as before with two main differences.
-First, our interaction term uses an indicator for whether household $j$ is a female same‐sex partnership, denoted by $FemSSP_{j}$.
-Second, $Y_{jct}$ now represents outcomes measured at the household level.
-The coefficient $\alpha_{2}$ is our parameter of interest and captures how female same‐sex households differ in their response to anti‑discrimination laws relative to male same‑sex households.
-The first outcome examined is an indicator for whether the household is a one‑earner family.
-We also analyze the intra‑household division of labor by measuring the absolute difference in weekly hours worked between partners, thereby capturing both the extensive and intensive margins of labor supply.
-Finally, we explore the potential effects on child rearing by examining both the presence of children in the household and the number of children present.
-We present the results in Table 13.
-Anti-discrimination laws do not affect the likelihood of same-sex partnerships becoming a one-earner household, but there is an effect on the intensive margin.
-There is a significant effect of anti-discrimination laws in the intra-household difference in hours worked for male same-sex and female same-sex households.
-An increase in the difference of hours worked within the household could suggest a greater specialization in market production for one person and a greater specialization in household production and working fewer hours for the other person in the partnership.
-Female same-sex households do see a significant increase in the likelihood of having any child and the number of children relative to male same-sex households.
-This evidence suggests that men and women in same-sex partnerships are responding differently to these anti-discrimination laws.
-Female same-sex households begin to have more children than male same-sex households.
-It may appear in the main results that anti-discrimination laws do not help women in female same-sex partnerships.
-However, their response in the labor market could be driven by a greater desire for children and subsequent changes in household labor allocation due to an increase in job security.
-The corresponding changes in the labor market could represent an increase in welfare for female same-sex households if they value their new children more than their lost hours worked and subsequently lost wages.
-Consistent with the theory and empirical findings, female same-sex households could become more secure with one woman being the primary earner following anti-discrimination laws.
+The estimates reveal meaningful heterogeneity by sex within same-sex partnerships.
+For men in male same-sex couples, ENDA adoption significantly reduces the labor supply differential relative to men in different-sex partnerships.
+By contrast, we do not detect statistically significant average effects for women in female same-sex partnerships.
+Importantly, the absence of an average effect for women does not imply that ENDAs only affect men or that a single mechanism must operate symmetrically across groups.
+Our heterogeneity tests show that the extensive-margin responses for men and women differ significantly: men exhibit clear increases in labor supply, whereas the corresponding female estimates are near zero (and negative in metro areas).
+The appropriate interpretation is therefore one of heterogeneous adjustment, not opposing effects from a unified mechanism.
+Men respond on labor-market margins, while women appear to adjust through channels outside those margins, such as fertility decisions, with subsequent evidence also showing increased specialization for both male and female households.
+A natural explanation draws on household models of specialization.
+If ENDAs reduce discrimination and increase job security for sexual minority workers, households may respond by reallocating labor between market and non-market activities.
+Prior work documents that reduced discrimination can strengthen the position of the primary earner within sexual minority households (badgett2006discrimination).
+In collective and bargaining models of the household, an improvement in one partner’s labor market opportunities increases the returns to specialization: the partner with stronger market prospects supplies more labor, while the other reallocates time toward household production (chiappori1992collective; lundberg1993separate).
+Such mechanisms can generate heterogeneous responses across male and female same-sex households.
+Female same-sex couples are substantially more likely to have children than male same-sex couples, and thus may be more sensitive to changes in expected job stability or income that facilitate childcare.
+A shift toward specialization in female households may therefore involve adjustments in fertility and household production, rather than intensive-margin labor supply.
+To test these mechanisms directly, we estimate a household-level difference-in-differences model comparing male and female same-sex partnerships:
+$Y_{jct}$ $=\alpha\;+\;\beta_{1}\bigl[\mathrm{FemSSP}_{j}\times\mathrm{ENDA}_{ct}\bigr]\;+\;\beta_{2}\,\mathrm{FemSSP}_{j}\;+\;\beta_{3}\,\mathrm{ENDA}_{ct}+X_{jct}^{\prime}\delta+\mu_{c}+\delta_{t}+\epsilon_{jct}$ (3)
+Here $\beta_{1}$ captures how female same-sex households respond relative to male same-sex households.
+We examine three sets of outcomes: (i) whether the household is a one-earner family, (ii) the absolute difference in weekly hours worked between partners (capturing intra-household specialization), and (iii) fertility outcomes.
+Table 11 shows that ENDAs do not affect the probability of being a one-earner household, but they significantly increase intra-household differences in hours worked for both male and female same-sex households.
+The interaction term is negative because male households experience a larger increase in specialization; combining the ENDA main effect and the interaction yields a positive, though smaller, specialization response for female households, consistent with the specialization hypothesis.
+Table 11: Responses between Male and Female Same-Sex Partnerships
+(1) (2) (3) (4)
+One Earner HH Diff in Hours Worked Any Children Number of Children
+ENDA $\times$ Female 0.0107 -1.0573 0.0338∗∗∗ 0.0562
+(0.0086) (0.5452) (0.0121) (0.0303)
+ENDA -0.0201 1.6692∗∗∗ -0.0056 -0.0289
+(0.0133) (0.4773) (0.0148) (0.0329)
+Female 0.0029 -0.8066∗∗∗ 0.1428∗∗∗ 0.2273∗∗∗
+(0.0061) (0.2587) (0.0068) (0.0157)
+Observations 41,382 41,382 41,382 41,382
+Mean (F) 0.2585 26.8599 0.3460 0.6036
+$R^{2}$ 0.0534 0.1349 0.1306 0.1216
+  • Notes: Data is derived from the Annual Community Surveys (ACS) spanning 2005-2019 comparing women in same-sex partnerships to men.
+The first column examines if the household has only one earner.
+The second column tests the difference in absolute terms of hours worked between the two partners.
+The third and fourth columns examine how households differ in having children.
+The coefficients show the differential effect of anti-discrimination between female and male same-sex households.
+The “Mean (F)” row reports pre-treatment averages of the outcome variables (including never-treated counties) for female same-sex couple households.
+Standard errors are clustered at the county level.
+$p<0.10$, ∗∗ $p<0.05$, $^{{}^{***}}$ $p<0.01$.
+Female same-sex households also exhibit significant increases in fertility, both in the likelihood of having any child and in the number of children, relative to male households.
+These results suggest that ENDA-induced improvements in job security may allow female same-sex households to realize fertility preferences and subsequently adjust labor supply to accommodate additional childcare needs.
+Taken together, these findings show that male and female same-sex households respond differently to anti-discrimination laws not because the laws are ineffective for women, but because the relevant channels of adjustment differ.
+Male same-sex households respond primarily through increased labor-market specialization, while female same-sex households adjust fertility and household production.
+If female households value these additional children more than forgone labor market hours, these responses may reflect higher welfare even in the absence of average improvements in women’s labor market outcomes.
+6.2 Sorting and Partnership Formation
+A remaining concern is whether ENDA adoption changes who appears in the same-sex partnered sample.
+If ENDAs affect partnership formation, cohabitation decisions, or relationship stability, then the composition of same-sex partnerships could shift endogenously with treatment.
+Such changes would complicate interpretation of the triple-differences estimates, since the observed sample would no longer represent the same underlying population before and after ENDA adoption.
+To assess this possibility, we examine whether the probability of being observed in a same-sex partnership changes following ENDA adoption, and we additionally test for sorting on observable demographics.
+Specifically we estimate:
+$\mathrm{Y}_{ict}$ $=\beta\,\mathrm{ENDA}_{ct}+\mu_{c}+\delta_{t}+\epsilon_{ict},$ (4)
+where $\mathrm{Y}_{ict}$ is an indicator for living in a same-sex partnership, age, or years of education.
+$\mu_{c}$ are county fixed effects, and $\delta_{t}$ are year fixed effects.
+Our core specification tests whether the fraction of adults observed in same-sex partnerships shifts systematically in treated counties relative to untreated counties.
+Using age and years of education as outcomes provides a direct check for demographic sorting: if ENDAs systematically attract (or repel) younger or more educated individuals, we would observe corresponding shifts in treated counties.
+Table 12 reports the results.
+Across all specifications, and separately for men and women, we find no evidence that ENDAs affect the likelihood of living in a same-sex partnership.
+Coefficients are small, precisely estimated, and statistically insignificant.
+These findings indicate that ENDA adoption does not meaningfully alter partnership formation or cohabitation patterns.
+Table 12: Sorting and Reporting
+(1) (2) (3) (4) (5)
+All SSP Male SSP Female SSP Age Years of Education
+ENDA -0.0002 0.0004 -0.0008 0.0257 -0.0212
+(0.0005) (0.0005) (0.0005) (0.0648) (0.0182)
+Observations 16,071,420 7,850,821 8,220,599 16,071,420 16,071,420
+Mean 0.0093 0.0087 0.0099 45.5421 13.6441
+$R^{2}$ 0.0043 0.0081 0.0024 0.0101 0.0467
+  • Notes: Data comes from the 2005-2019 yearly ACS seeing how the number of same-sex partnerships change in a county after the passage of an anti-discrimination law.
+The first column looks at both men and women with the next two columns separating the sexes.
+Column (4) investigates age and (5) looks at years of education.
+The “Mean” row reports pre-treatment averages of the outcome variables (including never-treated counties) for same-sex couples.
+Standard errors are clustered at the county level.
+$p<0.10$, ∗∗ $p<0.05$, $^{{}^{***}}$ $p<0.01$.
+The absence of sorting responses strengthens the interpretation of the main results.
+Because ENDAs do not appear to change who enters or exits same-sex partnerships, the coupled-sample restriction does not generate treatment-induced compositional bias.
+Thus, the heterogeneous labor supply responses documented above are best understood as behavioral adjustments within stable households, rather than as artifacts of changes in relationship structure or partner selection.
 6.3 Mechanisms
-The results of this research indicate that sexual orientation anti-discrimination laws could significantly impact the labor market gaps for men in male same-sex partnerships.
-It is possible that these laws are binding and eliminate significant discrimination in the labor market.
-However, it is relatively easy for an employer to create a reason to fire an employee rather than firing someone specifically for being gay.
-It is also difficult to accurately determine someone’s sexual orientation by looking at them, unlike race and sex.
-Although it may be easier to determine someone’s sexual orientation conditional on them being in a same-sex relationship.
-Sexual orientation anti-discrimination laws may not be binding in a meaningful way that eliminates discrimination.
+The results suggest that sexual-orientation anti-discrimination laws can meaningfully reduce labor-market gaps for men in male same-sex couples.
+One interpretation is that these laws are binding and help limit discrimination.
+At the same time, enforcing such protections is fundamentally different from enforcing protections based on race or sex, since sexual orientation is not always observable.
+These laws may become more binding when individuals are in publicly identifiable same-sex relationships, but less so otherwise.
+[^16] This idea is consistent with research on the saliency of protected class membership, such as armour2018disability, which shows that anti-discrimination protections are more effective when group membership is more readily perceived.
+Thus, while the evidence points to important effects for male same-sex couples, the extent to which anti-discrimination laws eliminate broader forms of sexual-orientation discrimination remains uncertain.
 One mechanism that the laws could be affecting labor market outcomes is through increased positive sentiment toward LGB workers.
@@
 6.4 Threats to External Validity
-One limitation of this research is that we are unable to view all LGB workers.
-The analysis is specifically conducted on individuals in same-sex partnerships and different-sex partnerships.
-The data does not allow for the identification of single LGB workers, whose outcomes may be significantly different than LGB workers in partnerships.
-Another pitfall to using partnerships to infer sexual orientation is that one could incorrectly infer someone’s sexual orientation from a partnership.
-Bisexual individuals represent a larger group than lesbian or gay individuals ((Gates, 2011)), and our specification erroneously classifies bisexual people in different-sex relationships as heterosexual.
-This misclassification biases estimates toward zero, as bisexual workers face significant wage penalties ((Mize, 2016)) and their inclusion in the reference group depresses the heterosexual average.
-Future work must disentangle these groups to avoid obscuring inequities.
-Another potential concern in using partnerships to infer sexual orientation is that people in a same-sex partnership may be the group that is most at risk of discrimination.
-It seems likely that single LGB workers can more plausibly stay in the closet with their co-workers compared to their counterparts in a same-sex partnership.
-This analysis may be capturing the effect of anti-discrimination laws on the group that is most likely to be affected.
-These results may have a larger effect size than what one would find in examining the generalized LGB population.
-These biases are for the broader LGB population and external validity.
-Notably, these potential biases do not affect internal validity.
-Translating the effect of anti-discrimination laws for people in same-sex partnerships to all other LGB workers is not immediately obvious.
-It is dependent on several factors that are unknowable in this analysis such as the wages and labor supply for single LGB workers and bisexual workers in different-sex partnerships and how discrimination affects LGB people for those in same-sex partnerships.
-Furthermore, it is important to emphasize that this study cannot identify transgender individuals in the ACS, nor can it estimate the impact of anti-discrimination laws specifically protecting gender minorities.
-This restricts the ability to generalize findings to the entire LGBTQ+ community.
+A primary limitation of this study is that we cannot observe the full LGB population.
+The analysis compares individuals in same-sex partnerships to those in different-sex partnerships, but the ACS does not identify single LGB workers, whose labor market outcomes may differ systematically from partnered LGB individuals.
+Partnership status is also an imperfect proxy for sexual orientation: bisexual individuals in different-sex relationships are classified as heterosexual, and some gay or lesbian workers may similarly appear in the control group.
+Because lesbian women tend to earn more than heterosexual women, gay men tend to earn less than heterosexual men, and bisexual workers earn less than both (mize2016sexual), such misclassification does not shift group means in a uniform direction.
+While the bias in level pay gaps is generally toward zero, the sign of the bias in the estimated ENDA coefficients is theoretically ambiguous—especially if bisexual workers respond differently to anti-discrimination protections than gay or lesbian workers.
+Misclassification that places gay or lesbian individuals into the control group reduces the contrast between groups and thus produces the usual attenuation mechanism.
+These issues highlight the need for future work that can distinguish gay, lesbian, and bisexual workers more precisely to avoid masking heterogeneity across subpopulations.
+A related concern is that individuals in same-sex partnerships may constitute the subset of LGB workers most vulnerable to discrimination.
+Single LGB workers may be more able to conceal their sexual orientation in the workplace, whereas partnered individuals face greater visibility.
+The effects documented here may therefore reflect treatment effects for the group most exposed to discrimination, potentially overstating impacts for the broader LGB population.
+Crucially, these limitations pertain to external validity rather than internal validity.
+Misclassification of sexual orientation does not affect the within-county, within-sex comparisons that identify $\beta_{1g}$, but it does restrict the generalizability of the results to groups not directly observed in the data—such as single LGB workers or bisexual individuals in different-sex partnerships.
+How ENDA protections translate to these populations depends on their baseline outcomes, patterns of disclosure, and exposure to discrimination, none of which can be observed in the ACS.
+Finally, the ACS does not allow us to identify transgender or nonbinary individuals, nor does it permit analysis of anti-discrimination laws that specifically protect gender minorities.
+This further limits generalizability to the broader LGBTQ+ community.
 7 Conclusion
-This analysis is the first quasi-experimental study examining the impact of both local and state anti-discrimination laws on labor supply and wage disparities between workers in same-sex and different-sex partnerships.
-Utilizing a novel city-level panel dataset capturing the granular detail of anti-discrimination laws, we find that these laws significantly reduce sexual orientation-based inequalities in labor market outcomes.
-Specifically, anti-discrimination laws decrease the labor force participation difference by 3.8 percentage points (78%) and employment difference by 3.7 percentage points (77%) for men in male same-sex partnerships, with insignificant effect for women in female same-sex partnerships.
-We also document important wage effects.
-Sexual orientation anti-discrimination laws lead to significant wage gains for men in same-sex partnerships, with their percentile rank in the hourly wage and annual earnings distribution improving on average by approximately 2.7 and 3 percentiles, respectively.
-Moreover, these laws increase the probability of ranking above the 25th percentile in hourly wages by 3.6 percentage points and above the 50th percentile by 3.4 percentage points, with similar improvements observed in the annual earnings distribution.
-By contrast, the wage distribution for women in same-sex partnerships remains largely unaffected by these laws.
-We identify increased positive societal sentiments toward LGB Americans, reflected in growing support for same-sex marriage, as a potential mechanism through which these laws improve labor market outcomes.
-Furthermore, we explain differential responses to these laws between male and female same-sex households using Becker’s household specialization model and provide empirical support for this explanation.
-Prior to the landmark Supreme Court decision in Bostock v.
-Clayton County (2020), many states permitted employers to fire employees based solely on sexual orientation.
-While Bostock extended federal protections under Title VII to LGBTQI+ workers, legislative momentum for anti-discrimination laws at the state level has slowed significantly since then.
-The Equality Act, reintroduced in the 118th Congress on June 21, 2023, seeks to expand federal civil rights laws to explicitly include sexual orientation and gender identity.
-As of March 2025, the bill is pending in the Senate, facing significant opposition from some senators, particularly concerning its implications for religious freedoms and women’s rights in certain contexts.
-At the state level, legislative momentum for anti-discrimination laws has slowed.
-Between 2009 and 2019, more states barred cities from protecting their residents from discrimination than those that expanded protections.
-Additionally, recent developments have seen states like Arkansas and Tennessee actively prohibiting localities from enacting their own anti-discrimination measures.
-Our research provides critical and timely insights into the impact of anti-discrimination laws on labor market outcomes for individuals in same-sex partnerships.
-By offering a comprehensive evaluation of these laws’ effects, our findings meaningfully inform ongoing policy discussions and judicial decisions, including deliberations by the Supreme Court of the United States, about extending employment protections to the LGB community.
-Appendix A Additional Robustness Checks
-A.1 Event Studies
-Figures A1, A2, A3, A4, A5 present the event study regressions corresponding to the robustness checks discussed in Sections 5.2, 5.3, 5.4, 5.5, 5.6.
-In each figure, we trace the evolution of labor supply impacts on the external margin relative to the enactment of anti-discrimination laws.
-Across all specifications, the pre-treatment coefficients are statistically insignificant, and the post-treatment dynamics closely mirror those observed in our main analyses.
-Specifically, estimates for women in metropolitan areas are smaller on average, unweighted models and those estimated before the legalization of same-sex marriage typically produce slightly lower coefficients, and both the post-2008 sample and non-movers yield somewhat larger effects.
-A.2 Same-Sex Marriage Fixed Effects
-The results in Table A1 reinforce the robustness of our primary findings for men.
-Although the extensive margin effects for men in same-sex partnerships are slightly lower (0.0267 for labor force participation and 0.0270 for employment) than in our primary specification, they remain statistically significant, with generally smaller intensive margin estimates.
-For women, the point estimates on the extensive margin are quite similar in Table A1 and Table 6.
-However, the employment effect for women turns negative and is marginally significant (–0.0212, p $<$ 0.10), which contrasts with the mostly null effects observed elsewhere, aside from the negative effects noted in metropolitan areas.
-This conflicting evidence suggests that further research is needed to better understand the impacts on women.
-Our hypothesis remains that high child-care costs in metropolitan areas may be an important contributing factor.
-A.3 Occupation Fixed Effects
-LGB workers tend to sort to different occupations and industries based on sex and sexual orientation [Black et al., 2007].
+The policy environment has evolved considerably in recent years.
+Prior to Bostock v.
+Clayton County (2020), many states allowed employees to be fired solely based on sexual orientation.
+Although Bostock extended federal Title VII protections to LGBTQI+ workers, new anti-discrimination statutes have stagnated at both the state and federal levels.
+The Equality Act, reintroduced in the 118th Congress on June 21, 2023, remains pending in the Senate, with critics voicing concerns related to religious liberties and women’s rights.
+Meanwhile, more states have chosen to preempt or restrict local anti-discrimination ordinances than those that have adopted new protections, reflecting persistent polarization.
+Against this backdrop, our study offers quasi-experimental evidence on how local and state-level anti-discrimination laws shape labor market outcomes for individuals in same-sex couples.
+Exploiting a novel city-level panel dataset that captures the granular evolution of these laws, we document significant reductions in sexual orientation-based labor market disparities.
+In particular, for men in male same-sex couples, our estimates indicate a 3.8 percentage-point (78%) decrease in the labor force participation gap and a 3.7 percentage-point (77%) decline in the employment gap relative to different-sex counterparts, whereas we find no significant changes for women in female same-sex couples.
+Beyond labor supply, our results reveal meaningful wage improvements for men in same-sex couples.
+Their average percentile rank in hourly wages and annual earnings increases by about 2.6 and 2.9 points, respectively, with a 3.7 percentage-point increase in the likelihood of surpassing the 25th wage percentile and a 3.3 percentage-point boost for exceeding the 50th percentile.
+These outcomes underscore the capacity of anti-discrimination laws to mitigate wage gaps rooted in sexual orientation, with no comparable impact observed for women in female same-sex couples.
+We identify increased positive societal sentiments toward LGB Americans, reflected in growing support for same-sex marriage, as a potential mechanism by which these laws improve labor market outcomes.
+Taken together, our results suggest that local and state-level anti-discrimination measures can significantly mitigate labor market disadvantages faced by men in same-sex couples.
+By illustrating how these laws reduce labor supply and wage gaps, we offer timely insight into ongoing policy debates and judicial deliberations about extending employment protections to sexual minorities, paving the way for future research on the broader implications of such measures.
+Appendix A County Identification from PUMAs
+The ACS reports Public Use Microdata Areas (PUMAs) rather than counties.
+We assign respondents to counties using the IPUMS PUMA–county crosswalk.
+In practice, metro PUMAs map nearly one-to-one to counties: 475 counties in our sample are uniquely identified.
+Rural PUMAs aggregate multiple counties within a state; in these cases, we follow ACS public-use geography and treat each rural PUMA as a single county-group.
+This yields 48 county-groups, for a total of 523 county (or county-group) units.
+All ENDA codes are constructed at the county (or county-group) $\times$ year level using this mapping.
+Appendix B Using Only State Laws
+Local ordinances provide meaningful identifying variation.
+Among the 332 counties in our sample that ever receive ENDA-type protection, 133 (about 40 percent) adopt a county-level ordinance before their state enacts a statewide ENDA.
+A state-only measure would classify all residents in these county–years as untreated, which affects roughly 30 percent of post-treatment ACS person-year observations.
+Ignoring county ordinances therefore induces non-classical measurement error in the treatment variable and mechanically attenuates the estimated effects.
+Table B.1 illustrates this attenuation by removing local-law variation.
+In Panel A, which replaces county fixed effects with state fixed effects, point estimates shrink relative to our baseline (Table 4), and, importantly, the male–female differences on the extensive margin become statistically insignificant.
+Panel B, which keeps the original fixed-effects structure but forces ENDA to vary only at the state level, shows a similar pattern: some effects shrink towards zero, and the male–female gap is less precisely estimated.
+Table B.1: Effect of Anti-Discrimination Laws: Extensive & Intensive Margin of Labor Supply (State Laws)
+(1) (2) (3) (4) (5) (6)
+Extensive Margin Intensive Margin
+Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
+Panel A: State-Fixed Effects
+ENDA $\times$ SSP: Female 0.0040 -0.0089 1.5019∗∗ -0.3407 1.0750∗∗∗ -1.4526∗∗∗
+(0.0136) (0.0189) (0.5864) (0.7646) (0.2932) (0.3630)
+ENDA $\times$ SSP: Male 0.0270∗∗ 0.0243 0.2947 -0.1546 -0.2827 -1.0244∗∗∗
+(0.0125) (0.0125) (0.3015) (0.5411) (0.3488) (0.3585)
+$p$-value: M - F 0.2831 0.2062 0.0793 0.8527 0.0053 0.3204
+Panel B: County-Fixed Effects
+ENDA $\times$ SSP: Female -0.0067 -0.0158 1.5380∗∗∗ -0.3971 1.2360∗∗∗ -1.3461∗∗∗
+(0.0106) (0.0140) (0.5051) (0.6383) (0.3126) (0.3535)
+ENDA $\times$ SSP: Male 0.0326∗∗ 0.0321 0.6748∗∗ 0.2175 -0.2064 -0.7743∗∗
+(0.0153) (0.0169) (0.3292) (0.5922) (0.5124) (0.3419)
+$p$-value: M - F 0.0294 0.0754 0.1967 0.5485 0.0117 0.1645
+Employed Only X X
+Observations 7,781,358 7,781,358 7,781,358 7,781,358 5,765,580 5,765,580
+Mean (F) 0.818 0.782 34.42 39.51 41.25 47.27
+Mean (M) 0.825 0.790 35.53 39.96 42.41 47.69
+  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
+We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment.
+For this table, we abstract from local laws and use only state-level ENDA variation.
+The regressions are estimated jointly, allowing the treatment effect to differ for men and women through interaction terms.
+Columns (1–2) focus on the extensive margin, examining changes in labor force participation and employment.
+Columns (3–6) explore the intensive margin using variation in weeks worked per year and usual weekly hours; Columns (5–6) restrict the sample to employed individuals.
+The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples, respectively.
+Reported $p$-values test whether the ENDA effect differs significantly between men and women.
+Panel A replaces county fixed effects with state fixed effects in Equation (1), while Panel B reports the original specification.
+Standard errors are clustered at the state-level.
+$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
+These results confirm that omitting local ordinances discards meaningful within-state variation and leads to substantial attenuation of the estimated ENDA effects, particularly on the extensive margin where the gender differences weaken considerably.
+Appendix C Additional Robustness Checks
+C.1 Placebo Tests and Alternative Estimators
+We next conduct a series of placebo tests and apply alternative estimators to assess the impact of anti‑discrimination laws on labor market outcomes among both same‑sex and different‑sex couples.
+Specifically, we utilize four distinct methods to assess the impact of anti‑discrimination laws on labor market outcomes among same‑sex and different‑sex couples: TWFE, callaway2021difference, sun2021estimating, and borusyak2021revisiting.
+As a placebo test, we apply the DiD framework to different‑sex couples, who are not the target of these laws, and thus we expect no significant changes in their labor market outcomes.
+Panel (a) of Tables C.1 and C.2 report the DiD estimates for different‑sex couples, while Panels (a) and (c) of Figures C.1 and C.2 display the corresponding event study plots.
+Our placebo tests show that the DiD estimates yield no statistically significant impacts on labor force participation or employment rates for different‑sex couples.
+The pre‑ENDA coefficients are insignificant, confirming the absence of differential trends prior to the law’s enactment.
+Most post‑ENDA coefficients remain statistically insignificant, with only a few minor exceptions of small magnitude, largely consistent with a 5% false positive rate.
+Next, Panel (b) of Tables C.1 and C.2 presents the impact of anti‑discrimination laws on same‑sex couples, while Panels (b) and (d) of Figures C.1 and C.2 display the corresponding event studies.
+In these analyses, we compare labor market outcomes for same‑sex couples in regions that have enacted the laws with those in regions that have not.
+The results reveal that male same‑sex couples in areas with anti‑discrimination laws experience significant improvements in labor market outcomes, particularly through increased labor force participation and higher employment rates.
+For female same‑sex couples, however, we observe no significant effects on the extensive margin; instead, the positive and significant impact is confined to the intensive margin, as evidenced by an increase in average weekly hours worked.
+Hence, our alternative estimators yield similar point estimates and comparable confidence intervals, reinforcing the robustness of our findings.
+((a)) Labor Force Participation (Female Different-Sex Couples)
+((b)) Labor Force Participation (Female Same-Sex Couples)
+((c)) Labor Force Participation (Male Different-Sex Couples)
+((d)) Labor Force Participation (Male Same-Sex Couples)
+Figure C.1: Impacts on Labor Force Participation by Partnership Type
+Notes: Event study plots on the effects of ENDAs on people in same-sex couples and different-sex couples broken down by sex following the county-level and state-level anti-discrimination laws.
+Coefficients with 95% confidence intervals.
+Standard errors are clustered at the county level.
+Reference year is -4.
+((a)) Employment (Female Different-Sex Couples)
+((b)) Employment (Female Same-Sex Couples)
+((c)) Employment (Male Different-Sex Couples)
+((d)) Employment (Male Same-Sex Couples)
+Figure C.2: Impacts on Employment by Partnership Type
+Notes: Event study plots on the effects of ENDAs on people in same-sex couples and different-sex couples broken down by sex following the county-level and state-level anti-discrimination laws.
+Coefficients with 95% confidence intervals.
+Standard errors are clustered at the county level.
+Reference year is -4.
+Table C.1: Effect of Anti-Discrimination Laws on Men in Same-Sex and Different-Sex Partnerships
+(1) (2) (3) (4) (5) (6)
+Extensive Margin Intensive Margin
+ENDA $\times$ SSP Estimates Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
+Panel A: Men in Different-Sex Partnerships
+TWFE 0.0004 -0.0002 0.0629 0.0642 -0.0022 -0.0233
+(0.0015) (0.0018) (0.0834) (0.0809) (0.0602) (0.0448)
+callaway2021difference 0.0008 -0.0012 0.1038 0.0596 0.0797 0.0016
+(0.0013) (0.0017) (0.0961) (0.0817) (0.0725) (0.0595)
+sun2021estimating 0.0008 -0.0005 0.1160 0.0827 0.0349 -0.0239
+(0.0015) (0.0019) (0.0852) (0.0911) (0.0570) (0.0496)
+borusyak2021revisiting 0.0023 0.0006 0.1898∗∗ 0.1345 0.0554 -0.0415
+(0.0014) (0.0020) (0.0879) (0.0958) (0.0552) (0.0465)
+Observations 3,763,442 3,763,442 3,763,442 3,763,442 3,046,376 3,046,376
+Employed Only X X
+Panel B: Men in Same-Sex Partnerships
+TWFE 0.0284∗∗ 0.0263 0.8832 1.1614 -0.1199 -0.2060
+(0.0137) (0.0144) (0.5618) (0.6476) (0.4351) (0.3773)
+callaway2021difference 0.0463∗∗ 0.0428∗∗ 1.2572∗∗ 1.3140 0.1736 -0.1863
+(0.0184) (0.0213) (0.6368) (0.8613) (0.5461) (0.4090)
+sun2021estimating 0.0337∗∗ 0.0332∗∗ 0.9484 1.3165∗∗ -0.1495 -0.1319
+(0.0145) (0.0152) (0.5357) (0.6657) (0.3888) (0.3290)
+borusyak2021revisiting 0.0420∗∗∗ 0.0414∗∗∗ 1.3754∗∗∗ 1.7464∗∗∗ -0.0756 -0.1261
+(0.0129) (0.0137) (0.5001) (0.6246) (0.3827) (0.2779)
+Observations 35,194 35,194 35,194 35,194 27,504 27,504
+Employed Only X X
+  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
+We compare men in counties with ENDAs with those without within four years of ENDA enactment.
+The regressions are estimated separately by partnership: Panel A for men in different-sex partnerships, while Panel B for men in same-sex partnerships.
+The coefficients indicate the impact of anti-discrimination laws.
+The analysis is divided into two parts: Columns (1-2) focus on the extensive margin, examining changes in labor force participation and employment status as a result of the laws.
+Columns (3-6), on the other hand, explore the intensive margin by leveraging variations in the number of weeks worked per year and average weekly working hours.
+Standard errors are clustered at the county level.
+Results are shown using six different estimators: Two-Way Fixed Effects (TWFE), callaway2021difference, sun2021estimating, and borusyak2021revisiting.
+$p<0.10$, ∗∗ $p<0.05$, $^{{}^{***}}$ $p<0.01$.
+Table C.2: Effect of Anti-Discrimination Laws on Women in Same-Sex and Different-Sex Partnerships
+(1) (2) (3) (4) (5) (6)
+Extensive Margin Intensive Margin
+ENDA $\times$ SSP Estimates Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
+Panel A: Women in Different-Sex Partnerships
+TWFE 0.0010 0.0008 0.0093 0.0045 -0.0882 -0.1083∗∗
+(0.0019) (0.0017) (0.0781) (0.0756) (0.0566) (0.0485)
+callaway2021difference 0.0017 0.0020 -0.0127 -0.0105 -0.0693 -0.0602
+(0.0019) (0.0023) (0.1021) (0.1016) (0.0733) (0.0783)
+sun2021estimating 0.0010 0.0010 -0.0152 -0.0159 -0.0732 -0.0705
+(0.0017) (0.0018) (0.0710) (0.0830) (0.0618) (0.0602)
+borusyak2021revisiting 0.0015 0.0013 0.0057 -0.0130 -0.0637 -0.1021
+(0.0016) (0.0016) (0.0733) (0.0845) (0.0620) (0.0641)
+Observations 3,942,881 3,942,881 3,942,881 3,942,881 2,660,191 2,660,191
+Employed Only X X
+Panel B: Women in Same-Sex Partnerships
+TWFE 0.0099 0.0090 1.3686∗∗∗ 0.4999 0.5733 -0.4873
+(0.0092) (0.0105) (0.4386) (0.5327) (0.3579) (0.3743)
+callaway2021difference -0.0042 -0.0088 0.9758 -0.6803 0.8353 -1.2359∗∗
+(0.0120) (0.0125) (0.7503) (0.6246) (0.4648) (0.4891)
+sun2021estimating 0.0022 -0.0009 1.1645∗∗ -0.0565 0.5538 -0.8987∗∗
+(0.0103) (0.0109) (0.4920) (0.5785) (0.3148) (0.3529)
+borusyak2021revisiting 0.0000 -0.0013 1.0359∗∗ -0.2617 0.5542 -0.9571∗∗∗
+(0.0111) (0.0129) (0.4939) (0.5924) (0.2977) (0.3228)
+Observations 40,071 40,071 40,071 40,071 31,707 31,707
+Employed Only X X
+  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
+We compare women in counties with ENDAs with those without within four years of ENDA enactment.
+The regressions are estimated separately by partnership: Panel A for women in different-sex partnerships, while Panel B for women in same-sex partnerships.
+The coefficients indicate the impact of anti-discrimination laws.
+The analysis is divided into two parts: Columns (1-2) focus on the extensive margin, examining changes in labor force participation and employment status as a result of the laws.
+Columns (3-6), on the other hand, explore the intensive margin by leveraging variations in the number of weeks worked per year and average weekly working hours.
+Standard errors are clustered at the county level.
+Results are shown using six different estimators: Two-Way Fixed Effects (TWFE), callaway2021difference, sun2021estimating, and borusyak2021revisiting.
+$p<0.10$, ∗∗ $p<0.05$, $^{{}^{***}}$ $p<0.01$.
+C.2 Same-Sex Marriage Fixed Effects
+The results in Table C.3 reinforce the robustness of our primary findings for men.
+Although the extensive-margin effects for men in same-sex couples are slightly lower (0.0267 for labor force participation and 0.0270 for employment) than in our primary specification, they remain statistically significant, with generally smaller intensive-margin estimates.
+For women, the point estimates on the extensive margin are quite similar in Table C.3 and Table 4, but the employment effect turns negative and is marginally significant (0.0212-0.0212, $p<0.10$), which contrasts with the mostly null extensive-margin effects observed elsewhere, aside from the negative responses noted in metropolitan areas.
+Moreover, the joint model indicates that the ENDA effects on the extensive margin are significantly larger for men than for women, as reflected in the $p$-values reported in Columns (1) and (2).
+This conflicting evidence for women suggests that further research is needed to better understand the mechanisms at play; our hypothesis remains that high child-care costs in metropolitan areas may be an important contributing factor.
+Table C.3: Effects of Anti-Discrimination Laws on Labor Supply: Analysis Without Same-Sex Marriage Interactions
+(1) (2) (3) (4) (5) (6)
+Extensive Margin Intensive Margin
+Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
+ENDA $\times$ SSP: Female -0.0130 -0.0212 0.8477 -0.7081 0.9893∗∗∗ -0.7657∗∗
+(0.0108) (0.0115) (0.5167) (0.5762) (0.3083) (0.3165)
+ENDA $\times$ SSP: Male 0.0267∗∗ 0.0270∗∗ 0.5377 0.8852 -0.4480 -0.3068
+(0.0129) (0.0133) (0.4987) (0.6178) (0.3627) (0.2820)
+$p$-value: M - F 0.0042 0.0034 0.6404 0.0259 0.0035 0.2688
+Employed Only X X
+Observations 7,781,493 7,781,493 7,781,493 7,781,493 5,765,693 5,765,693
+Mean (F) 0.818 0.782 34.42 39.51 41.25 47.27
+Mean (M) 0.825 0.790 35.53 39.96 42.41 47.69
+  • Notes: Data are derived from the American Community Survey (ACS), 2005–2019.
+We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment.
+The regressions are jointly estimated, allowing ENDA effects to differ by sex.
+This table excludes the fully interacted same-sex-marriage term from Equation (1).
+Columns (1–2) report extensive-margin effects; Columns (3–6) report intensive-margin effects, with Columns (5–6) restricted to employed individuals.
+The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples, respectively.
+Reported $p$-values test equality of male and female coefficients.
+Standard errors are clustered at the county level.
+$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
+C.3 Occupation Fixed Effects
+LGB workers tend to sort to different occupations and industries based on sex and sexual orientation [black2007economics].
 They show that gay men tend to sort to more female-dominated fields and lesbian women sort to more male-dominated fields relative to their heterosexual counterparts.
 It is typical to not consider those factors because discrimination affects occupation and industry choice, which would make occupation and industry choice endogenous.
+Table C.4: Anti-Discrimination Laws on Labor Supply with Occupation FEs
+(1) (2) (3) (4) (5) (6)
+Extensive Margin Intensive Margin
+Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
+ENDA $\times$ SSP: Female -0.0064 -0.0132 0.5614 -0.8344 0.5854∗∗ -0.9524∗∗∗
+(0.0098) (0.0111) (0.4011) (0.4630) (0.2943) (0.3218)
+ENDA $\times$ SSP: Male 0.0228∗∗ 0.0246∗∗ 0.5170 0.8743∗∗ -0.1462 -0.0550
+(0.0093) (0.0100) (0.3722) (0.4336) (0.3306) (0.2799)
+$p$-value: M - F 0.0105 0.0075 0.9335 0.0013 0.1021 0.0237
+Employed Only X X
+Observations 7,781,358 7,781,358 7,781,358 7,781,358 5,765,580 5,765,580
+Mean (F) 0.818 0.782 34.42 39.51 41.25 47.27
+Mean (M) 0.825 0.790 35.53 39.96 42.41 47.69
+  • Notes: Data are from the American Community Survey (ACS), 2005–2019.
+We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment.
+Regressions are estimated jointly with four-digit occupation fixed effects, allowing ENDA effects to differ by sex through interaction terms.
+Columns (1–2) report extensive-margin outcomes (labor force participation and employment).
+Columns (3–6) report intensive-margin outcomes (usual weekly hours and weeks worked), with Columns (5–6) restricting to employed individuals.
+The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples, respectively.
+Reported $p$-values test equality of male and female coefficients.
+Standard errors are clustered at the county level.
+$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
+Table C.5: Anti-Discrimination Laws on Pay with Occupation FEs
+Hourly Wage Annual Earnings
+(1) (2) (3) (4) (5) (6) (7) (8)
+Percentile $\geq$25th $\geq$50th $\geq$75th Percentile $\geq$25th $\geq$50th $\geq$75th
+ENDA $\times$ SSP: Female -0.9026 -0.0024 -0.0182 -0.0116 -1.1521 -0.0018 -0.0274∗∗ -0.0059
+(0.6991) (0.0097) (0.0121) (0.0146) (0.6522) (0.0097) (0.0111) (0.0155)
+ENDA $\times$ SSP: Male 1.8922∗∗ 0.0243∗∗ 0.0244 0.0200 2.2207∗∗∗ 0.0242∗∗ 0.0391∗∗∗ 0.0308∗∗
+(0.9191) (0.0113) (0.0145) (0.0141) (0.8332) (0.0113) (0.0152) (0.0128)
+$p$-value: M - F 0.0204 0.0699 0.0508 0.0983 0.0017 0.0778 0.0015 0.0375
+Employed Only
+Observations 7,781,358 7,781,358 7,781,358 7,781,358 7,781,358 7,781,358 7,781,358 7,781,358
+Mean (F) 48.73 0.799 0.513 0.247 48.68 0.798 0.515 0.241
+Mean (M) 51.51 0.791 0.557 0.305 51.69 0.791 0.561 0.305
+  • Notes: Data are from the 2005–2019 ACS.
+We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment.
+Regressions are estimated separately by sex and include four-digit occupation fixed effects.
+Columns (1–4) report effects on hourly real wage percentiles; Columns (5–8) report effects on annual real earnings percentiles.
+Columns (2–4) and (6–8) estimate the probability of being above the 25th, 50th, and 75th percentiles.
+All wage and earnings variables are in 2019 dollars.
+The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples, respectively.
+Reported $p$-values test equality of male and female coefficients.
+Standard errors are clustered at the county level.
+$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
 Although controlling for occupation can pose selection concerns, because individuals not in the labor force typically lack occupation codes, we nonetheless examine within-occupation effects of anti-discrimination laws on wages and labor supply.
@@
 Nonetheless, by creating a distinct “no occupation” category, we preserve a measure of occupational affiliation for those who are employed while ensuring that non-employed individuals remain part of the sample.
-One potential mechanism is that workers in male same-sex partnerships may transition into more in-demand occupations, thereby mitigating relative labor supply and wage gaps.
-However, as shown in Tables A2 and A3, even after accounting for occupation, men in same-sex partnerships exhibit significant changes in both labor supply and wages.
+One potential mechanism is that workers in male same-sex couples may transition into more in-demand occupations, thereby mitigating relative labor supply and wage gaps.
+However, as shown in Tables C.4 and C.5, even after accounting for occupation, men in same-sex couples exhibit significant changes in both labor supply and wages.
 These findings suggest that the impact of anti-discrimination laws is not merely a by-product of occupational sorting.
 Nonetheless, caution is warranted when interpreting these results, given the inherent limitations of conditioning on occupation in a broader labor force context.
-Appendix Tables
-Table A1: Effects of Anti-Discrimination Laws on Labor Supply: Analysis of Same-Sex Marriage Fixed Effects
+C.4 Clustering
+Our baseline results report standard errors clustered at the county level, which is the level of treatment variation for local ENDA ordinances.
+As CameronMiller2015Practitioners emphasize, however, serial correlation at higher levels of aggregation can bias standard errors downward when treatment varies at both local and state levels.
+To address this concern, we re-estimate all specifications using (i) state-level clustering, and (ii) two-way clustering by state and year.
+Table C.6 shows that clustering at the state level produces narrower confidence intervals than the original county-clustered results.
+Two-way clustering by state $\times$ year yields standard errors nearly identical to state clustering.
+Across all outcomes, the qualitative patterns remain unchanged: men in same-sex partnerships exhibit positive and statistically significant extensive-margin responses to ENDA, while women’s responses are smaller and generally not statistically distinguishable from zero.
+Overall, the results are robust to alternative clustering choices, and inference is, if anything, more conservative under our original county-level specification.
+Table C.6: Effect of Anti-Discrimination Laws: Extensive & Intensive Margin of Labor Supply (Clustering)
+(1) (2) (3) (4) (5) (6)
+Extensive Margin Intensive Margin
+Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
+Panel A: Clustered by State
+ENDA $\times$ SSP: Female -0.0037 -0.0050 0.9519∗∗ -0.4143 0.7017∗∗∗ -0.8915∗∗∗
+(0.0100) (0.0141) (0.4116) (0.5690) (0.2426) (0.3293)
+ENDA $\times$ SSP: Male 0.0379∗∗∗ 0.0370∗∗∗ 1.1289∗∗∗ 1.5089∗∗∗ -0.0920 -0.0719
+(0.0106) (0.0096) (0.3741) (0.3729) (0.3834) (0.2740)
+$p$-value: M - F 0.0011 0.0191 0.7778 0.0086 0.1114 0.0365
+Panel B: Clustered by State-Year
+ENDA $\times$ SSP: Female -0.0037 -0.0050 0.9519 -0.4143 0.7017∗∗ -0.8915∗∗∗
+(0.0103) (0.0120) (0.4907) (0.5992) (0.3487) (0.3277)
+ENDA $\times$ SSP: Male 0.0379∗∗∗ 0.0370∗∗∗ 1.1289∗∗∗ 1.5089∗∗∗ -0.0920 -0.0719
+(0.0098) (0.0104) (0.4205) (0.5173) (0.4045) (0.2842)
+$p$-value: M - F 0.0059 0.0128 0.8063 0.0191 0.1434 0.0632
+Employed Only X X
+Observations 7,781,358 7,781,358 7,781,358 7,781,358 5,765,580 5,765,580
+Mean (F) 0.818 0.782 34.42 39.51 41.25 47.27
+Mean (M) 0.825 0.790 35.53 39.96 42.41 47.69
+  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
+We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment.
+The regressions are estimated jointly, allowing the treatment effect to differ for men and women through interaction terms.
+Columns (1–2) focus on the extensive margin, examining changes in labor force participation and employment.
+Columns (3–6) explore the intensive margin using variation in weeks worked per year and usual weekly hours; Columns (5–6) restrict the sample to employed individuals.
+The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples, respectively.
+Reported $p$-values test whether the ENDA effect differs significantly between men and women.
+Panel A clusters at the state level, Panel B at the state-year level.
+$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
+C.5 Weighting
+Although the ACS recommends applying sample weights to ensure representativeness, we conduct a robustness check by re‐estimating our models without these weights, following solon2015we.
+This approach allows us to verify that our main results are not driven solely by the weighting scheme.
+Table C.7: Effect of Anti-Discrimination Laws: Extensive & Intensive Margin of Labor Supply (Unweighted)
+(1) (2) (3) (4) (5) (6)
+Extensive Margin Intensive Margin
+Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
+ENDA $\times$ SSP: Female -0.0085 -0.0112 0.7029 -0.5207 0.6289∗∗ -0.7879∗∗
+(0.0102) (0.0118) (0.4868) (0.5804) (0.3008) (0.3315)
+ENDA $\times$ SSP: Male 0.0265 0.0251 0.8506 1.0802 -0.0666 0.0419
+(0.0144) (0.0149) (0.4873) (0.6126) (0.3540) (0.2871)
+$p$-value: M - F 0.0091 0.0345 0.7768 0.0102 0.0909 0.0344
+Employed Only X X
+Observations 7,781,358 7,781,358 7,781,358 7,781,358 5,765,580 5,765,580
+Mean (F) 0.818 0.782 34.42 39.51 41.25 47.27
+Mean (M) 0.825 0.790 35.53 39.96 42.41 47.69
+  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
+We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment.
+The regressions are estimated jointly, allowing the treatment effect to differ for men and women through interaction terms.
+Columns (1–2) focus on the extensive margin, examining changes in labor force participation and employment.
+Columns (3–6) explore the intensive margin using variation in weeks worked per year and usual weekly hours; Columns (5–6) restrict the sample to employed individuals.
+The reported $p$-values test whether the ENDA effect differs significantly between men and women.
+Standard errors are clustered at the county level.
+$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$.
+The results in Table C.7 indicate that the overall impacts of anti‑discrimination laws on labor supply remain robust even when ACS weights are removed.
+For men in same‑sex couples, the magnitude of the estimated effects on both labor force participation and employment is slightly smaller compared to the weighted estimates, but they continue to be statistically significant.
+This suggests that the positive impact of these laws for men is not driven solely by the weighting procedure.
+For women, the pattern of the results is similar to the weighted sample, with some variations in magnitude across different labor outcomes, though most effects remain small and statistically insignificant.
+C.6 Event Studies
+Figures C.3C.8 present the event study regressions corresponding to the robustness checks discussed in Sections 5.25.6 and Appendix C.5.
+In each figure, we trace the evolution of labor supply impacts on the external margin relative to the enactment of anti-discrimination laws.
+Across all specifications, the pre-treatment coefficients are statistically insignificant, and the post-treatment dynamics closely mirror those observed in our main analyses.
+Specifically, estimates for women in metropolitan areas are smaller on average, unweighted models and those estimated before the legalization of same-sex marriage typically produce slightly lower coefficients, and both the post-2008 sample and non-movers yield somewhat larger effects.
+((a)) Labor Force Participation
+((b)) Employment
+Figure C.3: Labor Supply Impacts on the Extensive Margin: Same-Sex Marriage Legalization
+Notes: Event study plots on the difference in labor supply between people in same-sex couples and different-sex couples broken down by sex following the county-level and state-level anti-discrimination laws.
+Sample is limited to observations before same-sex marriage legalization.
+Coefficients with 95% confidence intervals.
+Standard errors are clustered at the county level.
+Reference year is -4.
+((a)) Labor Force Participation
+((b)) Employment
+Figure C.4: Labor Supply Impacts on the Extensive Margin: Non-Movers
+Notes: Event study plots on the difference in labor supply between people in same-sex couples and different-sex couples broken down by sex following the county-level and state-level anti-discrimination laws.
+Sample is limited to individuals who have not moved in the past year.
+Coefficients with 95% confidence intervals.
+Standard errors are clustered at the county level.
+Reference year is -4.
+((a)) Labor Force Participation
+((b)) Employment
+Figure C.5: Labor Supply Impacts on the Extensive Margin: Metropolitan Areas
+Notes: Event study plots on the difference in labor supply between people in same-sex couples and different-sex couples broken down by sex following the county-level and state-level anti-discrimination laws.
+Sample is limited to individuals who live in metropolitan areas.
+Coefficients with 95% confidence intervals.
+Standard errors are clustered at the county level.
+Reference year is -4.
+((a)) Labor Force Participation
+((b)) Employment
+Figure C.6: Labor Supply Impacts on the Extensive Margin: Work Residence same as Home Residence
+Notes: Event study plots on the difference in labor supply between people in same-sex couples and different-sex couples broken down by sex following the county-level and state-level anti-discrimination laws.
+Sample is limited to individuals who live in the same county as they work in.
+Coefficients with 95% confidence intervals.
+Standard errors are clustered at the county level.
+Reference year is -4.
+((a)) Labor Force Participation
+((b)) Employment
+Figure C.7: Labor Supply Impacts on the Extensive Margin: 2008-2019
+Notes: Event study plots on the difference in labor supply between people in same-sex couples and different-sex couples broken down by sex following the county-level and state-level anti-discrimination laws.
+Sample is limited to cover 2008 through 2019.
+Coefficients with 95% confidence intervals.
+Standard errors are clustered at the county level.
+Reference year is -4.
+((a)) Labor Force Participation
+((b)) Employment
+Figure C.8: Labor Supply Impacts on the Extensive Margin: Weighting Schemes
+Notes: Event study plots on the difference in labor supply between people in same-sex couples and different-sex couples broken down by sex following the county-level and state-level anti-discrimination laws.
+We compare estimates weighted by ACS weights and without.
+Coefficients with 95% confidence intervals.
+Standard errors are clustered at the county level.
+Reference year is -4.
+C.7 Sensitivity Analysis for Parallel Trends
+To address potential violations of the parallel trends assumption, we conduct a sensitivity analysis following RambachanRoth2023.
+Instead of relying on potentially underpowered pre-trend tests (roth2022pretest), we construct robust confidence intervals using the “Relative Magnitude” restriction.
+This approach bounds the potential post-treatment bias to be no larger than $\bar{M}$ times the maximum observed pre-treatment trend deviation.
+We estimate 90% robust confidence intervals ($\alpha=0.1$) for the main effects reported in Columns (1-4) of Table 4 across sensitivity parameters $\bar{M}\in[0.1,2.0]$.
+As shown in Table C.8, male extensive margin estimates (Panel A) remain statistically significant up to $\bar{M}=0.4$, indicating robustness to moderate trend violations, while intensive margin effects are slightly more sensitive (robust up to $\bar{M}=0.3$).
+Conversely, estimates for women (Panel B) are statistically indistinguishable from zero across all sensitivity parameters, confirming the main null results hold even under relaxed assumptions.
+Table C.8: Sensitivity Analysis for Parallel Trends (HonestDiD)
+(1) (2) (3) (4)
+Extensive Margin Intensive Margin
+Sensitivity ($M$) Labor Force Employed Weekly Hours Weeks Worked
+Panel A: Men
+Original ($M=0$) [0.017, 0.059] [0.015, 0.059] [0.310, 1.948] [0.478, 2.540]
+$M=0.1$ [0.013, 0.062] [0.013, 0.061] [0.233, 2.044] [0.325, 2.693]
+$M=0.2$ [0.010, 0.065] [0.010, 0.064] [0.144, 2.155] [0.128, 2.916]
+$M=0.3$ [0.006, 0.069] [0.007, 0.067] [0.012, 2.308] [-0.205, 3.249]
+$M=0.4$ [0.001, 0.073] [0.001, 0.073] [-0.185, 2.526] [-0.600, 3.644]
+$M=0.5$ [-0.004, 0.078] [-0.005, 0.079] [-0.406, 2.768] [-1.005, 4.049]
+$M=0.6$ [-0.010, 0.084] [-0.012, 0.086] [-0.652, 3.015] [-1.447, 4.492]
+$M=0.7$ [-0.016, 0.090] [-0.019, 0.093] [-0.902, 3.266] [-1.872, 4.917]
+$M=0.8$ [-0.022, 0.096] [-0.026, 0.100] [-1.156, 3.521] [-2.335, 5.381]
+$M=0.9$ [-0.028, 0.102] [-0.033, 0.107] [-1.414, 3.780] [-2.781, 5.827]
+$M=1.0$ [-0.035, 0.108] [-0.039, 0.114] [-1.676, 4.043] [-3.208, 6.283]
+$M=1.5$ [-0.067, 0.140] [-0.076, 0.150] [-3.006, 5.378] [-5.512, 8.560]
+$M=2.0$ [-0.100, 0.172] [-0.113, 0.187] [-4.365, 6.741] [-7.794, 10.844]
+Panel B: Women
+Original ($M=0$) [-0.023, 0.015] [-0.023, 0.015] [-0.023, 0.015] [-0.023, 0.015]
+$M=0.1$ [-0.025, 0.017] [-0.025, 0.017] [-0.025, 0.017] [-0.025, 0.017]
+$M=0.2$ [-0.027, 0.019] [-0.027, 0.019] [-0.027, 0.019] [-0.027, 0.019]
+$M=0.3$ [-0.029, 0.021] [-0.029, 0.021] [-0.029, 0.021] [-0.029, 0.021]
+$M=0.4$ [-0.032, 0.023] [-0.032, 0.023] [-0.032, 0.023] [-0.032, 0.023]
+$M=0.5$ [-0.036, 0.027] [-0.036, 0.027] [-0.036, 0.027] [-0.036, 0.027]
+$M=1.0$ [-0.059, 0.049] [-0.059, 0.049] [-0.059, 0.049] [-0.059, 0.049]
+$M=1.5$ [-0.083, 0.073] [-0.083, 0.073] [-0.083, 0.073] [-0.083, 0.073]
+$M=2.0$ [-0.108, 0.098] [-0.108, 0.098] [-0.108, 0.098] [-0.108, 0.098]
+  • Notes: This table presents 90% robust confidence intervals calculated using the methods proposed by RambachanRoth2023.
+We assume limited violations of parallel trends based on the relative magnitude of pre-treatment trends.
+We revisit estimates from Table 4 Columns (1-4).
+The parameter $M$ governs the amount of slope violation allowed between consecutive periods, relative to the maximum observed violation in the pre-treatment period.
+$M=0$ corresponds to the standard DDD estimate (assuming perfect parallel trends).
+A bolded entry indicates the value of $M$ at which the confidence interval first includes zero (i.e., the result becomes statistically insignificant).
+Panel A reports estimates for men, while Panel B for women.
+Online Appendix*: Effects of State and Local Sexual Orientation Anti-Discrimination Laws on Labor Market Differentials
+In an earlier version of the paper, we estimated all regressions separately by gender.
+The point estimates and standard errors are identical to those from the fully interacted specification, but the separate-gender approach does not allow us to test for differences efficiently.
+For transparency, we report those results here.
+Table O1: Effect of Anti-Discrimination Laws: Extensive & Intensive Margin of Labor Supply
+(1) (2) (3) (4) (5) (6)
+Extensive Margin Intensive Margin
+Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
+Panel A: Men
+ENDA $\times$ SSP 0.0379∗∗∗ 0.0370∗∗∗ 1.1289∗∗ 1.5089∗∗ -0.0920 -0.0719
+(0.0130) (0.0135) (0.4978) (0.6268) (0.3727) (0.2820)
+Observations 3,798,542 3,798,542 3,798,542 3,798,542 3,073,797 3,073,797
+Mean (M) 0.825 0.790 35.53 39.96 42.41 47.69
+Employed Only X X
+Panel B: Women
+ENDA $\times$ SSP -0.0037 -0.0050 0.9519 -0.4143 0.7017∗∗ -0.8915∗∗∗
+(0.0116) (0.0135) (0.5324) (0.6230) (0.3074) (0.3350)
+Observations 3,982,816 3,982,816 3,982,816 3,982,816 2,691,783 2,691,783
+Mean (F) 0.818 0.782 34.42 39.51 41.25 47.27
+Employed Only X X
+  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
+We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment.
+The regressions are estimated separately by sex: Panel A for men, while Panel B for women.
+The coefficients indicate the impact of anti-discrimination laws.
+The analysis is divided into two parts: Columns (1-2) focus on the extensive margin, examining changes in labor force participation and employment status as a result of the laws.
+Columns (3-6), on the other hand, explore the intensive margin by leveraging variations in the number of weeks worked per year and average weekly working hours.
+The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples, respectively.
+Standard errors are clustered at the county level.
+$p<0.10$, ∗∗ $p<0.05$, $^{{}^{***}}$ $p<0.01$.
+Table O2: Effect of Anti-Discrimination Laws: Wages and Earnings
+Hourly Wage Annual Earnings
+(1) (2) (3) (4) (5) (6) (7) (8)
+Percentile $\geq$25th $\geq$50th $\geq$75th Percentile $\geq$25th $\geq$50th $\geq$75th
+Panel A: Men
+ENDA $\times$ SSP 2.6310∗∗ 0.0366∗∗∗ 0.0326 0.0232 2.9315∗∗∗ 0.0365∗∗∗ 0.0467∗∗ 0.0334∗∗
+(1.1766) (0.0138) (0.0182) (0.0178) (1.1014) (0.0139) (0.0185) (0.0154)
+Observations 3,798,542 3,798,542 3,798,542 3,798,542 3,798,542 3,798,542 3,798,542 3,798,542
+Mean (M) 51.51 0.791 0.557 0.305 51.69 0.791 0.561 0.305
+Panel B: Women
+ENDA $\times$ SSP -0.3333 0.0045 -0.0117 -0.0061 -0.5994 0.0051 -0.0204 -0.0016
+(0.8442) (0.0118) (0.0143) (0.0154) (0.8307) (0.0118) (0.0134) (0.0161)
+Observations 3,982,816 3,982,816 3,982,816 3,982,816 3,982,816 3,982,816 3,982,816 3,982,816
+Mean (F) 48.73 0.799 0.513 0.247 48.68 0.798 0.515 0.241
+  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
+We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment.
+The regressions are estimated separately by sex: Panel A for men, while Panel B for women.
+The coefficients in the first row indicate the impact of anti-discrimination laws.
+Columns (1–4) report estimates for hourly real wage ranks, while Columns (5–8) report estimates for annual real wage ranks.
+In Columns 2–4 and 6–8, the estimates are indicator variables for being above the 25th, 50th, and 75th percentiles, respectively.
+All income variables are measured in 2019 dollars and expressed as percentile ranks.
+The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples, respectively.
+Standard errors are clustered at the county level.
+$p<0.10$, ∗∗ $p<0.05$, $^{{}^{***}}$ $p<0.01$.
+Table O3: Effects of Anti-Discrimination Laws on Labor Supply: Analysis of Same-Sex Marriage Legalization
+(1) (2) (3) (4) (5) (6)
+Extensive Margin Intensive Margin
+Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
+Panel A: Men
+ENDA $\times$ SSP 0.0283 0.0311 0.8892 0.4945 -0.1655 -0.7909∗∗
+(0.0186) (0.0193) (0.6504) (0.8716) (0.4310) (0.3950)
+Observations 2,553,515 2,553,515 2,553,515 2,553,515 2,058,535 2,058,535
+Mean (M) 0.818 0.778 35.39 39.44 42.40 47.27
+Employed Only X X
+Panel B: Women
+ENDA $\times$ SSP -0.0157 -0.0215 1.1714 -0.6654 1.0376∗∗∗ -1.2330∗∗∗
+(0.0153) (0.0158) (0.6724) (0.7928) (0.3679) (0.3998)
+Observations 2,665,702 2,665,702 2,665,702 2,665,702 1,802,056 1,802,056
+Mean (F) 0.817 0.774 34.44 39.13 41.28 46.85
+Employed Only X X
+  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
+We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment.
+The regressions are estimated separately by sex: Panel A for men, while Panel B for women.
+Sample is limited to observations before same-sex marriage legalization.
+The coefficients indicate the impact of anti-discrimination laws.
+The analysis is divided into two parts: Columns (1-2) focus on the extensive margin, examining changes in labor force participation and employment status as a result of the laws.
+Columns (3-6), on the other hand, explore the intensive margin by leveraging variations in the number of weeks worked per year and average weekly working hours.
+The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples, respectively.
+Standard errors are clustered at the county level.
+$p<0.10$, ∗∗ $p<0.05$, $^{{}^{***}}$ $p<0.01$.
+Table O4: Effects of Anti-Discrimination Laws on Labor Supply: Analysis of Non-Movers
+(1) (2) (3) (4) (5) (6)
+Extensive Margin Intensive Margin
+Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
+Panel A: Men
+ENDA $\times$ SSP 0.0421∗∗∗ 0.0424∗∗∗ 1.2112∗∗ 1.7705∗∗ -0.0583 0.2447
+(0.0145) (0.0153) (0.5686) (0.7283) (0.4027) (0.2938)
+Observations 3,414,649 3,414,649 3,414,649 3,414,649 2,748,449 2,748,449
+Mean (M) 0.823 0.791 35.44 39.99 42.45 47.89
+Employed Only X X
+Panel B: Women
+ENDA $\times$ SSP 0.0050 0.0048 1.1290 0.1140 0.7276∗∗ -0.5031
+(0.0142) (0.0147) (0.6188) (0.7469) (0.3082) (0.3944)
+Observations 3,611,058 3,611,058 3,611,058 3,611,058 2,437,454 2,437,454
+Mean (F) 0.815 0.783 34.34 39.57 41.32 47.55
+Employed Only X X
+  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
+We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment.
+The regressions are estimated separately by sex: Panel A for men, while Panel B for women.
+Sample is limited to individuals who have not moved in the past year.
+The coefficients indicate the impact of anti-discrimination laws.
+The analysis is divided into two parts: Columns (1-2) focus on the extensive margin, examining changes in labor force participation and employment status as a result of the laws.
+Columns (3-6), on the other hand, explore the intensive margin by leveraging variations in the number of weeks worked per year and average weekly working hours.
+The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples, respectively.
+Standard errors are clustered at the county level.
+$p<0.10$, ∗∗ $p<0.05$, $^{{}^{***}}$ $p<0.01$.
+Table O5: Effects of Anti-Discrimination Laws on Labor Supply: Analysis of Metro Areas
+(1) (2) (3) (4) (5) (6)
+Extensive Margin Intensive Margin
+Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
+Panel A: Men
+ENDA $\times$ SSP 0.0352∗∗ 0.0398∗∗ 0.8139 1.4503 -0.4661 -0.2430
+(0.0159) (0.0162) (0.5417) (0.7402) (0.3735) (0.3454)
+Observations 2,278,474 2,278,474 2,278,474 2,278,474 1,904,970 1,904,970
+Mean (M) 0.847 0.812 36.50 41.09 42.62 47.96
+Employed Only X X
+Panel B: Women
+ENDA $\times$ SSP -0.0262∗∗ -0.0284∗∗ -0.4601 -1.3804∗∗ -0.4479 -1.3355∗∗∗
+(0.0115) (0.0129) (0.7241) (0.6505) (0.6670) (0.3471)
+Observations 2,376,723 2,376,723 2,376,723 2,376,723 1,625,459 1,625,459
+Mean (F) 0.834 0.801 35.20 40.42 41.43 47.52
+Employed Only X X
+  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
+We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment.
+The regressions are estimated separately by sex: Panel A for men, while Panel B for women.
+Sample is limited to individuals who live in metropolitan areas.
+The coefficients indicate the impact of anti-discrimination laws.
+The analysis is divided into two parts: Columns (1-2) focus on the extensive margin, examining changes in labor force participation and employment status as a result of the laws.
+Columns (3-6), on the other hand, explore the intensive margin by leveraging variations in the number of weeks worked per year and average weekly working hours.
+The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples, respectively.
+Standard errors are clustered at the county level.
+$p<0.10$, ∗∗ $p<0.05$, $^{{}^{***}}$ $p<0.01$.
+Table O6: Effects of Anti-Discrimination Laws on Labor Supply: Analysis of ACS Redesign
+(1) (2) (3) (4) (5) (6)
+Extensive Margin Intensive Margin
+Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
+Panel A: Men
+ENDA $\times$ SSP 0.0459 0.0492 1.5759 3.1075∗∗ -0.2548 0.8834∗∗∗
+(0.0255) (0.0267) (1.2059) (1.2909) (0.7122) (0.3361)
+Observations 2,966,542 2,966,542 2,966,542 2,966,542 2,396,977 2,396,977
+Mean (M) 0.837 0.802 35.66 40.48 42.41 48.11
+Employed Only X X
+Panel B: Women
+ENDA $\times$ SSP 0.0217 0.0272 1.3482 0.6313 0.5888 -0.1525
+(0.0230) (0.0245) (1.1363) (1.2235) (0.4428) (0.4731)
+Observations 3,119,738 3,119,738 3,119,738 3,119,738 2,100,379 2,100,379
+Mean (F) 0.824 0.787 34.27 39.72 41.09 47.51
+Employed Only X X
+  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
+We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment.
+The regressions are estimated separately by sex: Panel A for men, while Panel B for women.
+Sample is limited to cover 2008 through 2019.
+The coefficients indicate the impact of anti-discrimination laws.
+The analysis is divided into two parts: Columns (1-2) focus on the extensive margin, examining changes in labor force participation and employment status as a result of the laws.
+Columns (3-6), on the other hand, explore the intensive margin by leveraging variations in the number of weeks worked per year and average weekly working hours.
+The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples, respectively.
+Standard errors are clustered at the county level.
+$p<0.10$, ∗∗ $p<0.05$, $^{{}^{***}}$ $p<0.01$.
+Table O7: Effects of Anti-Discrimination Laws on Labor Supply: Analysis of Unweighted Estimates
+(1) (2) (3) (4) (5) (6)
+Extensive Margin Intensive Margin
+Labor Force Employed Weekly Hours Weeks Worked Weekly Hours Weeks Worked
+Panel A: Men
+ENDA $\times$ SSP 0.0265 0.0251 0.8506 1.0802 -0.0666 0.0419
+(0.0144) (0.0149) (0.4860) (0.6116) (0.3540) (0.2870)
+Observations 3,798,542 3,798,542 3,798,542 3,798,542 3,073,797 3,073,797
+Mean (M) 0.825 0.790 35.53 39.96 42.41 47.69
+Employed Only X X
+Panel B: Women
+ENDA $\times$ SSP -0.0085 -0.0112 0.7029 -0.5207 0.6289∗∗ -0.7879∗∗
+(0.0102) (0.0118) (0.4860) (0.5798) (0.3004) (0.3316)
+Observations 3,982,816 3,982,816 3,982,816 3,982,816 2,691,783 2,691,783
+Mean (F) 0.818 0.782 34.42 39.51 41.25 47.27
+Employed Only X X
+  • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
+We compare individuals in same-sex partnerships with those in different-sex partnerships within four years of ENDA enactment.
+The regressions are estimated separately by sex: Panel A for men, while Panel B for women.
+The coefficients indicate the impact of anti-discrimination laws.
+The analysis is divided into two parts: Columns (1-2) focus on the extensive margin, examining changes in labor force participation and employment status as a result of the laws.
+Columns (3-6), on the other hand, explore the intensive margin by leveraging variations in the number of weeks worked per year and average weekly working hours.
+Standard errors are clustered at the county level.
+$p<0.10$, ∗∗ $p<0.05$, $^{{}^{***}}$ $p<0.01$.
+Table O8: Effects of Anti-Discrimination Laws on Labor Supply: Analysis of Same-Sex Marriage Fixed Effects
 (1) (2) (3) (4) (5) (6)
@@
 Observations 3,798,599 3,798,599 3,798,599 3,798,599 3,073,844 3,073,844
-Mean of DV 0.8254 0.7903 35.5714 39.9865 42.4118 47.6751
+Mean (M) 0.825 0.790 35.53 39.96 42.41 47.69
 Employed Only X X
@@
 Observations 3,982,894 3,982,894 3,982,894 3,982,894 2,691,849 2,691,849
-Mean of DV 0.8200 0.7846 34.5069 39.5937 41.2556 47.2540
+Mean (F) 0.818 0.782 34.42 39.51 41.25 47.27
 Employed Only X X
@@
 Columns (3-6), on the other hand, explore the intensive margin by leveraging variations in the number of weeks worked per year and average weekly working hours.
-We exclude $\sum_{s=1}^{S}\lambda_{s}\bigl{[}\mathrm{SSP}_{ipct}\times\mathrm{SSM}_{st}\times\mathbf{1}_{\text{state}(c)=s}\bigr{]}$ from Equation (1) for this table.
+This table excludes the fully interacted same-sex-marriage term from Equation (1).
+The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples, respectively.
 Standard errors are clustered at the county level.
 $p<0.10$, ∗∗ $p<0.05$, $^{{}^{***}}$ $p<0.01$.
-Table A2: Anti-Discrimination Laws on Labor Supply with Occupation FEs
+Table O9: Anti-Discrimination Laws on Labor Supply with Occupation FEs
 (1) (2) (3) (4) (5) (6)
@@
 Observations 3,798,542 3,798,542 3,798,542 3,798,542 3,073,797 3,073,797
-Mean of DV 0.8254 0.7903 35.5714 39.9865 42.4118 47.6751
+Mean (M) 0.825 0.790 35.53 39.96 42.41 47.69
 Employed Only X X
@@
 Observations 3,982,816 3,982,816 3,982,816 3,982,816 2,691,783 2,691,783
-Mean of DV 0.8200 0.7846 34.5069 39.5937 41.2556 47.2540
+Mean (F) 0.818 0.782 34.42 39.51 41.25 47.27
 Employed Only X X
@@
 Columns (3-6), on the other hand, explore the intensive margin by leveraging variations in the number of weeks worked per year and average weekly working hours.
+The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples, respectively.
 Standard errors are clustered at the county level.
 $p<0.10$, ∗∗ $p<0.05$, $^{{}^{***}}$ $p<0.01$.
-Table A3: Anti-Discrimination Laws on Pay with Occupation FEs
+Table O10: Anti-Discrimination Laws on Pay with Occupation FEs
 Hourly Wage Annual Earnings
 (1) (2) (3) (4) (5) (6) (7) (8)
-Percentile $>$25th $>$50th $>$75th Percentile $>$25th $>$50th $>$75th
+Percentile $\geq$25th $\geq$50th $\geq$75th Percentile $\geq$25th $\geq$50th $\geq$75th
 Panel A: Men
-ENDA $\times$ SSP 1.9105∗∗ 0.0241∗∗ 0.0262 0.0239 2.2973∗∗∗ 0.0241∗∗ 0.0445∗∗∗ 0.0340∗∗
+ENDA $\times$ SSP 1.8922∗∗ 0.0243∗∗ 0.0244 0.0200 2.2207∗∗∗ 0.0242∗∗ 0.0391∗∗∗ 0.0308∗∗
 (0.9189) (0.0113) (0.0146) (0.0141) (0.8331) (0.0113) (0.0152) (0.0128)
 Observations 3,798,542 3,798,542 3,798,542 3,798,542 3,798,542 3,798,542 3,798,542 3,798,542
-Mean of DV 52.0273 0.7917 0.5653 0.3121 52.4128 0.7917 0.5746 0.3163
+Mean (M) 51.51 0.791 0.557 0.305 51.69 0.791 0.561 0.305
 Panel B: Women
-ENDA $\times$ SSP -0.9428 -0.0023 -0.0232∗∗ -0.0096 -1.1544 -0.0023 -0.0217 -0.0106
+ENDA $\times$ SSP -0.9026 -0.0024 -0.0182 -0.0116 -1.1521 -0.0018 -0.0274∗∗ -0.0059
 (0.6989) (0.0097) (0.0121) (0.0146) (0.6519) (0.0097) (0.0111) (0.0155)
 Observations 3,982,816 3,982,816 3,982,816 3,982,816 3,982,816 3,982,816 3,982,816 3,982,816
-Mean of DV 49.3832 0.8002 0.5246 0.2564 49.5588 0.8002 0.5324 0.2551
+Mean (F) 48.73 0.799 0.513 0.247 48.68 0.798 0.515 0.241
   • Notes: Data is derived from the American Community Survey (ACS) conducted annually between 2005 and 2019.
@@
 All income variables are measured in 2019 dollars and expressed as percentile ranks.
+The “Mean (F)” and “Mean (M)” rows report pre-treatment averages of the outcome variables (including never-treated counties) for women and men in same-sex couples, respectively.
 Standard errors are clustered at the county level.
 $p<0.10$, ∗∗ $p<0.05$, $^{{}^{***}}$ $p<0.01$.
-Appendix Figures
-((a)) Labor Force Participation
-((b)) Employment
-Figure A1: Labor Supply Impacts on the Extensive Margin: Same-Sex Marriage Legalization
-Notes: Event study plots on the difference in labor supply between people in same-sex partnerships and different-sex partnerships broken down by sex following the county-level and state-level anti-discrimination laws.
-Sample is limited to observations before same-sex marriage legalization.
-Coefficients with 95% confidence intervals.
-Standard errors are clustered at the county level.
-((a)) Labor Force Participation
-((b)) Employment
-Figure A2: Labor Supply Impacts on the Extensive Margin: Non-Movers
-Notes: Event study plots on the difference in labor supply between people in same-sex partnerships and different-sex partnerships broken down by sex following the county-level and state-level anti-discrimination laws.
-Sample is limited to individuals who have not moved in the past year.
-Coefficients with 95% confidence intervals.
-Standard errors are clustered at the county level.
-((a)) Labor Force Participation
-((b)) Employment
-Figure A3: Labor Supply Impacts on the Extensive Margin: Metropolitan Areas
-Notes: Event study plots on the difference in labor supply between people in same-sex partnerships and different-sex partnerships broken down by sex following the county-level and state-level anti-discrimination laws.
-Sample is limited to individuals who live in metropolitan areas.
-Coefficients with 95% confidence intervals.
-Standard errors are clustered at the county level.
-((a)) Labor Force Participation
-((b)) Employment
-Figure A4: Labor Supply Impacts on the Extensive Margin: 2008-2019
-Notes: Event study plots on the difference in labor supply between people in same-sex partnerships and different-sex partnerships broken down by sex following the county-level and state-level anti-discrimination laws.
-Sample is limited to cover 2008 through 2019.
-Coefficients with 95% confidence intervals.
-Standard errors are clustered at the county level.
-((a)) Labor Force Participation
-((b)) Employment
-Figure A5: Labor Supply Impacts on the Extensive Margin: Weighting Schemes
-Notes: Event study plots on the difference in labor supply between people in same-sex partnerships and different-sex partnerships broken down by sex following the county-level and state-level anti-discrimination laws.
-We compare estimates weighted by ACS weights and without.
-Coefficients with 95% confidence intervals.
-Standard errors are clustered at the county level.
-Online Appendix*: Effects of State and Local Sexual Orientation Anti-Discrimination Laws on Labor Market Differentials
-In the original version of our paper, we employed a traditional difference-in-differences (DiD) setup, incorporating a range of controls to assess the impact of anti-discrimination legislation on labor market outcomes for LGB workers.
-This conventional approach served as the foundational analysis for our initial findings.
-To ensure transparency and thoroughness in our research, we are including these initial estimates in this section.
-Furthermore, in response to recent advancements in DiD methodologies, we use Borusyak et al.
-[2021] in the main text to estimate the impact of anti-discrimination laws.
-Our analysis reveals that the application of these modern DiD techniques does not substantially alter the point estimates we initially obtained.
-This consistency across methodological frameworks underscores the robustness of our findings, reinforcing the reliability of our conclusions regarding the effects of anti-discrimination laws on LGB labor market outcomes.
-In the TWFE setup, we control for sentiment toward LGB workers in a given state at a given year by including polling information on support for same-sex marriages.
-Unfortunately, we cannot observe polling at a local level, as such the polling measure is an incomplete proxy for sentiment toward LGB workers.
-We collected every poll that Pew Research Center published from 2005-2016 to pull each poll regarding support for same-sex marriage.
-The 28 polls asking about same-sex marriage were aggregated to get the percentage of people that supported same-sex marriage by state by year.
-The polling information contains data on every state for every year, except for Alaska and Hawaii, which are missing polling information for 2005-2008.
-O.4 Identification Strategy & Estimation
-The differential roll-out of anti-discrimination laws by state and locality over time lends itself to a difference-in-differences framework with the main outcomes of interest as the change in labor supply and pay between two groups, workers in same-sex versus different-sex partnerships.
-This strategy uses the variation presented in Figure 1, which shows how the laws changed over time by state and county.
-Formally, we estimate this equation on labor supply:
-$LS_{ict}=\alpha_{0}ENDA_{ct}+\alpha_{1}SSP_{i}+\alpha_{2}ENDA_{ct}\cdot SSP_{i}+\alpha_{3}X_{i}+\alpha_{4}\gamma_{st}+\mu_{c}+\delta_{t}+\epsilon_{itcs}$ (4)
-Where $LS_{ict}$ is the labor force participation or employment status for person $i$ in county $c$ in year $t$.
-$ENDA_{ct}$ is an indicator for if county $c$ has an employment anti-discrimination law on sexual orientation or is covered by a state-wide anti-discrimination law in year $t$, and $SSP_{it}$ is an indicator for if person $i$ is in a same-sex partnership.
-$X_{i}$ is a vector of person-specific covariates including race, children, along with age and education level - each represented up to a quadratic term to capture potential nonlinear effects.
-$\gamma_{st}$ is a vector for state $s$ covariates on the legality of same-sex marriage in year $t$ as well as polling for the support of same-sex marriage.
-It includes the percentage of people that “strongly support,” “support,” “oppose,” and “strongly oppose” same-sex marriage in a given state in a given year.
-Finally, $\mu_{c}$ gives county fixed effects, $\delta_{t}$ gives year fixed effects, and $\epsilon_{itcs}$ is the error term [^19] Since the ACS only identifies metro counties, the county fixed effect is close to a pseudo-county fixed effect.
-The rural counties are all compared together with a true county fixed effect for metro area counties..
-Our sample is limited to people aged 25 to 65, and we estimate this model separately by sex using the ACS person weights.
-We cluster the standard errors at the county level, as that is the level of the treatment in this specification.
-[^20] Treatment status is computed using the earliest year of ENDAs at the county or state level, with individuals remaining classified as “treated” thereafter throughout the dataset.
-Given the ACS’s 23-month reference period for earnings, as discussed by Hansen et al.
-[2020], there’s a potential for bias in treatment assignment.
-This extended timeframe might not precisely align with the exact implementation dates of ENDAs, potentially biasing the immediate effects of these policies towards zero.
-The coefficients of interest are $\alpha_{2}$, which will give the effect of the law on the labor supply gap/premium, and $\alpha_{1}$, which gives the labor supply gap/premium in the absence of anti-discrimination law.
-Our specification assumes that the effect of the laws is the same across states and localities, and we will be estimating an average effect.
-We estimate a similar equation for wages:
-$\ln(Y_{ict}+1)=\beta_{0}\text{ENDA}_{ct}+\beta_{1}\text{SSP}_{i}+\beta_{2}\text{ENDA}_{ct}\cdot\text{SSP}_{i}+\beta_{3}X_{i}+\beta_{4}\gamma_{st}+\mu_{c}+\delta_{t}+\epsilon_{itcs}$ (5)
-Where $Y_{ict}$ is the real value, in 2005 dollars, of person $i$’s wages in county $c$ year $t$.
-The rest of the notation is the same as above.
-The coefficients of interest are $\beta_{2}$, which will give the effect of the law on the pay gap/premium, and $\beta_{1}$, which gives the pay gap/premium.
-[^21] The $\ln(Y_{ict}+1)$ specification is adopted to include both employed and unemployed individuals in the analysis, recognizing employment as endogenous to anti-discrimination laws.
-This approach ensures a comprehensive evaluation of the law’s impact on the entire income spectrum, capturing nuances missed by focusing solely on employed workers.
-O.5 TWFE Results
-In Table O1, we present the regression on the extensive margin of labor supply, with Panel A showing the effect of anti-discrimination laws on the labor supply of men in male same-sex partnerships, and Panel B showing the effect for women in female same-sex partnerships.
-We present the full specification in columns (1) and (2) and include a specification without accounting for local anti-discrimination laws in columns (3) and (4) to show how failing to account for local laws would lead to incorrect inference.
-Consistent with the previous literature, we find that women in female same-sex partnerships have a labor supply premium and men in male same-sex partnerships have a labor supply penalty on the extensive margin.
-Women in female same-sex partnerships are 7.3% more likely to be employed and men in male same-sex partnerships are 7.9% less likely to be employed than their different-sex counterparts.
-Anti-discrimination laws have a significant effect in reducing the labor force participation and employment gap for men in male same-sex partnerships.
-The labor force participation rate gap is reduced by 1.3 p.p.
-(18%), and the employment gap is reduced by 1.4 p.p.
-(18%).
-Conversely, the labor force participation and employment premium that women in female same-sex partnerships have over their different-sex counterparts is reduced by 1.3 p.p.
-(16%) and 1 p.p.
-(14%), respectively.
-As seen in columns (3) and (4) of Panel A, estimating the effect of anti-discrimination laws using only state laws may lead to the erroneous conclusion that these laws have no significant effect on the labor force participation and employment gaps for men in male same-sex partnerships.
-However, the effect on the extensive margin of labor supply for women in female same-sex partnerships looks similar when estimating the effect of just state anti-discrimination laws.
-In Table O2, we present the regression results on the intensive margin of labor supply as measured by weekly hours worked and weeks worked.
-Similar to Table 4, we include the results when only estimating the effects of the state laws in columns (3) and (4) to show the contribution of including local laws in estimating the effect of anti-discrimination laws.
-We replicate past findings where men in male same-sex partnerships supply significantly less labor and women in female same-sex partnerships supply significantly more labor on the intensive margin.
-As with Table O1, we find contrasting effects of anti-discrimination laws on male and female outcomes.
-We find that the intensive labor supply gap for men in male same-sex partnerships as measured by weekly hours worked and weeks worked is significantly reduced, and the intensive labor supply premium for women in female same-sex partnerships is significantly reduced as well.
-The effect of anti-discrimination laws on weekly hours worked becomes insignificant for men in male same-sex partnerships and women in female same-sex partnerships when only estimating off of state-wide anti-discrimination laws.
-In Table O3, we present the regression results on the wage gap as measured by hourly wages and annual wage earnings.
-These variables received a log-transformation, to handle zero values so the changes can be interpreted in percentage terms.
-We run the regressions separately by sex and present the regressions showing how the sexual orientation anti-discrimination laws impact both employed and unemployed people in columns (1) and (2) and limit the regressions to only those who are employed in columns (3) and (4).
-Our preferred specification is to look at both employed and unemployed people in the sample as employment is endogenous.
-Conditioning on employed workers would create an endogeneity problem since labor force participation and employment are significantly impacted by anti-discrimination laws.
-Given Table O1, the sample of employed workers could be changing as a result of the law changes, which would lead to a change in the sample composition.
-We present the regressions limited to employed workers to replicate past work on the wage gap/premium where it is common to only examine employed workers.
-We include results by only using state anti-discrimination laws, excluding local laws.
-We present those results in columns (5) - (8).
-Men in male same-sex partnerships see a significant decline in the hourly wage gap by 2.8 p.p.
-(11%) at the 90% significance level and see an insignificant decline in the annual earnings gap by 5.8 p.p.
-On the other hand, women in female same-sex partnerships see a significant change in their annual earnings premium by 13 p.p.
-(16%) and an insignificant reduction in their hourly wage premium.
-Columns (3) and (4) replicate the past literature where employed men in male same-sex partnerships make about 8% to 11% less than their different-sex counterparts and employed women in female same-sex partnerships make about 5% to 15% more than their different-sex counterparts.
-The pay gap/premium is reduced quite significantly once the sample is limited to employed workers, which is unsurprising.
-As shown in Table O1 and Table O2, there are significant differences in labor supply for same-sex and different-sex workers.
-Men in male same-sex partnerships and women in different-sex partnerships are far more likely to be unemployed, out of the labor force, and work fewer hours and weeks, so including their zero wages in the regression will increase the gap/premium.
-There is a significant decrease in the hourly wage gap now and an insignificant decrease in the annual wage gap.
-There is also a significant decrease in the annual earnings premium for women in female same-sex partnerships but no change in their hourly wage premium.
-This result is consistent with the labor supply results in Table O1 and O2 where their intensive and extensive margin of labor supply premium are declining, so the annual earnings decline, but hourly wage does not.
-O.6 Occupation Fixed Effects
-We show that even when controlling for occupation there is a significant change in relative labor supply for men and women in same-sex partnerships and a significant change in relative pay for women in same-sex partnerships, indicating there is a significant effect of these laws even within occupation.
-We present the extensive labor supply and pay regressions with occupation fixed effects in Table O4 and O5.
-In these tables, there is a smaller pay gap for men in male same-sex partnerships and a smaller pay premium for women in female same-sex partnerships when controlling for occupation or industry.
-However, qualitatively the results seem to be similar to the main results in Table O1 and Table O3.
-Men in male same-sex partnerships see a significant reduction in the labor supply gap, and women in female same-sex partnerships see a significant reduction in their labor supply and annual wage premium.
-There is no longer a significant effect on the wage gap for men in male same-sex partnerships.
-These findings suggest that anti-discrimination laws have a significant effect even within occupations.
-The effect does not come from workers shifting into occupations with less discrimination, but even when conditional on occupation choice, there is a significant reduction in differences in labor supply across sexual orientations for men and women.
-O.7 Sorting and Increased Reporting
-One finding from Klawitter and Flatt [1998] suggests that LGB workers sort to areas with anti-discrimination laws.
-If LGB workers were sorting to areas that recently passed anti-discrimination laws, then it could violate the identifying assumption.
-It would be concerning if high-wage LGB workers migrated from unprotected areas to areas that recently passed anti-discrimination laws since it would lower the average wage in untreated counties and increase the average wage in treated counties.
-Another potential concern would be increased reporting for same-sex partnerships.
-Same-sex partnerships, while less stigmatized in 2005-2019 than in previous years, were still heavily stigmatized.
-Anti-discrimination laws could signal to LGB people that their community is more accepting, and people may be more likely to declare that they are in a same-sex partnership.
-This change in reporting could affect the composition of our sample, which in turn could violate the identifying assumption.
-To address this concern, we run the regression below to see if there is an increase in same-sex partnerships in a given county after an anti-discrimination law is passed:
-$\text{SSP}_{ict}=\beta_{0}\text{Laws}_{ct}+\beta_{1}X_{i}+\beta_{2}\gamma_{st}+\mu_{c}+\delta_{t}+\epsilon_{itcs}$ (6)
-In this regression, $\beta_{0}$ is the coefficient of interest, and it will give the effect of anti-discrimination laws passing on the number of same-sex partnerships in a given county whether that is from in-migration or increased reporting of same-sex partnerships.
-We present these results in Table O6.
-We find no significant effect of anti-discrimination laws on the number of same-sex partnerships at the county level, suggesting there is minimal sorting or change in reporting following the passage of anti-discrimination laws.
-Online Appendix Tables
-Table O1: Effect of Anti-Discrimination Laws: Extensive Margin of Labor Supply
-Panel A: Men
-(1) (2) (3) (4)
-Labor Force Employed Labor Force Employed
-Laws*SSP 0.013*** 0.014** 0.007 0.005
-(0.005) (0.005) (0.005) (0.006)
-SSP -0.074*** -0.079*** -0.068*** -0.073***
-(0.003) (0.004) (0.003) (0.004)
-State Laws Only X X
-Observations 6,287,441 6,287,441 6,287,441 6,287,441
-R-squared 0.135 0.114 0.135 0.114
-Panel B: Women
-(1) (2) (3) (4)
-Labor Force Employed Labor Force Employed
-Laws*SSP -0.013** -$0$016*** -0.013**
-(0.006) (0.006) (0.006) (0.006)
-SSP 0.081*** 0.073*** 0.080*** 0.072***
-(0.004) (0.005) (0.004) (0.004)
-State Laws Only X X
-Observations 6,569,373 6,569,373 6,569,373 6,569,373
-R-squared 0.091 0.088 0.091 0.088
-  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex.
-All of the outcome are binary taking a value of 0 or 1.
-The first row of coefficients shows the effect of anti-discrimination on the labor supply gap or premium, and the second row of coefficients gives the labor supply gap or premium.
-Columns (3) and (4) present results when estimating only using state-wide anti-discrimination laws for comparison.
-Standard errors are clustered at the county level.
-$p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
-Table O2: Effect of Anti-Discrimination Laws: Intensive Margin of Labor Supply
-Panel A: Men
-(1) (2) (3) (4)
-Weekly Hours Weeks Worked Weekly Hours Weeks Worked
-Laws*SSP 0.273* 0.089 0.180 -0.0744
-(0.160) (0.080) (0.164) (0.0820)
-SSP -2.196*** -0.575*** -2.106*** -0.477***
-(0.114) (0.056) (0.0923) (0.0480)
-State Laws Only X X
-Observations 5,244,258 5,244,258 5,244,258 5,244,258
-R-squared 0.034 0.011 0.034 0.011
-Panel B: Women
-(1) (2) (3) (4)
-Weekly Hours Weeks Worked Weekly Hours Weeks Worked
-Laws*SSP -0.305** -0.184** -0.213 -0.171**
-(0.136) (0.084) (0.134) (0.0790)
-SSP 2.792*** 0.604*** 2.706*** 0.574***
-(0.105) (0.066) (0.0924) (0.0577)
-State Laws Only X X
-Observations 4,366,603 4,366,603 4,366,603 4,366,603
-R-squared 0.038 0.015 0.038 0.015
-  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex.
-All of the outcome are binary taking a value of 0 or 1.
-The first row of coefficients shows the effect of anti-discrimination on the labor supply gap or premium, and the second row of coefficients gives the labor supply gap or premium.
-Columns (3) and (4) present results when estimating only using state-wide anti-discrimination laws for comparison.
-Standard errors are clustered at the county level.
-$p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
-Table O3: Effect of Anti-Discrimination Laws: Wages and Earnings
-Panel A: Men
-(1) (2) (3) (4) (5) (6) (7) (8)
-Hourly Wage Annual Earnings Hourly Wage Annual Earnings Hourly Wage Annual Earnings Hourly Wage Annual Earnings
-Laws*SSP 0.028* 0.059 0.011 -0.041 0.0139 -0.00404 0.00639 -0.0581
-(0.017) (0.050) (0.015) (0.035) (0.0184) (0.0525) (0.0168) (0.0373)
-SSP -0.244*** -0.757*** -0.082*** -0.114*** -0.233*** -0.716*** -$0**$112***
-(0.012) (0.038) (0.009) (0.023) (0.0108) (0.0354) (0.00745) (0.0184)
-Employed Only X X X X
-State Laws Only X X X X
-Observations 6,287,441 6,287,441 5,244,258 5,244,258 5,244,258 6,287,441 5,244,258 6,287,441
-R-squared 0.157 0.121 0.127 0.041 0.127 0.157 0.041 0.121
-Panel B: Women
-(1) (2) (3) (4) (5) (6) (7) (8)
-Hourly Wage Annual Earnings Hourly Wage Annual Earnings Hourly Wage Annual Earnings Hourly Wage Annual Earnings
-Laws*SSP -0.019 -0.129** -0.024** -0.062** -0.0203 -0.153** -0.0144 -0.0453
-(0.020) (0.066) (0.011) (0.029) (0.0202) (0.0668) (0.0112) (0.0297)
-SSP 0.220*** 0.815*** 0.053*** 0.149*** 0.218*** 0.809*** 0.0448*** 0.132***
-(0.014) (0.049) (0.007) (0.021) (0.0116) (0.0409) (0.00635) (0.0180)
-Employed Only X X X X
-State Laws Only X X X X
-Observations 6,569,373 6,569,373 4,366,603 4,366,603 6,569,373 6,569,373 4,366,603 4,366,603
-R-squared 0.136 0.106 0.123 0.036 0.136 0.106 0.123 0.036
-  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex.
-Columns (3) and (4) are limited to those who are employed.
-All of the are in terms of 2019 dollars and received the log(x+1) transformation to have the interpretation of the coefficient be in terms of percent.
-The first row of coefficients show the effect of anti-discrimination laws on the pay gap or premium, and the second row of coefficients give the pay gap or premium.
-Columns (5) - (8) present results when estimating only using state-wide anti-discrimination laws for comparison.
-Standard errors are clustered at the county level.
-$p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
-Table O4: Anti-Discrimination Laws on Labor Supply with Occupation FEs
-Men Women
-(1) (2) (1) (2)
-Labor Force Employed Labor Force Employed
-Laws*SSP 0.009*** 0.009** -0.009*** -0.009***
-(0.003) (0.004) (0.003) (0.004)
-SSP -0.037*** -0.045*** 0.039*** 0.036***
-(0.002) (0.003) (0.002) (0.003)
-Observations 6,287,441 6,287,441 6,569,373 6,569,373
-R-squared 0.372 0.461 0.517 0.461
-  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex and include four-digit occupation fixed effects.
-All of the outcome are binary taking a value of 0 or 1.
-The first row of coefficients show the effect of anti-discrimination on the labor supply gap or premium, and the second row of coefficients give the labor supply gap or premium.
-Standard errors are clustered at the county level.
-$p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
-Table O5: Anti-Discrimination Laws on Pay with Occupation FEs
-Panel A: Men
-(1) (2) (3) (4)
-Hourly Wage Annual Earnings Hourly Wage Annual Earnings
-Laws*SSP 0.013 0.018 0.011 -0.028
-(0.013) (0.035) (0.013) (0.033)
-SSP -0.113*** -0.351*** -0.043*** -0.051**
-(0.008) (0.025) (0.008) (0.022)
-Employed only X X
-Observations 6,287,441 6,287,441 5,244,258 5,244,258
-R-squared 0.372 0.372 0.248 0.169
-Panel B: Women
-(1) (2) (3) (4)
-Hourly Wage Annual Earnings Hourly Wage Annual Earnings
-Laws*SSP -0.015 -0.104*** -0.024*** -0.063**
-(0.011) (0.034) (0.009) (0.025)
-SSP 0.117*** 0.437*** 0.056*** 0.172***
-(0.008) (0.028) (0.007) (0.018)
-Employed only X X
-Observations 6,569,373 6,569,373 4,366,603 4,366,603
-R-squared 0.513 0.527 0.303 0.233
-  • Notes: Data comes from the 2005-2016 yearly ACS comparing people in same-sex partnerships to people in different-sex partnerships with the regressions run separately by sex and include four-digit occupation fixed effects.
-All of the are in terms of 2019 dollars and received the log(x+1) transformation to have the interpretation of the coefficient be in terms of percent.
-The first row of coefficients show the effect of anti-discrimination on the pay gap or premium, and the second row of coefficients give the pay gap or premium.
-Standard errors are clustered at the county level.
-$p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
-Table O6: Sorting and Reporting
-(1) (2) (3) (4) (5)
-All SSP Male SSP Female SSP Age Years of Education
-ENDA -0.0002 0.0004 -0.0008 0.0257 -0.0212
-(0.0005) (0.0005) (0.0005) (0.0648) (0.0182)
-Observations 16,071,420 7,850,821 8,220,599 16,071,420 16,071,420
-Mean of DV 0.0122 0.0123 0.0122 45.5220 13.8539
-$R^{2}$ 0.0043 0.0081 0.0024 0.0101 0.0467
-  • Notes: Data comes from the 2005-2019 yearly ACS seeing how the number of same-sex partnerships change in a county after the passage of an anti-discrimination law.
-The first column looks at both men and women with the next two columns separating the sexes.
-Column (4) investigates age and (5) looks at years of education.
-Standard errors are clustered at the county level.
-$p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.
-Table O7: Differential Responses between Male and Female Same-Sex Partnerships
-(1) (2) (3) (4)
-One Earner HH Diff in Hours Worked Any Children Number of Children
-Laws*FemSSP -0.00308 0.947*** 0.0344*** 0.0544***
-(0.00868) (0.364) (0.00846) (0.0195)
-FemSSP 0.00445 -0.729** 0.123*** 0.193***
-(0.00717) (0.317) (0.00743) (0.0176)
-Observations 73,181 73,181 73,181 73,181
-R-squared 0.046 0.043 0.122 0.111
-  • Notes: Data comes from the 2005-2016 yearly ACS comparing women in same-sex partnerships to men in same-sex partnerships.
-The first column examines if the household has only one-earner.
-The second column tests the difference in absolute terms of hours worked between the two partners.
-The third and fourth columns examines how households differ with having children.
-The first row of coefficients show the differential effect of anti-discrimination between lesbian and gay households.
-Standard errors are clustered at the county level.
-$p<0.10$, ∗∗ $p<0.05$, ${}^{{}^{***}}$ $p<0.01$.